### v3

# iclr2025_conference.bib

Abstract
As diffusion probabilistic models (DPMs) are being employed as mainstream models for Generative Artificial Intelligence (GenAI), the study of their memorization of training data has attracted growing attention. Existing works in this direction aim to establish an understanding of whether or to what extent DPMs learn via memorization. Such an understanding is crucial for identifying potential risks of data leakage and copyright infringement in diffusion models and, more importantly, for trustworthy application of GenAI. Existing works revealed that conditional DPMs are more prone to training data memorization than unconditional DPMs, and the motivated data extraction methods are mostly for conditional DPMs. However, these understandings are primarily empirical, and extracting training data from unconditional models has been found to be extremely challenging. In this work, we provide a theoretical understanding of memorization in both conditional and unconditional DPMs under the assumption of model convergence. Our theoretical analysis indicates that extracting data from unconditional models can also be effective by constructing a proper surrogate condition. Based on this result, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) that leverages a time-dependent classifier trained on the generated data as a surrogate condition to extract training data from unconditional DPMs. Empirical results demonstrate that our SIDE can extract training data in challenging scenarios where previous methods fail, and it is, on average, over 50% more effective across different scales of the CelebA dataset.
1 Introduction
The diffusion probabilistic models (DPMs) ho2020denoising; sohl2015deep; song2019generative is a family of powerful generative models that learn the distribution of a dataset by first gradually destroying the structure of the data through an iterative forward diffusion process and then restoring the data structure via a reverse diffusion process. Due to their outstanding capability in capturing data distribution, DPMs have become the foundation models for many pioneering Generative Artificial Intelligence (GenAI) products such as Stable Diffusion rombach2022high, DALL-E 3 BetkerImprovingIG, and Sora videoworldsimulators2024. Despite the widespread adoption of DPMs, a potential risk they face is data memorization, i.e., the risk of memorizing a certain proportion of the raw training samples. This could result in the generation of memorized (rather than new) samples via direct copying, causing data leakage, biased generation, privacy breaches, or copyright infringement, as highlighted in the literature Somepalli2022DiffusionAO; somepalli2023understanding; asay2020independent; cooper2024files; perera2023analyzing. A current case argues that Stable Diffusion is a 21st-century collage tool, remixing the copyrighted creations of countless artists whose works were included in the training databutterick2023stable. Furthermore, data memorization also gives rise to data extraction attacks, which is one type of privacy attacks that attempt to extract the training data from a well-trained model. Notably, recent work by carlini2023extracting demonstrated the feasibility of extracting training data samples from DPMs like Stable Diffusion rombach2022high, revealing the potential dangers associated with these models.
Several works have investigated the data memorization phenomenon in diffusion models. For example, it has been observed that there exists a strong correlation between training data memorization and conditional DPMs, i.e., being conditional is more prone to memorization somepalli2023understanding. Gu2023OnMI investigates the influential factors on memorization behaviors via a comprehensive set of experiments. They found that conditioning on random-labeled data can significantly trigger memorization, and unconditional models memorize much less training data. Though inspiring, these understandings are primarily empirical. Moreover, without a unified theoretical understanding of memorization in both conditional and unconditional DPMs, it is extremely difficult to design an effective data extraction method for unconditional DPMs.
In this work, we first introduce a memorization metric to quantify the degree of memorization in DPMs by measuring the overlap between the generated and training data in a point-wise manner. Based on this metric, we present a theoretical framework that explains why conditional generative models memorize more data and why random labeling can lead to increased memorization. Our theoretical analysis indicates that a classifier trained on the same or similar training data can serve as a surrogate condition for unconditional DPMs. By further making the classifier time-dependent (to the diffusion sampling process), we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) to extract training data from unconditional DPMs. We empirically verify the effectiveness of SIDE on CIFAR-10 and different scales of the CelebA dataset (attack results in Figure 1), confirming the accuracy of the theoretical framework.
In summary, our main contributions are:
  • We introduce a novel metric to measure the degree of point-wise memorization in DPMs and present a theoretical framework that explains 1) why conditional DPMs are more prone to memorization, 2) why random labels can lead to more memorization, and 3) implicit labels (e.g., the learned clusters) can serve as a surrogate condition for unconditional DPMs.
  • Based on our theoretical understanding, we propose a novel training data extraction method SIDE that leverages the implicit labels learned by a time-dependent classifier to extract training data from unconditional DPMs.
  • We evaluate the effectiveness of SIDE on CIFAR-10 and various scales of CelebA datasets, and show that, on average, it can outperform the baseline method proposed by carlini2023extracting by more than 50%.
Figure 1: A few examples of the extracted images from a DDPM trained on a subset of the CelebA dataset using our SIDE method. Top: training images; bottom: extracted images.
2 Related Work
Diffusion Probabilistic Models  DPMs sohl2015deep such as Stable Diffusion rombach2022high, DALL-E 3 BetkerImprovingIG, Sora videoworldsimulators2024, Runway rombach2022high, and Imagen saharia2022photorealistic have achieved state-of-the-art performance in image/video generation across a wide range of benchmarks dhariwal2021diffusion. These models can be viewed from two perspectives. The first is score matching song2019generative, where diffusion models learn the gradient of the image distribution song2020score. The second perspective involves denoising DPMs ho2020denoising, which add Gaussian noise at various time steps to clean images and train models to denoise them. To conditionally sample from diffusion models, dhariwal2021diffusion utilizes a classifier to guide the denoising process at each sampling step. Additionally, ho2022classifier introduces classifier-free guidance for conditional data sampling using DPMs.
Memorization in Diffusion Models  Early exploration of memorization in large models was focused on language models carlini2022quantifying; jagielski2022measuring, which has motivated more in-depth research on DPMs somepalli2023understanding; Gu2023OnMI; rahman2024frame. Notably, a recent research on image-generation DPMs: Somepalli2022DiffusionAO found that 0.5-2% of generated images duplicate training samples, a result concurrently reported by carlini2023extracting in broader experiments for both conditional and unconditional diffusion models. Further studies somepalli2023understanding and Gu2023OnMI linked memorization to model conditioning, showing that conditional models are more prone to memorization. Several works have proposed methods to detect and mitigate memorization to address this issue. Wen et al. wen2024detecting proposed a method to detect whether a prompt can trigger memorization by inspecting the magnitude of text-conditional predictions. It delivers high accuracy even at the first generation step, with a single generation per prompt to detect whether a prompt can trigger memorization. Ren et al. ren2024unveiling proposed two metrics based on findings on cross-attention in DPMs to detect memorization. To mitigate memorization, Chen et al.chen2024towards proposed anti-memorization guidance to mitigate memorization during the sampling process. Instead of using guidance, Ren et al. ren2024unveiling mitigate memorization by changing the attention score or masking out summary tokens in the cross-attention layer. Wen et al.wen2024detecting do so by minimizing the magnitude during inference. So far, the motivated data extraction attack or defend methods from these empirical understandings are mostly for conditional DPMs carlini2023extracting; webster2023reproducible, and studies have shown that extracting training data from unconditional DPMs can be much more challenging than that on conditional DPMs Gu2023OnMI. Although the empirical understandings are inspiring, a theoretical understanding of the memorization behaviors of DPMs is missing from the current literature. A recent attempt by ross2024a uses the local intrinsic dimension (LID) metric to characterize memorization in DPMs. However, the finding that lower LID leads to more memorization is only verified on a toy 1-dimensional dataset and only a few generated images, and it fails to address the data extraction challenge on unconditional DPMs. This challenge is exemplified by the brute-force methods currently employed, as discussed in Somepalli2022DiffusionAO; carlini2023extracting, highlighting the inadequacies of existing approaches for extracting data from unconditional diffusion models. In response to these challenges, we introduce a theoretical framework to characterize memorization in DPMs, which further motivates a novel data extraction method for unconditional DPMs.
3 Proposed Theory
In this section, we introduce a memorization metric and provide a theoretical explanation for the universality of data memorization in both conditional and unconditional diffusion models.
3.1 Memorization Metric
Intuitively, the memorization of fixed training data points (i.e., pointwise memorization) can be quantified by the degree of overlap between the generated distribution and the distributions centered at each training data point. Given a generative model and training dataset, we propose the following memorization metric to quantify the memorization of the training data points in the model.
Definition 1 (Pointwise Memorization)
Given a generative model $p_{\theta}$ with parameters $\theta$ and training dataset ${\mathcal{D}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, the degree of memorization in $p_{\theta}$ of ${\mathcal{D}}$ is defined as:
${\mathcal{M}}_{point}({\mathcal{D}};\theta)=\sum_{{\bm{x}}_{i}\in{\mathcal{D}}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}},$ (1)
where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ is the $i$-th training sample, $N$ is the total number of training samples, $p_{\theta}({\bm{x}})$ represents the probability density function (PDF) of the generated samples, and $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the probability distribution centered at training sample ${\bm{x}}_{i}$ within a small radius $\epsilon$.
A straightforward choice for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the Dirac delta function centered at training data point ${\bm{x}}_{i}$: $q({\bm{x}},{\bm{x}}_{i},\epsilon)=\delta({\bm{x}}-{\bm{x}}_{i})$. However, this would make equation (1) uncomputable as the Dirac delta function is zero beyond the $\epsilon$-neighborhood. Alternatively, we could use the Gaussian distribution with a covariance matrix $\epsilon\textit{I}$ (I is the identity matrix) for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$:
$q({\bm{x}},{\bm{x}}_{i},\epsilon)=\frac{1}{\sqrt{(2\pi\epsilon)^{d}}}\exp\left\{-\frac{1}{2\epsilon}({\bm{x}}-{\bm{x}}_{i})^{\top}({\bm{x}}-{\bm{x}}_{i})\right\}.$ (2)
Note that in equation (1), a smaller value of ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ indicates more overlap between the two distributions and thus more memorization. As $\epsilon$ tends to 00, the measured memorization becomes more accurate, with the limit characterizing the intrinsic memorization capability of model $p_{\theta}$.
Semantic Memorization
As a metric, ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ should be monotonous with the actual semantic memorization effect, which refers to the model’s tendency to reproduce unique features of the training samples rather than generating near-duplicate examples. Here, we define semantic memorization in the latent space as it distills the unique semantic information of each training sample.
Definition 2 (Semantic Memorization)
Let ${\mathcal{D}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$ be the dataset, $p_{\theta}(x)$ be the PDF of the generated samples by model $p_{\theta}$, and $\bm{z}$ be the learned latent code for data sample $\bm{x}$. The semantic memorization of model $p_{\theta}$ on ${\mathcal{D}}$ is:
[…]
where $\bm{z_{i}}$ denotes the ground truth latent code of $\bm{x_{i}}$.
${\mathcal{M}}_{semantic}\left({\mathcal{D}};\theta\right)$ measures semantic memorization because it evaluates how well the learned representations ${\bm{z}}$ of training samples align with their ground truth latent codes, capturing the underlying structure of the data rather than simply memorizing the exact duplicates of the training data points. In practice, the ground truth latent codes are known but can be approximated by an independent encoder.
The following theorem formulates the relationship between pointwise and semantic memorization.
Theorem 1
Pointwise memorization ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ is monotonic to semantic memorization, formally:
$\frac{\partial{\mathcal{M}}_{point}({\mathcal{D}};\theta)}{\partial{\mathcal{M}}_{semantic}({\mathcal{D}};\theta)}=\frac{1}{2\epsilon}+\frac{\operatorname{Tr}(\Sigma_{p_{\theta}}^{-1})}{2}>0$ (4)
where $\Sigma_{p_{\theta}}$ is the covariance matrix of the learned latent distribution of the training data.
We first build the relationship map ${\mathcal{M}}_{point}$ into latent space, and then break down equation (4) into separate components and derive each component separately. So, the equation (4) can be simplified to $\frac{1}{2\epsilon}$. The detailed proof can be found in Appendix A.2. Note that the memorization effect studied in this work refers specifically to the pointwise memorization ${\mathcal{M}}_{point}$.
3.2 Theoretical Framework
Based on pointwise memorization, here we present a theoretically framework that explains why conditional DPMs memorize more data. Our theoretical framework is based on the concept of informative labels, which refers to information that can differentiate subsets of data samples. We first give a formal definition of informative labels and then prove their two key properties: 1) they facilitate tighter clustering of samples around their respective means, and 2) they reduce variance in the latent representations. Building upon the two properties, we theoretically show that conditional DPMs memorize more data.
Informative Labels
The concept of informative labels has been previously discussed as class labels Gu2023OnMI. In this work, we introduce a more general definition that encompasses both class labels and random labels as special cases. We define an informative label as follows.
Definition 3 (Informative Label)
Let ${\mathcal{Y}}=\{y_{i},y_{2},\cdots,y_{C}\}$ be the label set for training dataset ${\mathcal{D}}$ with $C$ unique labels. $y_{i}$ is the associated label with ${\bm{x}}_{i}$, and ${\mathcal{D}}_{y=c}=\{{\bm{x}}_{i}:{\bm{x}}_{i}\in\mathcal{D},y_{i}=c\}$ is the subset of training samples shared the same label $y=c$. A label $y=c$ is an informative label if
$|{\mathcal{D}}_{y=c}|<|{\mathcal{D}}|.$ (5)
Here, the labels are not limited to the conventional class labels; they can also be text captions, features, or cluster information that can be used to group the training samples into subsets. The above definition states that an information label should be able to differentiate a subset of samples from others. An extreme case is that all samples have the same label; in this case, the label is not informative. According to our definition, class-wise and random labels are special cases of informative labels. Informative labels can be either explicit like class/random labels and text captions, or implicit like silent features or clusters. Next, we will explore the correlation between informative labels and the clustering effect in the latent space of a generative model with an encoder and decoder. In diffusion models, the encoder represents the forward diffusion process, while the decoder represents the reverse process. Notably, the latent space can be defined at earlier timesteps according to ho2020denoising.
Memorization in Conditional DPMs
Informative labels cause a clustering effect in the latent space of DPMs by providing contextual information that allows the encoder to better differentiate between data samples. When data points are associated with informative labels, the encoder can map them to a latent distribution that accurately reflects their shared characteristics. This results in tighter clusters in the latent space as samples with the same informative label become more concentrated around their respective means. Consequently, the latent representations of these samples exhibit reduced variance, leading to a more structured and organized latent space. We prove these two properties under the assumption of model convergence.
Suppose we have an encoder $f_{\theta_{E}}({\bm{x}})$ and a decoder $f_{\theta_{D}}({\bm{z}})$. The encoder $f_{\theta_{E}}({\bm{x}})$ maps data samples ${\bm{x}}\in{\mathcal{D}}$ to the latent distribution ${\bm{z}}$, which follows a normal distribution ${\mathcal{N}}(\mathbf{{\bm{\mu}}},\mathbf{\Sigma})$: $p_{\theta}({\bm{z}})={\mathcal{N}}(\mathbf{{\bm{\mu}}},\mathbf{\Sigma})$. For ${\bm{x}}_{i}\in{\mathcal{D}}_{y=c}$, the encoder maps ${\bm{x}}_{i}$ to a latent distribution ${\bm{z}}_{c}$ subject to ${\mathcal{N}}(\mathbf{{\bm{\mu}}_{c}},\mathbf{\Sigma_{c}})$. The decoder $f_{\theta_{D}}({\bm{z}})$ maps ${\bm{z}}$ back to the original data samples ${\bm{x}}$. $y_{i}$ is the label of data sample ${\bm{x}}_{i}$. Training a generative $p_{\theta}$ is to optimize the following:
$\min_{\theta}-\sum_{{{\bm{x}}}_{\mathrm{i}}\in\mathrm{{\mathcal{D}}}}{\log{p}_{\theta}\left({{\bm{x}}_{i}}|y_{i}\right)}.$ (6)
Assumption 1
Given suffix training on ${\mathcal{D}}$, the generative model $p_{\theta}$ converges to an optimal solution for equation (6) : $\theta^{*}=\arg\min_{\theta}-\sum_{{\bm{x}}_{\mathrm{i}}\in\mathcal{D}}\log p_{\theta}({\bm{x}}_{i}|y_{i})$.
Based on the above assumption, we can derive the following Proposition 1, with detailed proof deferred to Appendix A.3
Proposition 1
Let ${\bm{z}}$ be the latent space of generative model $p_{\theta}$ conditioned on an informative label $y=c$, the latent representations learned by $p_{\theta}$ under Assumption 1 satisfy:
$\sum_{{\bm{z}}_{i}\in\mathcal{D}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})\leq\sum_{{\bm{z}}_{i}\in\mathcal{D}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}),$ (7)
$\|\mathbf{\Sigma}_{c}\|_{*}\leq\|\mathbf{\Sigma}\|_{*},$ (8)
where $\Sigma$ is the covariance matrix, $\Sigma_{c}$ is the covariance matrix conditioned on informative label $y=c$, $\bm{\mu}_{c}$ denotes the mean of the latent representations ${\bm{z}}$ conditioned on $y=c$ and $\bm{\mu}$ denotes the mean on the overall dataset ${\mathcal{D}}$.
Proposition 1 describes two properties of the learned latent space driven by informative labels: 1) tighter clustering as defined in equation (7) and 2) reduced variance as defined in equation (8). Tighter clustering means that the data samples associated with the same informative label are more closely clustered, which allows the model to more effectively capture and memorize the specific features and patterns relevant to those labels. This proximity in the latent space enhances the model’s ability to recall memorized samples during generation, as the representations are organized around distinct means. Additionally, reduced variance in these clustered representations leads to greater stability, ensuring that the model can consistently reproduce memorized outputs from the clustered latent codes. Based on this understanding, we formalize the relationship between informative labels and memorization in conditional DPMs via the following theorem.
Theorem 2
A generative model $p_{\theta}$ incurs a higher degree of pointwise memorization when conditioned on informative labels $y=c$, mathematically expressed as:
$\lim_{\epsilon\to 0}\frac{{\mathcal{M}}_{point}({\mathcal{D}}_{y=c},\theta_{y=c})}{{\mathcal{M}}_{point}({\mathcal{D}}_{y=c},\theta)}\leq 1$ (9)
where $\theta_{y=c}$ denotes the parameters of the model when trained on dataset ${\mathcal{D}}_{y=c}$.
The proof can be found in Appendix A.4. Theorem 2 states that any form of information labels can incur more memorization in DPMs, including conventional class labels and random labels. As shown in our proof, the informative nature of a label reduces the entropy (or uncertainty) of the data distribution conditioned on that label, leading to a more focused and memorable data representation. This explains the two empirical observations made in existing works Gu2023OnMI: 1) conditional DPMs are more prone to memorization and 2) even random labels can lead to more memorization. It also explains the findings that unconditional models do not replicate data and that text conditioning increases memorization Somepalli2022DiffusionAO; chen2024towards.
Memorization in Unconditional DPMs
According to Theorem 2, one could leverage informative labels to extract training data from conditional DPMs. Intuitively, the Text captions and class labels commonly used to train conditional DPMs are valid informative labels, which we call explicit informative labels. However, unconditional DPMs do not have explicit informative labels and thus are more difficult to extract training data from. Nevertheless, our theory indicates that the learned representation clusters by an unconditional DPM can also serve as a type of informative labels, which we call implict information labels. It means that if we can formulate the clustering information in the training data, we could construct implicit informative labels to help extract training data from unconditional DPMs. This motivates us to propose a new data extraction method SIDE for unconditional DPMs in the next section.
4 Proposed SIDE Method
In this section, we will construct implicit informative labels for unconditional DPMs, convert the implicit labels into explicit ones, and then leverage the explicit labels to extract training data.
4.1 Constructing Implicit Informative Labels
Intuitively, one could use a classifier to generate (predict) implicit labels $y_{I}$ during the sampling process of the diffusion model. The classifier can be a normal classifier trained on the same dataset as the diffusion model. When such a classifier is not available, random labels or representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can also be used as the implicit labels, according to our theoretical analysis in Section 3.2. We assume an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process defined as:
$\mathrm{d}{\bm{x}}=\left[f({\bm{x}},t)-g(t)^{2}\left(\nabla_{\bm{x}}\log p_{\theta}^{t}({\bm{x}})+\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\right)\right]\mathrm{d}t+g(t)\mathrm{d}w,$ (10)
where $x$ represents the state vector, $f(x,t)$ denotes the drift coefficient, $g(t)$ is the diffusion coefficient, $\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$ denotes the gradient of model $p_{\theta}$ given ${\bm{x}}$ at time $t$, and $\mathrm{d}w$ corresponds to the increment of the Wiener process.
We can use the classifier that generates the implicit labels to approximate the gradient in equation (10). However, a known challenge associated with neural network classifiers is their tendency towards miscalibration guo2017calibration. Specifically, the classifier could be overconfident or underconfident about its predictions. To mitigate the potential impact of miscalibration on the sampling process, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output on the diffusion path using power prior as follows:
$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right).$ (11)
Then, we have:
$\mathrm{d}{\bm{x}}=\left[f({\bm{x}},t)-g(t)^{2}\left(\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\right)\right]\mathrm{d}t+g(t)\mathrm{d}w.$ (12)
[…]
Figure 2: An illustration of our proposed Time-Dependent Knowledge Distillation (TDKD) method that trains a time-dependent classifier on a pseudo-labeled synthetic dataset.
4.2 Time-dependent Classifier
In equation (151), the classifier is denoted by $p_{\theta}^{t}(y|{\bm{x}})$ with $t$ being the timestep, implying its time-dependent nature. However, we do not have a time-dependent classifier but only a time-independent classifier. To address this problem, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier. The distillation process is illustrated in Figure 2. TDKD equips the classifier with time-dependent guidance during sampling. It operates in two steps: first, the network architecture is adjusted to accommodate time-dependent inputs, and the structure of the time-dependent module and modification are illustrated in Appendix C; second, a synthetic dataset and pseudo labels are created to facilitate knowledge distillation from the normal classifier to its time-dependent counterpart.
As the original training dataset is unknown, we employ the target DPM to generate a synthetic dataset, following the generative data augmentation techniques chen2023data; chen2024comprehensive. We then use the normal classifier trained on the original dataset to generate pseudo labels for the synthetic dataset. Finally, we train a time-dependent classifier on the labelled synthetic dataset. The objective of this training is defined as following:
$\mathcal{L}_{distil}=D_{KL}\left(p_{\theta}(y_{I}|{\bm{x}}),p_{\theta}^{t}(y_{I}|{\bm{x}}_{t})\right).$ (13)
Overall Procedure of SIDE
With the trained time-dependent classifier $p_{\theta}^{t}(y|{\bm{x}}_{t})$ and the target DPM, our SIDE extracts training data from the model following a conditional generation process. Assume we condition on the label $y=c$. First, we pick a set of values for $\lambda$, called $\mathcal{S}_{\lambda}$, to use in the SIDE attack. Then, we gather $N_{G}$ data samples for each value of $\lambda$ from the set. During each sampling timestep $t$, we compute the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ ($CE(\cdot)$ is the cross-entropy loss), then we use the gradient and the target DPM to reverse the diffusion process. Third, we compute the similarity score for each generated image. Lastly, we evaluate the attack performance using evaluation metrics. Defer to Appendix E for SIDE’s pseudocode.
5 Experiments
In this section, we first describe our experimental setup, introduce the performance metrics, and then present the main evaluation results of our SIDE method. We empirically verify that the memorized images are not from the classifier. We also conduct an ablation study and hyper-parameter analysis to help understand the working mechanism of SIDE.
5.1 Experimental Setup
We evaluate the effectiveness of our method on three datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) na2022unrestricted which consists of 5478 images, 2) a subset of the CelebA (CelebA-25000) liu2015faceattributes which contains 25,000 images and 2) CIFAR-10 containing 50,000 images. All the images are normalized to [-1,1]. We use the AdamW optimizer loshchilov2018decoupled with a learning rate of 1e-4 to train the time-dependent classifier. We train denoising DPMs with a discrete denoising scheduler (DDIM song2021denoising) using the HuggingFace implementation diffusers. All DPMs are trained with batch size 64. We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, respectively. We use ResNet34 He2015DeepRL as the normal classifier.
5.2 Performance Metrics
Determining whether a generated image is a memorized copy of a training image is difficult, as $L_{p}$ distances in the pixel space are ineffective. Previous research addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection Somepalli2022DiffusionAO; Gu2023OnMI. However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization. Here, we propose two new memorization scores to solve these issues: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS metrics).
The AMS and UMS metrics are defined as follows:
$AMS\left({{\mathcal{D}}}_{{\text{gen}}},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{\text{gen}}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)}}{N_{G}},$ (14)
$UMS\left({{\mathcal{D}}}_{{\text{gen}}},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{\text{gen}}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)|}{N_{G}},$ (15)
where ${{\mathcal{D}}}_{\text{gen}}$ is the generated dataset, ${\mathcal{D}}_{\text{train}}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring. $\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria. $\phi({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria. Mathematically, we can represent $\mathcal{F}(x_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)=\mathbb{1}[\max_{x_{j}\in{\mathcal{D}}_{\text{train}}}\gamma(x_{i},x_{j})\geq\alpha\;\&\;\gamma(x_{i},x_{j})\leq\beta]$ $\varphi(x_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)=\{j:x_{j}\in{\mathcal{D}}_{\text{train}},\gamma(x_{i},x_{j})\geq\alpha\;\&\;\gamma(x_{i},x_{j})\leq\beta\}$. $\gamma$ represents the similarity/distance function. For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$ following carlini2023extracting, while for high-resolution datasets, we use the SSCD score as $\gamma$. In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity. The thresholds for the normalized $L_{2}$ ditance carlini2023extracting are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
The AMS averages the similarity scores across generated images, ensuring that memorized images are not overlooked. In contrast, the UMS quantifies distinct memorized instances by evaluating unique matches, thereby accounting for the uniqueness of the memorized images. By further categorizing the two scores into three levels—low, mid, and high—we obtain more comprehensive measurements for different types of memorization.
While carlini2023extracting; chen2024towards introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is essential for assessing different types of copyright infringement, such as character or style copying lee2023talkin; sag2023copyright; sobel2023elements. Additionally, the UMS considers the number of generated images $N_{G}$, which was overlooked in carlini2023extracting. The significance of $N_{G}$ lies in its non-linear impact on the UMS (see Appendix B for proof), indicating that UMS scores should not be compared across different values of $N_{G}$.
Figure 3: A comparison between the original training images (top row) and generated images (bottom row) by our SIDE method. The matches are classified into three categories: low similarity (SSCD score < 0.5), mid similarity (SSCD score between 0.5 and 0.6), and high similarity (SSCD score > 0.6). This classification highlights varying degrees of semantic resemblance among the image pairs.
5.3 Main Results
We compare SIDE with a random baseline and a variant of SIDE that substitutes the time-dependent classifier with a standard (time-independent) classifier. Here, "TD" refers to the time-dependent classifier trained using our proposed TDKD method, "TI" denotes the time-independent classifier, and "OL" indicates training with the original dataset labels.
The "Random" baseline generates images directly using the target unconditional DPM, as described in carlini2023extracting. We average the results across various values of $\lambda$ (defined in equation (151)), ranging from 5 to 9, with a detailed analysis provided in Section 5.3. It is important to note that $\lambda=0$ corresponds to the "Random" baseline.
For each $\lambda$, including $\lambda=0$, we generate 50,000 images for CelebA and 10,000 images for CIFAR-10 to validate our theoretical analysis and the proposed SIDE method. This effort results in one of the largest generated image datasets to date for studying the memorization of DPMs.
Table 1: The AMS (%) and UMS (%) results at low (top), mid (middle), and high (bottom) levels. ‘Random’ denotes the baseline that generates images directly using the target unconditional DPM, while OL-TI is a variant of SIDE which is not trained using TDKD.
Dataset Method Low Similarity Mid Similarity High Similarity
AMS(%) UMS(%) AMS(%) UMS(%) AMS(%) UMS(%)
CelebA-HQ-FI Random 11.656 2.120 0.596 0.328 0.044 0.040
OL-TI 2.649 0.744 0.075 0.057 0.005 0.005
SIDE (Ours) 15.172 2.342 1.115 0.444 0.054 0.044
CelebA-25000 Random 5.000 4.240 0.100 0.100 0.000 0.000
OL-TI 0.164 0.152 0.000 0.000 0.000 0.000
SIDE (Ours) 8.756 6.940 0.224 0.212 0.012 0.012
CIFAR-10 Random 2.470 1.770 0.910 0.710 0.510 0.420
OL-TI 2.460 1.780 0.800 0.680 0.420 0.370
SIDE (Ours) 5.325 2.053 2.495 0.860 1.770 0.560
Effectiveness of SIDE
The AMS and UMS results for the three datasets are presented in Table 1. As shown, SIDE is highly effective in extracting memorized data across all similarity levels, particularly on the CelebA-25000 dataset, a task previously deemed unfeasible due to its large scale.
On CelebA-HQ-FI, SIDE increases mid-level AMS by 87% to 1.115% and UMS by 37% to 0.444%, with an average improvement of 20% across other levels. For the CelebA-25000 dataset, SIDE dramatically enhances AMS and UMS, achieving increases of 75% and 63% for low similarity and 124% and 112% for mid similarity. In the high similarity, SIDE excels at extracting memorized data.
On CIFAR-10, SIDE also outperforms the baselines across all similarity levels. For low similarity, it achieves an AMS of 5.325%, more than double Random’s 2.470%, and a UMS of 2.05% compared to 1.780%. For mid similarity, SIDE reaches 2.495% AMS and 0.860% UMS, significantly higher than Random’s 0.910% AMS and 0.710% UMS, respectively. In the high similarity category, SIDE achieves 1.770% AMS and 0.560% UMS, well ahead of Random’s 0.510% AMS and 0.420% UMS.
Effectiveness of TDKD
As shown in Table  1, time-independent classifiers perform significantly worse than their time-dependent counterparts on the two high-resolution CelebA datasets, achieving only about 10% of the effectiveness of classifiers trained using our TDKD method. This discrepancy arises because time-independent classifiers can provide accurate gradients only at the final timestep, whereas time-dependent classifiers deliver accurate gradients at each timestep. However, the performance gap narrows on the CIFAR-10 dataset due to its simplicity, consisting of only 10 classes. Thus, the gradients produced by the time-independent classifier are less prone to inaccuracy, which mitigates the limitations of the time-independent classifier.
Memorized Images Are Not From the Classifiers
There might be a concern that the extracted images may originate from the classifier rather than the target DPM. We argue that even if a time-dependent classifier could disclose its training images, these images would still originate from the target DPM, as they can also be extracted with the time-dependent classifiers. Furthermore, Table  1 demonstrates that employing a simple classifier reduces memorization compared to the baseline extraction method, indicating that the memorized images do not originate from the classifier.
Figure 4: Hyper-parameter ($\lambda$) analysis on CelebA-HQ-FI. For high similarity, the best $\lambda$ for AMS and UMS are 16 and 13. For other similarity levels, the best $\lambda$ for AMS and UMS is 13.
Impact of the Classifier on SIDE
The classifier used to train SIDE is associated with a certain number of classes. Here, we conduct experiments to explore the relationship between the number of classes (of the classifier) and the extraction performance at a low similarity level, using 1,200 images per class with $\lambda=5$ on the CelebA-HQ-FI dataset. As shown in Table  2 and Figure 7 (Appendix), there exists a strong positive correlation: as the number of classes increases, both AMS and UMS improve. In summary, increasing number of classes positively affects both AMS and UMS, with a stronger impact on AMS. The UMS values reported here differ significantly from Table  1. This is because here, we only generated 1,200 images per class, whereas 50,000 images per class in the previous experiment. Increasing the number of classes improves AMS more than UMS because a higher class count enables the classifier to better differentiate fine details, leading to more accurate matches at low similarity levels. UMS is less affected since it relies more on the diversity across images, which is constrained by the smaller number of images per class in this experiment.
Table 2: This table presents the results of fitting a linear model to the relationship between the number of classes and both AMS and UMS.
Relationship Coefficient (×$\times 10^{-5}$) Intercept $R^{2}$ Correlation Coefficient
Class and AMS 7.4 0.115 0.637 0.80
Class and UMS 6.1 0.090 0.483 0.70
Hyper-parameter Analysis
Here, we test the sensitivity of SIDE to its hyper-parameter $\lambda$. We generate 50,000 images for each integer value of $\lambda$ within the range of [0, 50]. As shown in Figure 4, the memorization score increases at first, reaching its highest, then decreases as $\lambda$ increases. This can be understood from sampling SDE Equation (151). Starting from 0, the diffusion models are unconditional. As $\lambda$ increases, the diffusion models become conditional, and according to Theorem 2, the memorization effect will be triggered. However, when $\lambda$ is too large, the generated images will overfit the classifier’s decision boundaries, leading to a low diversity and ignoring the data distribution. Consequently, the memorization score decreases.
6 Conclusion
In this paper, we introduced a pointwise memorization metric to quantify memorization effects in DPMs. We provided a theoretical analysis of conditional memorization, offering a generalized definition of informative labels and clarifying that random labels can also be informative. We distinguish between explicit labels and implicit labels and propose a novel method, SIDE, to extract training data from unconditional diffusion models by constructing a surrogate condition. The key to this approach is training a time-dependent classifier using our TDKD technique. We empirically validate SIDE on subsets of the CelebA and CIFAR-10 datasets with two new memorization scores: AMS and UMS. We aim for our work to enhance the understanding of memorization mechanisms in diffusion models and inspire further methods to mitigate memorization.
Broader Impacts
Our work proposes a training data extraction method targeted at diffusion models and the effectiveness of our proposed method has far exceeded random methods. This may be exploited by potential malicious users to attack existing generative diffusion models. By effectively avoiding the introduction of additional or unnecessary informative label information into the training process, the risk of a diffusion model being attacked by data extraction can be greatly reduced. Therefore, the contribution of our work to the community outweighs the risks it poses, especially considering that the scale of our extraction method has not reached the same scale of state-of-the-art diffusion models. Moreover, the theoretical and experimental understandings established in this work can help the community build a deeper understanding of the memorization mechanism of diffusion models, and inspire better defense methods.
[…]
Appendix A Proofs
A.1 Preliminaries
If $p(x)$ and $q(x)$ are normal distributions:
$p(\bm{x})=\frac{1}{\sqrt{(2\pi)^{d}\operatorname{det}(\bm{\Sigma}_{p})}}\exp\left\{-\frac{1}{2}(\bm{x}-\bm{\mu_{p}})^{\top}\bm{\Sigma}_{p}^{-1}(\bm{x}-\bm{\mu}_{p})\right\}$ (16)
$q(\bm{x})=\frac{1}{\sqrt{(2\pi)^{d}\operatorname{det}(\bm{\Sigma}_{q})}}\exp\left\{-\frac{1}{2}(\bm{x}-\bm{\mu}_{q})^{\top}\bm{\Sigma}_{q}^{-1}(\bm{x}-\bm{\mu}_{q})\right\}$ (17)
Then we have:
$\mathbb{E}_{\bm{x}\sim p(\bm{x})}\left[\left(\bm{x}-\bm{\mu}_{q}\right)^{\top}\mathbf{\Sigma}_{q}^{-1}\left({x}-\bm{\mu}_{q}\right)\right]$ (18)
$=\mathrm{Tr}\left(\mathbf{\Sigma}_{q}^{-1}\mathbf{\Sigma}_{p}\right)+\left(\bm{\mu}_{p}-\bm{\mu}_{q}\right)^{\top}\mathbf{\Sigma}_{q}^{-1}\left(\bm{\mu}_{p}-\bm{\mu}_{q}\right)$ (19)
$\mathbb{E}_{\bm{x}\sim q\left(\bm{x}\right)}\left[\left(\bm{x}-\bm{\mu}_{q}\right)^{\top}\mathbf{\Sigma}_{q}^{-1}\left(\bm{x}-\bm{\mu}_{q}\right)\right]=d$ (21)
The entropy of $p(x)$:
$H_{p}\left(\bm{x}\right)=\mathbb{E}_{\bm{x}\sim p\left(\bm{x}\right)}[-\log p(x)]=\frac{n}{2}(1+\log 2\pi)+\frac{1}{2}\log\det\mathrm{(}\bm{\Sigma}_{p})$ (22)
The KL divergence between the two distributions is:
$D_{KL}(p(\bm{x})\|q(\bm{x}))$ (23)
$=\frac{1}{2}\left[\left(\bm{\mu}_{p}-\bm{\mu}_{q}\right)^{\top}\bm{\Sigma}_{q}^{-1}\left(\bm{\mu}_{p}-\bm{\mu}_{q}\right)-\log\operatorname{det}\left(\bm{\Sigma}_{q}^{-1}\bm{\Sigma}_{p}\right)+\operatorname{Tr}\left(\bm{\Sigma}_{q}^{-1}\bm{\Sigma}_{p}\right)-d\right]$ (24)
A.2 Proof for theorem 1
We begin by assuming that we have an encoder $f_{\theta_{E}}({\bm{x}})$ and a decoder $f_{\theta_{D}}({\bm{z}})$. The encoder $f_{\theta_{E}}({\bm{x}})$ maps the data samples ${\bm{x}}$ into the latent distribution ${\bm{z}}$, which is modeled by a normal distribution $N\left(\bm{\mu},\bm{\Sigma}\right)$, where ${\bm{z}}\in\mathbb{R}^{d}$ is the latent space of dimension $d$. The decoder $f_{\theta_{D}}({\bm{z}})$ maps the latent variables ${\bm{z}}$ back to the original data samples ${\bm{x}}$. This structure forms the basis of many variational autoencoder (VAE) frameworks, where we aim to optimize the relationship between ${\bm{x}}$ and ${\bm{z}}$ through probabilistic modeling.
Transformation of Probability Distributions
Based on the transformation of probability density functions (PDF) and the method of change of variables for multiple integrals, we can express the likelihoods as follows:
1. Conditional Probability of ${\bm{x}}$ given $y=c$:
$p_{\theta}\left({\bm{x}}\mid y=c\right)$ $=p_{\theta}\left({\bm{z}}\mid y=c\right)\det\left(\frac{\partial{\bm{z}}}{\partial{\bm{x}}}\right)$ (25)
$=p_{\theta}\left({\bm{z}}\mid y=c\right)\det\left(\frac{\partial f_{\theta_{E}}({\bm{x}})}{\partial{\bm{x}}}\right)$ (26)
Here, $p_{\theta}\left({\bm{x}}\mid y=c\right)$ is the probability of the data ${\bm{x}}$ conditioned on the label $y=c$, which depends on the latent variable ${\bm{z}}$. The term $\det\left(\frac{\partial f_{\theta_{E}}({\bm{x}})}{\partial{\bm{x}}}\right)$ is the determinant of the Jacobian matrix of the encoder function $f_{\theta_{E}}({\bm{x}})$, which accounts for the change of variables from ${\bm{x}}$ to ${\bm{z}}$.
2. Marginal Probability of ${\bm{x}}$:
$p_{\theta}\left({\bm{x}}\right)$ $=p_{\theta}\left({\bm{z}}\right)\det\left(\frac{\partial{\bm{z}}}{\partial{\bm{x}}}\right)$ (27)
$=p_{\theta}\left({\bm{z}}\right)\det\left(\frac{\partial f_{\theta_{E}}({\bm{x}})}{\partial{\bm{x}}}\right)$ (28)
This expression captures the marginal distribution of ${\bm{x}}$, which is obtained by marginalizing over the latent variable ${\bm{z}}$.
3. Volume Element Change:
$d{\bm{x}}$ $=\det\left(\frac{\partial{\bm{x}}}{\partial{\bm{z}}}\right)d{\bm{z}}$ (29)
$=\det\left(\frac{\partial f_{\theta_{D}}({\bm{z}})}{\partial{\bm{z}}}\right)d{\bm{z}}$ (30)
$=\det\left(\frac{\partial{\bm{x}}}{\partial f_{\theta_{E}}({\bm{x}})}\right)d{\bm{z}}$ (31)
This represents the volume element transformation between the latent space ${\bm{z}}$ and the data space ${\bm{x}}$. The determinant of the Jacobian of the decoder $f_{\theta_{D}}({\bm{z}})$ relates the volume elements in ${\bm{z}}$ and ${\bm{x}}$.
Objective Function and Change of Variables
Given the above transformations, we can rewrite the memorization objective $\mathcal{M}_{point}(\mathcal{D};\theta)$ as follows:
1. Original Memorization Objective:
$\mathcal{M}_{point}(\mathcal{D};\theta)$ $=\sum_{\bm{x}_{i}\in\mathcal{D}}\int p_{\theta}(\bm{x})\log\frac{p_{\theta}(\bm{x})}{q(\bm{x},\bm{x}_{i},\epsilon)}\,d\bm{x}$ (32)
This measures the difference between the model distribution $p_{\theta}(\bm{x})$ and a perturbed distribution $q(\bm{x},\bm{x}_{i},\epsilon)$ at each point $\bm{x}_{i}$ in the dataset $\mathcal{D}$.
2. Transformed Memorization Objective in Latent Space:
$\mathcal{M}_{semantic}(\mathcal{D};\theta)$ $=\sum_{\bm{z}_{i}\in\mathcal{D}}\int p_{\theta}(\bm{z})\log\frac{p_{\theta}(\bm{z})}{q(\bm{z},\bm{z}_{i},\epsilon)}\,d\bm{z}$ (33)
By applying the change of variables, we transform the objective into the latent space, where the same logic applies, but now the integration is over the latent variables $\bm{z}$ instead of the data space $\bm{x}$.
Monotonicity Derivation
We now derive the monotonicity of the memorization objective with respect to the content-related part $\mathcal{M}_{semantic}(\mathcal{D};\theta)$. Specifically, we want to show that the objective increases monotonically with $\mathcal{M}_{semantic}$:
1. Partial Derivative of the Objective:
$\frac{\partial\mathcal{M}(\mathcal{D};\theta)}{\partial\mathcal{M}_{semantic}(\mathcal{D};\theta)}$ $=\frac{\partial}{\partial\mathcal{M}_{semantic}(\mathcal{D};\theta)}\left(\sum_{\bm{x}_{i}\in\mathcal{D}}\int p_{\theta}(\bm{x})\log\frac{p_{\theta}(\bm{x})}{q(\bm{x},\bm{x}_{i},\epsilon)}\,d\bm{z}\right)$ (34)
2. Expanding the Derivatives:
[…]
3. Applying the Chain Rule:
[…]
4. Evaluating Each Term:
[…]
5. Final Form:
$=\frac{1}{2\epsilon}+\frac{\operatorname{Tr}(\Sigma_{p_{\theta}}^{-1})}{2}$ (38)
A.2.1 Derive $\frac{\partial\mathcal{H}(p_{\theta})}{\partial\mathcal{M}_{semantic}}$
We are given:
  • A multivariate normal distribution $p_{\theta}(\bm{z})=\mathcal{N}(\bm{\mu},\Sigma_{p_{\theta}})$, where $\bm{\mu}$ is the mean vector and $\Sigma_{p_{\theta}}$ is the covariance matrix.
  • The entropy of this distribution is:
    $H(p_{\theta})=\frac{1}{2}\ln\left((2\pi e)^{d}|\Sigma_{p_{\theta}}|\right)=\frac{d}{2}\ln(2\pi e)+\frac{1}{2}\ln|\Sigma_{p_{\theta}}|$ (39)
  • The sum expression:
    ${\mathcal{M}}_{semantic}=\sum_{\bm{z}_{i}\in\mathcal{D}}\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]=|\mathcal{D}|\operatorname{Tr}(\Sigma_{p_{\theta}})+\sum_{i=1}^{|\mathcal{D}|}\|\bm{\mu}-\bm{z}_{i}\|^{2}$ (40)
    where $|\mathcal{D}|$ is the number of data points in $\mathcal{D}$.
Our goal is to find $\frac{\partial H(p_{\theta})}{\partial{\mathcal{M}}_{semantic}}$.
Entropy of a Multivariate Normal Distribution
The entropy of a multivariate normal distribution $p_{\theta}(\bm{z})=\mathcal{N}(\bm{\mu},\Sigma_{p_{\theta}})$ is given by:
[…]
$=\frac{1}{2}\ln\left((2\pi e)^{d}|\Sigma_{p_{\theta}}|\right)$ (42)
Where:
  • $d$ is the dimensionality of the vector $\bm{z}$.
  • $|\Sigma_{p_{\theta}}|$ denotes the determinant of the covariance matrix $\Sigma_{p_{\theta}}$.
Thus, we can write:
$H(p_{\theta})=\frac{d}{2}\ln(2\pi e)+\frac{1}{2}\ln|\Sigma_{p_{\theta}}|$ (43)
Computing the Sum ${\mathcal{M}}_{semantic}$
We have:
${\mathcal{M}}_{semantic}=\sum_{\bm{z}_{i}\in\mathcal{D}}\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]$ (44)
First, compute $\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]$ for each $\bm{z}_{i}$:
Expanding the Squared Norm
$\|\bm{z}-\bm{z}_{i}\|^{2}$ $=(\bm{z}-\bm{z}_{i})^{\top}(\bm{z}-\bm{z}_{i})$ (45)
$=\bm{z}^{\top}\bm{z}-2\bm{z}^{\top}\bm{z}_{i}+\bm{z}_{i}^{\top}\bm{z}_{i}$ (46)
Taking the Expectation
$\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]$ $=\mathbb{E}_{p_{\theta}}\left[\bm{z}^{\top}\bm{z}\right]-2\bm{z}_{i}^{\top}\mathbb{E}_{p_{\theta}}\left[\bm{z}\right]+\bm{z}_{i}^{\top}\bm{z}_{i}$ (47)
$=\operatorname{Tr}\left(\mathbb{E}_{p_{\theta}}\left[\bm{z}\bm{z}^{\top}\right]\right)-2\bm{z}_{i}^{\top}\bm{\mu}+\|\bm{z}_{i}\|^{2}$ (48)
Computing the Expectations
  • First Term: The second moment of $\bm{z}$:
    $\mathbb{E}_{p_{\theta}}\left[\bm{z}\bm{z}^{\top}\right]=\Sigma_{p_{\theta}}+\bm{\mu}\bm{\mu}^{\top}$ (49)
    Therefore:
    $\operatorname{Tr}\left(\mathbb{E}_{p_{\theta}}\left[\bm{z}\bm{z}^{\top}\right]\right)=\operatorname{Tr}(\Sigma_{p_{\theta}})+\operatorname{Tr}(\bm{\mu}\bm{\mu}^{\top})=\operatorname{Tr}(\Sigma_{p_{\theta}})+\|\bm{\mu}\|^{2}$ (50)
  • Second Term: The mean of $\bm{z}$:
    $\mathbb{E}_{p_{\theta}}\left[\bm{z}\right]=\bm{\mu}$ (51)
  • Third Term: Constant term involving $\bm{z}_{i}$:
    $\|\bm{z}_{i}\|^{2}=\bm{z}_{i}^{\top}\bm{z}_{i}$ (52)
Combining the Terms
Substitute back into the expectation:
$\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]$ $=\left(\operatorname{Tr}(\Sigma_{p_{\theta}})+\|\bm{\mu}\|^{2}\right)-2\bm{z}_{i}^{\top}\bm{\mu}+\|\bm{z}_{i}\|^{2}$ (53)
$=\operatorname{Tr}(\Sigma_{p_{\theta}})+\left(\|\bm{\mu}\|^{2}-2\bm{\mu}^{\top}\bm{z}_{i}+\|\bm{z}_{i}\|^{2}\right)$ (54)
$=\operatorname{Tr}(\Sigma_{p_{\theta}})+\|\bm{\mu}-\bm{z}_{i}\|^{2}$ (55)
Thus, for each $\bm{z}_{i}$:
$\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]=\operatorname{Tr}(\Sigma_{p_{\theta}})+\|\bm{\mu}-\bm{z}_{i}\|^{2}$ (56)
Summing Over All Data Points
The total sum ${\mathcal{M}}_{semantic}$ becomes:
${\mathcal{M}}_{semantic}$ $=\sum_{i=1}^{|\mathcal{D}|}\left(\operatorname{Tr}(\Sigma_{p_{\theta}})+\|\bm{\mu}-\bm{z}_{i}\|^{2}\right)$ (57)
$=|\mathcal{D}|\operatorname{Tr}(\Sigma_{p_{\theta}})+\sum_{i=1}^{|\mathcal{D}|}\|\bm{\mu}-\bm{z}_{i}\|^{2}$ (58)
Let’s define the sample variance $\text{Var}_{\text{data}}$:
$\text{Var}_{\text{data}}=\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\|\bm{\mu}-\bm{z}_{i}\|^{2}$ (59)
Then, ${\mathcal{M}}_{semantic}$ can be expressed as:
${\mathcal{M}}_{semantic}=|\mathcal{D}|\operatorname{Tr}(\Sigma_{p_{\theta}})+|\mathcal{D}|\text{Var}_{\text{data}}=|\mathcal{D}|\left(\operatorname{Tr}(\Sigma_{p_{\theta}})+\text{Var}_{\text{data}}\right)$ (60)
Attempting to Relate $H(p_{\theta})$ and ${\mathcal{M}}_{semantic}$
Our challenge is to express $H(p_{\theta})$ as a function of ${\mathcal{M}}_{semantic}$ so that we can compute $\frac{\partial H(p_{\theta})}{\partial{\mathcal{M}}_{semantic}}$. However, we face a difficulty:
  • The entropy $H(p_{\theta})$ depends on $\ln|\Sigma_{p_{\theta}}|$.
  • The sum ${\mathcal{M}}_{semantic}$ depends on $\operatorname{Tr}(\Sigma_{p_{\theta}})$.
For a general covariance matrix $\Sigma_{p_{\theta}}$, there is no direct algebraic relationship between $\operatorname{Tr}(\Sigma_{p_{\theta}})$ and $\ln|\Sigma_{p_{\theta}}|$. Therefore, we need to explore an alternative method.
Computing Derivatives with Respect to $\Sigma_{p_{\theta}}$
The entropy $H(p_{\theta})$ is given by:
$H(p_{\theta})=\frac{d}{2}\ln(2\pi e)+\frac{1}{2}\ln|\Sigma_{p_{\theta}}|$ (61)
To find the derivative of $H(p_{\theta})$ with respect to $\Sigma_{p_{\theta}}$, we proceed as follows:
$\frac{\partial H(p_{\theta})}{\partial\Sigma_{p_{\theta}}}$ $=\frac{\partial}{\partial\Sigma_{p_{\theta}}}\left(\frac{d}{2}\ln(2\pi e)+\frac{1}{2}\ln|\Sigma_{p_{\theta}}|\right)$ (62)
$=\frac{1}{2}\frac{\partial}{\partial\Sigma_{p_{\theta}}}\ln|\Sigma_{p_{\theta}}|$ (63)
$=\frac{1}{2}\Sigma_{p_{\theta}}^{-1}$ (64)
Recall that:
${\mathcal{M}}_{semantic}=|\mathcal{D}|\left(\operatorname{Tr}(\Sigma_{p_{\theta}})+\text{Var}_{\text{data}}\right)$ (65)
Since $\text{Var}_{\text{data}}$ does not depend on $\Sigma_{p_{\theta}}$, the derivative of ${\mathcal{M}}_{semantic}$ with respect to $\Sigma_{p_{\theta}}$ is:
$\frac{\partial{\mathcal{M}}_{semantic}}{\partial\Sigma_{p_{\theta}}}=|\mathcal{D}|\frac{\partial}{\partial\Sigma_{p_{\theta}}}\operatorname{Tr}(\Sigma_{p_{\theta}})$ (66)
Computing $\frac{\partial H(p_{\theta})}{\partial{\mathcal{M}}_{semantic}}$ Using the Chain Rule
Using the chain rule for derivatives:
$\frac{\partial H(p_{\theta})}{\partial{\mathcal{M}}_{semantic}}=\operatorname{Tr}\left(\frac{\partial H(p_{\theta})}{\partial\Sigma_{p_{\theta}}}\cdot\left(\frac{\partial{\mathcal{M}}_{semantic}}{\partial\Sigma_{p_{\theta}}}\right)^{-1}\right)$ (67)
Computing $\left(\frac{\partial{\mathcal{M}}_{semantic}}{\partial\Sigma_{p_{\theta}}}\right)^{-1}$
Since $\frac{\partial{\mathcal{M}}_{semantic}}{\partial\Sigma_{p_{\theta}}}=|\mathcal{D}|I$, its inverse is:
$\left(\frac{\partial{\mathcal{M}}_{semantic}}{\partial\Sigma_{p_{\theta}}}\right)^{-1}=\frac{1}{|\mathcal{D}|}I$ (68)
Combine together, we can derive
$\frac{\partial H(p_{\theta})}{\partial{\mathcal{M}}_{semantic}}=\frac{\operatorname{Tr}(\Sigma_{p_{\theta}}^{-1})}{2|\mathcal{D}|}$ (69)
Thus, we conclude that the memorization objective increases monotonically with respect to the memorization metric.
A.3 Proof for Proposition 1
A.3.1 Proof for equation 8
Covariance Definitions:
$\Sigma$ $=\text{Cov}(Z),\quad\Sigma_{c}=\text{Cov}(Z\mid Y)$ (70)
Here:
  • $\Sigma$ is the overall covariance matrix of $Z$,
  • $\Sigma_{c}$ is the conditional covariance matrix of $Z$ given $Y$.
The goal is to prove that the trace of the conditional covariance matrix is less than or equal to the trace of the overall covariance matrix, i.e.,
$\text{Tr}(\Sigma_{c})\leq\text{Tr}(\Sigma)$ (71)
for all realizations of $Y$.
Proof of Trace Inequality
$\Sigma$ $=\text{Cov}(Z),\quad\Sigma_{c}=\text{Cov}(Z\mid Y)$ (72)
Here:
  • $\Sigma$ is the overall covariance matrix of the random vector $Z\in\mathbb{R}^{n}$,
  • $\Sigma_{c}$ is the conditional covariance matrix of $Z$ given $Y$.
Goal Prove that:
$\text{Tr}(\Sigma_{c})\leq\text{Tr}(\Sigma)$ (73)
for all realizations of $Y$.
We begin by expressing the overall covariance matrix $\Sigma$ in terms of the conditional covariance matrix $\Sigma_{c}$ and the covariance of the conditional expectation of $Z$ given $Y$.
$\Sigma$ $=\text{Cov}(Z)$ (74)
$=\mathbb{E}\left[(Z-\mathbb{E}[Z])(Z-\mathbb{E}[Z])^{T}\right]$ (75)
$=\mathbb{E}\left[\left(Z-\mathbb{E}[Z\mid Y]+\mathbb{E}[Z\mid Y]-\mathbb{E}[Z]\right)\left(Z-\mathbb{E}[Z\mid Y]+\mathbb{E}[Z\mid Y]-\mathbb{E}[Z]\right)^{T}\right]$ (76)
Expanding the Product Inside the Expectation
We expand the product inside the expectation:
$\Sigma$ $=\mathbb{E}\Bigg{[}\underbrace{(Z-\mathbb{E}[Z\mid Y])(Z-\mathbb{E}[Z\mid Y])^{T}}_{\text{Term 1}}$
$\quad+\underbrace{(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}}_{\text{Term 2}}$
$\quad+\underbrace{(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])(Z-\mathbb{E}[Z\mid Y])^{T}}_{\text{Term 3}}$
$\quad+\underbrace{(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}}_{\text{Term 4}}\Bigg{]}$ (77)
Analyzing Each Term
Term 1:
$\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(Z-\mathbb{E}[Z\mid Y])^{T}\right]=\text{Cov}(Z\mid Y)=\Sigma_{c}$ (78)
Term 2:
$\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\right]$ (79)
To evaluate Term 2, we condition on $Y$:
$\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\right]=\mathbb{E}\left[\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\mid Y\right]\right]$ (80)
Inside the inner expectation, $\mathbb{E}[Z\mid Y]-\mathbb{E}[Z]$ is treated as a constant with respect to $Z$, so:
$\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\mid Y\right]$ (81)
$=\left(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z]\right)\mathbb{E}\left[Z-\mathbb{E}[Z\mid Y]\mid Y\right]$ (82)
$=\left(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z]\right)\cdot 0=0$ (83)
Thus:
$\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\right]=0$ (84)
Term 3:
$\mathbb{E}\left[(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])(Z-\mathbb{E}[Z\mid Y])^{T}\right]$ (85)
Similarly, we condition on $Y$:
$\mathbb{E}\left[(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])(Z-\mathbb{E}[Z\mid Y])^{T}\right]$ (86)
$=\mathbb{E}\left[(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])^{T}\mid Y\right]\right]$ (87)
$=\mathbb{E}\left[(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])\cdot 0^{T}\right]$ (88)
$=0$ (89)
Term 4:
$\mathbb{E}\left[(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\right]=\text{Cov}\left(\mathbb{E}[Z\mid Y]\right)=\text{Cov}(\mathbb{E}[Z\mid Y])$ (90)
Combining All Terms
Putting all the terms together:
$\Sigma$ $=\Sigma_{c}+0+0+\text{Cov}(\mathbb{E}[Z\mid Y])$ (91)
$=\Sigma_{c}+\text{Cov}(\mathbb{E}[Z\mid Y])$ (92)
Taking the Trace
Taking the trace on both sides of the covariance decomposition:
$\text{Tr}(\Sigma)$ $=\text{Tr}\left(\Sigma_{c}\right)+\text{Tr}\left(\text{Cov}(Z\mid Y)\right)$ (93)
$=\left[\text{Tr}(\Sigma_{c})\right]+\text{Tr}\left(\text{Cov}(\mathbb{E}[Z\mid Y])\right)$ (94)
Since the trace of a covariance matrix is non-negative:
$\text{Tr}\left(\text{Cov}(\mathbb{E}[Z\mid Y])\right)\geq 0$ (95)
Thus:
$\text{Tr}(\Sigma_{c})\leq\text{Tr}(\Sigma)$ (96)
Conclusion
We have proven that the trace of the conditional covariance matrix $\Sigma_{c}$ is less than or equal to the trace of the overall covariance matrix $\Sigma$. Formally,
$\text{Tr}(\Sigma_{c})\leq\text{Tr}(\Sigma)$ (97)
This result is a multivariate generalization of the variance decomposition, showing that the expected conditional variability of $Z$ given $Y$ does not exceed the overall variability of $Z$.
A.3.2 Proof for equation 7
Centering and Decomposing
By the definition of variance, we can express the sum of squared distances for the conditional mean $\bm{\mu}_{c}$ and overall mean $\bm{\mu}$:
$\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu})$ (98)
$\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu})$ $=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}\left(({\bm{z}}_{i}-\bm{\mu}_{c}+\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c}+\bm{\mu}_{c}-\bm{\mu})\right)$ (99)
$=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}\Big{[}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})$
$\quad+({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$
$\quad+(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})$
$\quad+(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})\Big{]}$ (100)
$=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})$
$\quad+\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$
$\quad+\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})$
$\quad+\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$ (101)
$=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})$
$\quad+(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}1$ (102)
$=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})+|{\mathcal{D}}_{y=c}|(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$ (103)
Expanding this expression, we have:
$=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}\left(({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})+2({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})+(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})\right).$ (104)
Step 2: Simplifying the Inequality
The term $(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$ is a constant for the conditional samples, and the term $2({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$ sums to zero when averaged over the samples in ${\mathcal{D}}_{y=c}$ due to the definition of $\bm{\mu}_{c}$.
Thus, we have:
$\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu})=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})+|{\mathcal{D}}_{y=c}|(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu}),$ (105)
where $|{\mathcal{D}}_{y=c}|$ is the number of samples with label $y=c$.
Step 3: Establishing the Inequality
Since the variance (sum of squared distances) from the conditional mean will always be less than or equal to the variance from the overall mean, we conclude:
$\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})\leq\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}).$ (106)
This establishes that the latent space representation conditioned on an informative label exhibits reduced variance, confirming our initial claim.
Generalized Variance Comparison
In this section, we examine a broader comparison of variance that does not restrict the analysis to the subset ${\mathcal{D}}_{y=c}$ but rather considers the entire dataset ${\mathcal{D}}$. Namely, we prove that:
$\sum_{\bm{z}_{i}\in{\mathcal{D}}_{y=c}}{(}\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})\leq\sum_{\bm{z}_{i}\in{\mathcal{D}}}{(}\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (107)
For a set of data points $D$, the sum of squared deviations from the mean is given by:
$\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (108)
Derivation:
First, let’s write out the sums of squared deviations for the unconditional and conditional cases.
Unconditional Sum of Squared Deviations:
$\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (109)
where $D$ represents the entire dataset of $\bm{z}$ values.
Conditional Sum of Squared Deviations:
$\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})$ (110)
where $D_{y=c}$ represents the subset of data points $\bm{z}_{i}$ where $y=c$.
Covariance Matrix and Sum of Squared Deviations:
The covariance matrix can be related to the sum of squared deviations. For the unconditional case:
$\Sigma=\frac{1}{n}\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})(\bm{z}_{i}-\bm{\mu})^{T}$ (111)
Taking the trace on both sides:
$\text{tr}(\Sigma)=\frac{1}{n}\text{tr}\left(\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})(\bm{z}_{i}-\bm{\mu})^{T}\right)$ (112)
Since the trace of a sum is the sum of the traces:
$\text{tr}(\Sigma)=\frac{1}{n}\sum_{\bm{z}_{i}\in D}\text{tr}\left((\bm{z}_{i}-\bm{\mu})(\bm{z}_{i}-\bm{\mu})^{T}\right)$ (113)
The trace of the outer product of a vector with itself is the sum of squared elements of the vector:
$\text{tr}\left((\bm{z}_{i}-\bm{\mu})(\bm{z}_{i}-\bm{\mu})^{T}\right)=(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (114)
Therefore:
$\text{tr}(\Sigma)=\frac{1}{n}\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (115)
Similarly, for the conditional case:
$\Sigma_{c}=\frac{1}{n_{c}}\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})(\bm{z}_{i}-\bm{\mu}_{c})^{T}$ (116)
Taking the trace:
$\text{tr}(\Sigma_{c})=\frac{1}{n_{c}}\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})$ (117)
Inequality of Traces:
Given that conditioning on $y=c$ provides information about $\bm{z}$, it generally reduces the variance of $\bm{z}$. Mathematically, this can be expressed as:
$\text{tr}(\Sigma_{c})\leq\text{tr}(\Sigma)$ (118)
In terms of sums of squared deviations:
$\frac{1}{n_{c}}\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})\leq\frac{1}{n}\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (119)
Multiplying both sides by their respective sample sizes $n_{c}$ and $n$:
$\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})\leq\frac{n_{c}}{n}\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (120)
Since $n_{c}\leq n$, this further simplifies to:
$\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})\leq\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (121)
A.4 Proof for theorem 2
This section will detail the proof for the theorem 2.
$\lim_{\epsilon\to 0}\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{y=c}}{\int{p_{\theta}}}({\bm{x}}|y=c)\log\frac{p_{\theta}({\bm{x}}|y=c)}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,dx}{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{y=c}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}}}\leq 1$ (122)
Define ${\mathcal{D}}_{y=c}^{z}=\{{\bm{z}}_{i}:f_{\theta_{E}}({\bm{x}}_{i})\in{\mathcal{D}}_{y=c}\}$
${\mathcal{D}}_{y=c}=\{{\bm{x}}_{i}:{\bm{x}}_{i}\in\mathcal{D},y_{i}=c\}$
By using the change of variable theorem, the (122) becomes:
$\lim_{\epsilon\to 0}\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{y=c}}{\int{p_{\theta}}}({\bm{x}}|y=c)\log\frac{p_{\theta}({\bm{x}}|y=c)}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,dx}{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{y=c}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}}}\leq 1$ (123)
$\Rightarrow\,\,\lim_{\epsilon\to 0}\frac{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\int{p_{\theta}}}({\bm{z}}|y=c)\log\frac{p_{\theta}({\bm{z}}|y=c)}{q({\bm{z}},{\bm{z}}_{i},\epsilon)}\,d{\bm{z}}}{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\int{p_{\theta}}}({\bm{z}})\log\frac{p_{\theta}({\bm{z}})}{q({\bm{z}},{\bm{z}}_{i},\epsilon)}\,d{\bm{z}}}\leq 1$ (124)
Since $p_{\theta}({\bm{z}})=N\left(\bm{\mu},\bm{\Sigma}\right)$, it is reasonable to assume that its conditional distribution is also a normal distribution. Then:
$p_{\theta}\left({\bm{z}}|y=c\right)=N\left(\bm{\mu}_{c},\bm{\Sigma}_{c}\right)$ (125)
where $\bm{\mu}_{c}\in\mathbb{R}^{d},\bm{\Sigma}_{c}\in\mathbb{R}^{d\times d}$. Moreover, because $p_{\theta}\left({\bm{z}}|y=c\right)$ depends on label $c$, we can derive the following:
$\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\left({\bm{z}}_{i}-\bm{\mu}_{c}\right)^{\mathrm{T}}\left({\bm{z}}_{i}-\bm{\mu}_{c}\right)}\leq\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\left(\bm{z}_{i}-\bm{\mu}\right)^{\mathrm{T}}\left(\bm{z}_{i}-\bm{\mu}\right)}$ (126)
where $\forall{\bm{z}}_{i}$, $f_{\theta_{D}}({\bm{z}}_{i})\in y_{c}$.
Intuitively, (126) means that the latent code of each training sample conditioned on the label $y=c$ is more centered around the learned latent space of distribution $p_{\theta}({\bm{z}}|y=c)$ than around the distribution $p_{\theta}({\bm{z}})$.
We now look into the KL divergence:
$\int{p_{\theta}}(\bm{z}|y=c)\log\frac{p_{\theta}(\bm{z}|y=c)}{q(\bm{z};\bm{z}_{i})}\,d\bm{z}$ (127)
[…]
$=-\frac{d}{2}\left(1+\log 2\pi\right)-\frac{1}{2}\log\det\left(\bm{\Sigma}_{c}\right)+\mathbb{E}_{\bm{z}\sim p_{\theta}(\bm{z}|y=c)}\left(-\log q\left(\bm{z};\bm{z}_{i}\right)\right)$ (129)
$=\frac{1}{2}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}_{c}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)}{\epsilon}-\log\frac{\det\left(\mathbf{\Sigma}_{c}\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\mathbf{\Sigma}_{c}\right)}{\epsilon}-d\right]$ (130)
We use the SVD decomposition to decompose $\bm{\Sigma}_{c}$:
$\bm{\Sigma}_{c}=U_{c}\varLambda_{c}U_{c}^{\mathrm{T}}$ (131)
And:
$\log\det\mathbf{\Sigma}_{c}=\log\det U_{c}\varLambda U_{c}^{\mathrm{T}}=\log|U_{c}||\varLambda_{c}||U_{c}^{\mathrm{T}}|=\log|\varLambda_{c}|$ (132)
$\mathrm{Tr}\left(\mathbf{\Sigma}_{c}\right)=\mathrm{Tr}\left(U_{c}\varLambda_{c}U_{c}^{\mathrm{T}}\right)=\mathrm{Tr}\left(\varLambda_{c}U_{c}U_{c}^{\mathrm{T}}\right)=\mathrm{Tr}\left(\varLambda_{c}\right)$ (133)
Thus, (130) simplifies to:
$\frac{1}{2}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}_{c}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)}{\epsilon}-\log\frac{\det\left(\varLambda_{c}\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\varLambda_{c}\right)}{\epsilon}-d\right]$ (134)
Similarly:
$\int{p_{\theta}}(\bm{z})\log\frac{p_{\theta}(\bm{z})}{q(\bm{z};\bm{z}_{i})}\,d\bm{z}$ (136)
$=\frac{1}{2}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}\right)}{\epsilon}-\log\frac{\det\left(\mathbf{\Sigma}_{c}\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\mathbf{\Sigma}_{c}\right)}{\epsilon}-d\right]$ (137)
$=\frac{1}{2}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}\right)}{\epsilon}-\log\frac{\det\left(\varLambda\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\varLambda\right)}{\epsilon}-d\right]$ (138)
where
$\bm{\Sigma}=U\varLambda U^{\mathrm{T}}$ (139)
According to the equation (8), the nuclear norm of the two covariance matrices differs. Specifically:
$\|\mathbf{\Sigma}_{c}\|_{*}\leq\|\mathbf{\Sigma}\|_{*}$ (140)
Thus, according to the definition of the nuclear norm, we have:
$\mathrm{Tr}\left(\varLambda_{c}\right)\leq\mathrm{Tr}\left(\varLambda\right)$ (141)
Therefore:
$\lim_{\epsilon\to 0}\frac{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\int{p_{\theta}}({\bm{z}}|y=c)\log\frac{p_{\theta}({\bm{z}}|y=c)}{q({\bm{z}},{\bm{z}}_{i},\epsilon)}\,d{\bm{z}}}}{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\int{p_{\theta}}({\bm{z}})\log\frac{p_{\theta}({\bm{z}})}{q({\bm{z}},{\bm{z}}_{i},\epsilon)}\,d{\bm{z}}}}$ (142)
$\Rightarrow\lim_{\epsilon\to 0}\frac{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}_{c}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)}{\epsilon}-\log\frac{\det\left(\varLambda_{c}\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\varLambda_{c}\right)}{\epsilon}-d\right]}{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}\right)}{\epsilon}-\log\frac{\det\left(\varLambda\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\varLambda\right)}{\epsilon}-d\right]}$ (143)
Using L’Hospital’s rule:
$\lim_{\epsilon\rightarrow 0}\frac{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left[-\frac{\left(\bm{z}_{i}-\bm{\mu}_{c}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)}{\epsilon^{2}}+\frac{d}{\epsilon}-\frac{\mathrm{Tr}\left(\varLambda_{c}\right)}{\epsilon^{2}}\right]}{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left[-\frac{\left(\bm{z}_{i}-\bm{\mu}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}\right)}{\epsilon^{2}}+\frac{d}{\epsilon}-\frac{\mathrm{Tr}\left(\varLambda\right)}{\epsilon^{2}}\right]}$ (144)
$=\frac{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)+\mathrm{Tr}\left(\varLambda_{c}\right)}{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left(\bm{z}_{i}-\bm{\mu}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}\right)+\mathrm{Tr}\left(\varLambda\right)}$ (145)
$\leq 1$ (146)
Appendix B Explanation of the $N_{G}$’s importance in fair comparison
By highlighting the significance of $N_{G}$, UMS allows a fair comparison among extraction methods. We explain this both theoretically and experimentally. Let $N_{\mathrm{umem}}$ denote the number of uniquely memorized images, $M$ the dataset size, and $p_{\gamma}(i)$ the probability that image $i$ is memorized. To find $\mathbb{E}(N_{\mathrm{umem}})$, we introduce a new variable $I_{i}$, representing the generation of image $i$ in $N_{G}$ trials:
$p(I_{i})=1-\left(1-p_{\gamma}(i)\right)^{N_{G}}$ (147)
Using the linearity of expectation, we derive: $\mathbb{E}\left(N_{\mathrm{mem}}\right)$
$\mathbb{E}\left(N_{\mathrm{umem}}\right)$ (148)
$=\mathbb{E}\left(I_{1}\right)+\mathbb{E}\left(I_{2}\right)+\cdots+\mathbb{E}\left(I_{M}\right)$ (149)
$=\sum_{i=1}^{M}{1-\left(1-p_{\gamma}\left(i\right)\right)^{N_{G}}}$ (150)
The importance of $N_{G}$ lies in its impact on the non-linear expectation of uniquely memorized images. Comparing UMS across different $N_{G}$ values is flawed because varying $N_{G}$ leads to different outcomes, underscoring the need for consistent $N_{G}$ values to ensure fair comparisons. In carlini2023extracting, the significance of $N_{G}$ was overlooked, as they only reported the number of uniquely extracted images. Our theoretical analysis accurately aligns with the behavior observed in the experimental data in Fig 5.
Figure 5: Validation of $N_{G}$ ’s significance
Appendix C Refinement ResNet block
Figure 6: Refinement ResNet block with time-dependent module integration. This block diagram depicts the insertion of a time module within a conventional ResNet block architecture, allowing the network to respond to the data’s timesteps. Image $x_{BN}$ is the image processed after the first Batch Normalization Layer.
The integration of the time module directly after batch normalization within the network architecture is a reasonable design choice rooted in the functionality of batch normalization itself. Batch normalization standardizes the inputs to the network layer, stabilizing the learning process by reducing internal covariate shifts. By positioning the time module immediately after this normalization process, the model can introduce time-dependent adaptations to the already stabilized features. This placement ensures that the temporal adjustments are applied to a normalized feature space, thereby enhancing the model’s ability to learn temporal dynamics effectively.
Moreover, the inclusion of the time module at a singular point within the network strikes a balance between model complexity and temporal adaptability. This singular addition avoids the potential redundancy and computational overhead that might arise from multiple time modules. It allows the network to maintain a streamlined architecture while still gaining the necessary capacity to handle time-varying inputs.
Appendix D Results on Classifier Choice
Figure 7: Scatter plots showing the relationship between the number of classes and two performance metrics, AMS (left) and UMS (right). The fitted regression lines demonstrate a positive correlation in both cases.
The analysis of classifier performance across varying numbers of classes reveals interesting patterns for both AMS and UMS. As shown in Figure 7, the scatter plots highlight a positive relationship between the number of classes and the performance metrics. For AMS, the regression line suggests a stronger relationship (R² = 0.637), implying that as the number of classes increases, the AMS metric improves with a moderately strong association. In contrast, the relationship between the number of classes and UMS, while still positive, exhibits a slightly weaker connection (R² = 0.483).
These results suggest that classifier performance, particularly as measured by AMS, benefits more significantly from an increase in the number of classes compared to UMS. The shaded regions in the plots represent the 95% confidence intervals, indicating the range of uncertainty around the fitted regression lines. Overall, the findings imply that the choice of classifier could have a notable impact on AMS , with a less pronounced but still meaningful effect on UMS.
Appendix E Pseudocode for SIDE method
The following Algorithm 1 outlines the detailed steps of the SIDE method for extracting training data from unconditional diffusion models. This algorithm combines the construction of implicit informative labels, the training of a time-dependent classifier, and the conditional generation process to effectively extract valuable training samples.
Algorithm 1 SIDE Method for Extracting Training Data from Unconditional Diffusion Models
1:Input:
  • Unconditional Diffusion Probabilistic Model (DPM) $p_{\theta}(\bm{x})=\mathcal{N}(\bm{\mu},\Sigma_{p_{\theta}})$
  • Pre-trained Classifier $p_{\theta}(y_{I}\mid\bm{x})$
  • Hyperparameter set $\mathcal{S}_{\lambda}$
  • Number of generated samples per $\lambda$, $N_{G}$
  • Target label $y=c$
2:Output: Extracted training data $\mathcal{D}_{extracted}$
3:Construct Implicit Informative Labels
4:Input: Unconditional DPM $p_{\theta}(\bm{x})$, Classifier $p_{\theta}(y_{I}\mid\bm{x})$
5:Define sampling process with implicit labels:
$\mathrm{d}\bm{x}$ $=\left[f(\bm{x},t)-g(t)^{2}\left(\nabla_{\bm{x}}\log p_{\theta}^{t}(\bm{x})+\lambda\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}\mid\bm{x})\right)\right]\mathrm{d}t+g(t)\mathrm{d}\bm{w}$ (151)
6:Calibrate Classifier Output
7:Adjust classifier probabilities using power prior:
$p^{t}_{\theta}\left(\bm{x}\mid y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}\mid\bm{x}\right)p^{t}_{\theta}\left(\bm{x}\right)$ (152)
8:Train Time-Dependent Classifier via TDKD
9:Input: Pre-trained Classifier $p_{\theta}(y_{I}\mid\bm{x})$, Synthetic Dataset $\mathcal{D}_{synthetic}$
10:Output: Time-Dependent Classifier $p_{\theta}^{t}(y_{I}\mid\bm{x}_{t})$
11:Initialize Time-Dependent Classifier architecture with time-dependent modules
12:Generate synthetic dataset using DPM:
$\mathcal{D}_{synthetic}=\{\bm{x}^{(i)}\}_{i=1}^{N_{synthetic}}\sim p_{\theta}(\bm{x})$ (153)
13:Generate pseudo labels using pre-trained classifier:
$y_{I}^{(i)}=p_{\theta}(y_{I}\mid\bm{x}^{(i)})$ (154)
14:Train Time-Dependent Classifier by minimizing KL divergence:
$\mathcal{L}_{distil}=D_{KL}\left(p_{\theta}(y_{I}\mid\bm{x}^{(i)})\,\|\,p_{\theta}^{t}(y_{I}\mid\bm{x}_{t}^{(i)})\right)$ (155)
15:Overall SIDE Procedure
16:Input: Trained Time-Dependent Classifier $p_{\theta}^{t}(y_{I}\mid\bm{x}_{t})$, Target Label $y=c$, Hyperparameter set $\mathcal{S}_{\lambda}$, Number of samples $N_{G}$
17:Initialize empty dataset $\mathcal{D}_{extracted}$
18:for each $\lambda\in\mathcal{S}_{\lambda}$ do
19:     for $i=1$ to $N_{G}$ do
20:         Sample $\bm{x}_{0}$ conditioned on $y=c$ using equation (151)
21:         Compute gradient of cross-entropy loss:
$\nabla_{\bm{x}_{t}}\mathcal{L}_{CE}(c,p_{\theta}^{t}(y\mid\bm{x}_{t}))$ (156)
22:         Reverse diffusion process using computed gradient
23:         Generate similarity score for $\bm{x}_{t}$
24:         Append $\bm{x}_{t}$ to $\mathcal{D}_{extracted}$ with similarity score
25:     end for
26:end for
27:Evaluate Attack Performance
28:Compute evaluation metrics on $\mathcal{D}_{extracted}$
29:return $\mathcal{D}_{extracted}$
Appendix F SSCD-related metrics
In our evaluation of the proposed method, it is essential to include a discussion of the 95th percentile SSCD metric alongside our newly introduced metrics, AMS (Average Matching Similarity) and UMS (Unconditional Matching Similarity). While 95th percentile SSCD metric has its limitations, it still serves as a useful reference point for assessing the relative similarity between extracted images and the original training dataset.
As shown in Table 4, we provide a comprehensive comparison of performance across different datasets.
In addition, we report detailed metrics in the context of our experiments on the CelebA-HQ-FI dataset, which further illustrates the effectiveness of our approach. These results underscore the significance of incorporating SSCD-related scores to complement AMS and UMS in providing a more nuanced understanding of the similarities between generated and training images.
Table 3: Performance Comparison on CelebA-HQ-FI and CelebA-25000 Datasets
Dataset Method Top 0.01% Top 0.05% Top 0.1% Top 0.5% Top 1.0% Top 5.0%
CelebA-HQ-FI Uncond 0.656 0.624 0.604 0.544 0.518 0.463
SIDE(ours) 0.680 0.639 0.618 0.567 0.544 0.485
CelebA-25000 Uncond 0.565 0.538 0.525 0.491 0.475 0.434
SIDE(ours) 0.585 0.554 0.539 0.506 0.491 0.450
Table 4: Generate Training Epoch: 3000 Dataset: CelebA-HQ-FI Generate Nums Per $\lambda$: 50000. The AMS and UMS is measured on Mid Similarity
AMS(%) UMS(%) Top 0.1% Top 0.5% Top 1.0% Top 5.0% Top 10.0%
$\lambda$
0 0.596 0.328 0.604 0.544 0.518 0.463 0.440
1 0.588 0.312 0.596 0.540 0.517 0.463 0.440
2 0.640 0.350 0.591 0.541 0.518 0.465 0.441
3 0.764 0.386 0.594 0.549 0.525 0.470 0.446
4 0.850 0.390 0.604 0.553 0.529 0.473 0.448
5 0.952 0.436 0.596 0.551 0.530 0.476 0.451
6 1.092 0.414 0.611 0.560 0.536 0.480 0.454
7 1.110 0.446 0.607 0.562 0.539 0.482 0.457
8 1.148 0.444 0.618 0.566 0.542 0.484 0.458
9 1.274 0.478 0.615 0.567 0.544 0.485 0.459
10 1.338 0.444 0.613 0.569 0.546 0.487 0.461
11 1.292 0.454 0.604 0.562 0.541 0.486 0.460
12 1.262 0.406 0.617 0.567 0.544 0.486 0.460
13 1.390 0.432 0.617 0.569 0.546 0.489 0.462
14 1.232 0.384 0.613 0.567 0.544 0.485 0.459
15 1.516 0.462 0.616 0.570 0.548 0.490 0.463
16 1.280 0.390 0.612 0.566 0.543 0.487 0.461
17 1.282 0.386 0.605 0.561 0.541 0.486 0.460
18 1.330 0.374 0.616 0.569 0.545 0.488 0.461
19 1.204 0.354 0.612 0.564 0.541 0.485 0.460
20 1.178 0.358 0.603 0.559 0.538 0.483 0.458
21 1.172 0.342 0.617 0.566 0.542 0.484 0.459
22 1.208 0.368 0.602 0.560 0.539 0.485 0.459
23 1.286 0.302 0.607 0.561 0.540 0.485 0.459
24 1.244 0.352 0.597 0.558 0.538 0.484 0.458
25 1.198 0.340 0.599 0.560 0.538 0.483 0.458
26 1.220 0.338 0.601 0.559 0.539 0.483 0.458
27 1.128 0.320 0.608 0.561 0.538 0.483 0.457
28 1.102 0.314 0.604 0.556 0.534 0.481 0.456
29 1.034 0.290 0.595 0.556 0.534 0.481 0.456
30 1.026 0.326 0.602 0.557 0.535 0.480 0.455
31 1.020 0.268 0.591 0.551 0.531 0.479 0.455
32 1.054 0.282 0.593 0.551 0.531 0.479 0.455
33 1.106 0.306 0.600 0.555 0.535 0.481 0.456
34 1.062 0.288 0.582 0.547 0.529 0.479 0.454
35 0.922 0.266 0.587 0.547 0.527 0.477 0.453
36 0.874 0.260 0.585 0.545 0.525 0.477 0.453
37 0.964 0.258 0.589 0.549 0.528 0.477 0.452
38 0.888 0.246 0.582 0.543 0.524 0.475 0.452
39 0.940 0.274 0.587 0.548 0.528 0.476 0.452
40 0.808 0.234 0.587 0.544 0.524 0.474 0.451
41 0.870 0.252 0.582 0.543 0.524 0.476 0.452
42 0.872 0.238 0.584 0.543 0.523 0.475 0.451
43 0.856 0.244 0.584 0.545 0.525 0.475 0.451
44 0.796 0.212 0.578 0.540 0.521 0.473 0.449
45 0.770 0.242 0.580 0.538 0.519 0.472 0.449
46 0.774 0.218 0.580 0.540 0.521 0.472 0.448
47 0.754 0.214 0.581 0.542 0.521 0.471 0.448
48 0.716 0.218 0.572 0.536 0.518 0.471 0.448
49 0.694 0.216 0.570 0.533 0.515 0.469 0.446
50 0.728 0.204 0.576 0.535 0.518 0.471 0.447

### v5

@@
-iclr2025_conference.bib
+appAppendix References
 Abstract
-As diffusion probabilistic models (DPMs) are being employed as mainstream models for Generative Artificial Intelligence (GenAI), the study of their memorization of training data has attracted growing attention.
-Existing works in this direction aim to establish an understanding of whether or to what extent DPMs learn via memorization.
+As diffusion probabilistic models (DPMs) are being employed as mainstream models for Generative Artificial Intelligence (GenAI), the study of their memorization has attracted growing attention.
+Existing works in this field aim to establish an understanding of whether or to what extent DPMs learn via memorization.
 Such an understanding is crucial for identifying potential risks of data leakage and copyright infringement in diffusion models and, more importantly, for trustworthy application of GenAI.
-Existing works revealed that conditional DPMs are more prone to training data memorization than unconditional DPMs, and the motivated data extraction methods are mostly for conditional DPMs.
-However, these understandings are primarily empirical, and extracting training data from unconditional models has been found to be extremely challenging.
-In this work, we provide a theoretical understanding of memorization in both conditional and unconditional DPMs under the assumption of model convergence.
-Our theoretical analysis indicates that extracting data from unconditional models can also be effective by constructing a proper surrogate condition.
-Based on this result, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) that leverages a time-dependent classifier trained on the generated data as a surrogate condition to extract training data from unconditional DPMs.
-Empirical results demonstrate that our SIDE can extract training data in challenging scenarios where previous methods fail, and it is, on average, over 50% more effective across different scales of the CelebA dataset.
+Existing works revealed that conditional DPMs are more prone to memorize training data than unconditional DPMs.
+And most data extraction methods developed so far target conditional DPMs.
+Although unconditional DPMs are less prone to data extraction, further investigation into these attacks remains essential since they serve as the foundation for conditional models like Stable Diffusion, and exploring these attacks will enhance our understanding of memorization in DPMs.
+In this work, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) that leverages a time-dependent classifier trained on generated data as surrogate conditions to extract training data from unconditional DPMs.
+Empirical results demonstrate that it can extract training data in challenging scenarios where previous methods fail, and it is, on average, over 50% more effective across different scales of the CelebA dataset.
+Furthermore, we provide a theoretical understanding of memorization in both conditional and unconditional DPMs and why SIDE is effective.
 1 Introduction
-The diffusion probabilistic models (DPMs) ho2020denoising; sohl2015deep; song2019generative is a family of powerful generative models that learn the distribution of a dataset by first gradually destroying the structure of the data through an iterative forward diffusion process and then restoring the data structure via a reverse diffusion process.
-Due to their outstanding capability in capturing data distribution, DPMs have become the foundation models for many pioneering Generative Artificial Intelligence (GenAI) products such as Stable Diffusion rombach2022high, DALL-E 3 BetkerImprovingIG, and Sora videoworldsimulators2024.
+The diffusion probabilistic models (DPMs) [16, 30, 34] is a family of powerful generative models that learn the distribution of a dataset by first gradually destroying the structure of the data through an iterative forward diffusion process and then restoring the data structure via a reverse diffusion process.
+Due to their outstanding capability in capturing data distribution, DPMs have become the foundation models for many pioneering Generative Artificial Intelligence (GenAI) products such as Stable Diffusion [26], DALL-E 3 [2], and Sora [3].
 Despite the widespread adoption of DPMs, a potential risk they face is data memorization, i.e., the risk of memorizing a certain proportion of the raw training samples.
-This could result in the generation of memorized (rather than new) samples via direct copying, causing data leakage, biased generation, privacy breaches, or copyright infringement, as highlighted in the literature Somepalli2022DiffusionAO; somepalli2023understanding; asay2020independent; cooper2024files; perera2023analyzing.
-A current case argues that Stable Diffusion is a 21st-century collage tool, remixing the copyrighted creations of countless artists whose works were included in the training databutterick2023stable.
+This could result in the generation of memorized (rather than new) samples via direct copying, causing data leakage, privacy breaches or copyright infringement, as highlighted in the literature [31, 32, 1, 11].
+A current case argues that Stable Diffusion is a 21st-century collage tool, remixing the copyrighted creations of countless artists whose works were included in the training data[4].
 Furthermore, data memorization also gives rise to data extraction attacks, which is one type of privacy attacks that attempt to extract the training data from a well-trained model.
-Notably, recent work by carlini2023extracting demonstrated the feasibility of extracting training data samples from DPMs like Stable Diffusion rombach2022high, revealing the potential dangers associated with these models.
+Notably, recent work by [5] demonstrated the feasibility of extracting training data samples from DPMs like Stable Diffusion [26], revealing the potential privacy and copyright dangers associated with them.
 Several works have investigated the data memorization phenomenon in diffusion models.
-For example, it has been observed that there exists a strong correlation between training data memorization and conditional DPMs, i.e., being conditional is more prone to memorization somepalli2023understanding.
-Gu2023OnMI investigates the influential factors on memorization behaviors via a comprehensive set of experiments.
+For example, it has been observed that there exists a strong correlation between training data memorization and conditional DPMs, i.e., conditional-trained DPMs are more prone to memorization [32].
+Gu et al.
+[13] investigates influential factors on memorization behaviors via comprehensive experiments.
 They found that conditioning on random-labeled data can significantly trigger memorization, and unconditional models memorize much less training data.
-Though inspiring, these understandings are primarily empirical.
-Moreover, without a unified theoretical understanding of memorization in both conditional and unconditional DPMs, it is extremely difficult to design an effective data extraction method for unconditional DPMs.
-In this work, we first introduce a memorization metric to quantify the degree of memorization in DPMs by measuring the overlap between the generated and training data in a point-wise manner.
-Based on this metric, we present a theoretical framework that explains why conditional generative models memorize more data and why random labeling can lead to increased memorization.
-Our theoretical analysis indicates that a classifier trained on the same or similar training data can serve as a surrogate condition for unconditional DPMs.
-By further making the classifier time-dependent (to the diffusion sampling process), we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) to extract training data from unconditional DPMs.
-We empirically verify the effectiveness of SIDE on CIFAR-10 and different scales of the CelebA dataset (attack results in Figure 1), confirming the accuracy of the theoretical framework.
+Though these works indicate that unconditional diffusion models are much less prone to data extraction attacks, investigating such attacks for unconditional diffusion models is still necessary.
+This is because 1) unconditional DPMs are the foundations of conditional generative models like Stable Diffusion, which adapts pre-trained unconditional DPMs as the essential part for conditional generation or combines them with conditional training approaches [26, 39]; 2) exploring data extraction attacks on unconditional DPMs enhances understanding of memorization mechanisms in both conditional and unconditional DPMs.
+In this work, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) to extract training data from unconditional DPMs.
+This attack method uses a classifier trained on the same or similar training data as a surrogate condition for unconditional DPMs.
+SIDE can extract training data effectively by making the classifier time-dependent (to the diffusion sampling process).
+We empirically verify the effectiveness of SIDE on CIFAR-10 and different scales of the CelebA dataset (attack results in Figure 1).
+Additionally, to offer a theoretical analysis, we introduce a memorization metric to quantify the degree of memorization in DPMs by measuring the overlap between the generated and training data in a point-wise manner.
+Based on this metric, we explain why conditional DPMs are more likely to memorize training data, why random labelling can increase this memorization and why the SIDE method is effective for extracting training data in unconditional DPMs.
 In summary, our main contributions are:
-  • We introduce a novel metric to measure the degree of point-wise memorization in DPMs and present a theoretical framework that explains 1) why conditional DPMs are more prone to memorization, 2) why random labels can lead to more memorization, and 3) implicit labels (e.g., the learned clusters) can serve as a surrogate condition for unconditional DPMs.
-  • Based on our theoretical understanding, we propose a novel training data extraction method SIDE that leverages the implicit labels learned by a time-dependent classifier to extract training data from unconditional DPMs.
-  • We evaluate the effectiveness of SIDE on CIFAR-10 and various scales of CelebA datasets, and show that, on average, it can outperform the baseline method proposed by carlini2023extracting by more than 50%.
+  • We propose a novel training data extraction method SIDE that leverages the implicit labels learned by a time-dependent classifier to extract training data from unconditional DPMs.
+  • We introduce a novel metric to measure the degree of point-wise memorization in DPMs and present a theoretical analysis that explains 1) why conditional DPMs are more prone to memorization, 2) why random labels can lead to more memorization, and 3) why the SIDE method is effective for extracting training data in unconditional DPMs.
+  • We evaluate the effectiveness of SIDE on CIFAR-10 and various scales of CelebA datasets and show that, on average, it can outperform the baseline method by more than 50%.
 Figure 1: A few examples of the extracted images from a DDPM trained on a subset of the CelebA dataset using our SIDE method. Top: training images; bottom: extracted images.
 2 Related Work
-Diffusion Probabilistic Models  DPMs sohl2015deep such as Stable Diffusion rombach2022high, DALL-E 3 BetkerImprovingIG, Sora videoworldsimulators2024, Runway rombach2022high, and Imagen saharia2022photorealistic have achieved state-of-the-art performance in image/video generation across a wide range of benchmarks dhariwal2021diffusion.
+Diffusion Probabilistic Models  DPMs [30] such as Stable Diffusion [26], DALL-E 3 [2], Sora [3], Runway [26], and Imagen [28] have achieved state-of-the-art performance in image/video generation across a wide range of benchmarks [12].
 These models can be viewed from two perspectives.
-The first is score matching song2019generative, where diffusion models learn the gradient of the image distribution song2020score.
-The second perspective involves denoising DPMs ho2020denoising, which add Gaussian noise at various time steps to clean images and train models to denoise them.
-To conditionally sample from diffusion models, dhariwal2021diffusion utilizes a classifier to guide the denoising process at each sampling step.
-Additionally, ho2022classifier introduces classifier-free guidance for conditional data sampling using DPMs.
-Memorization in Diffusion Models  Early exploration of memorization in large models was focused on language models carlini2022quantifying; jagielski2022measuring, which has motivated more in-depth research on DPMs somepalli2023understanding; Gu2023OnMI; rahman2024frame.
-Notably, a recent research on image-generation DPMs: Somepalli2022DiffusionAO found that 0.5-2% of generated images duplicate training samples, a result concurrently reported by carlini2023extracting in broader experiments for both conditional and unconditional diffusion models.
-Further studies somepalli2023understanding and Gu2023OnMI linked memorization to model conditioning, showing that conditional models are more prone to memorization.
+The first is score matching [34], where diffusion models learn the gradient of the image distribution [35].
+The second perspective involves denoising DPMs [16], which add Gaussian noise at various time steps to clean images and train models to denoise them.
+To conditionally sample from diffusion models, [12] utilizes a classifier to guide the denoising process at each sampling step.
+Additionally, [17] introduces classifier-free guidance for conditional data sampling using DPMs.
+Memorization in Diffusion Models  Early exploration of memorization in large models was focused on language models [6, 19], which has motivated more in-depth research on DPMs [32, 13, 24].
+Notably, a recent research on image-generation DPMs: [31] found that 0.5-2% of generated images duplicate training samples, a result concurrently reported by [5] in broader experiments for both conditional and unconditional diffusion models.
+Further studies [32] and [13] linked memorization to model conditioning, showing that conditional models are more prone to memorization.
 Several works have proposed methods to detect and mitigate memorization to address this issue.
 Wen et al.
-wen2024detecting proposed a method to detect whether a prompt can trigger memorization by inspecting the magnitude of text-conditional predictions.
+[38] proposed a method to detect whether a prompt can trigger memorization by inspecting the magnitude of text-conditional predictions.
 It delivers high accuracy even at the first generation step, with a single generation per prompt to detect whether a prompt can trigger memorization.
 Ren et al.
-ren2024unveiling proposed two metrics based on findings on cross-attention in DPMs to detect memorization.
-To mitigate memorization, Chen et al.chen2024towards proposed anti-memorization guidance to mitigate memorization during the sampling process.
+[25] proposed two metrics based on findings on cross-attention in DPMs to detect memorization.
+To mitigate memorization, Chen et al.[7] proposed anti-memorization guidance to mitigate memorization during the sampling process.
 Instead of using guidance, Ren et al.
-ren2024unveiling mitigate memorization by changing the attention score or masking out summary tokens in the cross-attention layer.
-Wen et al.wen2024detecting do so by minimizing the magnitude during inference.
-So far, the motivated data extraction attack or defend methods from these empirical understandings are mostly for conditional DPMs carlini2023extracting; webster2023reproducible, and studies have shown that extracting training data from unconditional DPMs can be much more challenging than that on conditional DPMs Gu2023OnMI.
-Although the empirical understandings are inspiring, a theoretical understanding of the memorization behaviors of DPMs is missing from the current literature.
-A recent attempt by ross2024a uses the local intrinsic dimension (LID) metric to characterize memorization in DPMs.
-However, the finding that lower LID leads to more memorization is only verified on a toy 1-dimensional dataset and only a few generated images, and it fails to address the data extraction challenge on unconditional DPMs.
-This challenge is exemplified by the brute-force methods currently employed, as discussed in Somepalli2022DiffusionAO; carlini2023extracting, highlighting the inadequacies of existing approaches for extracting data from unconditional diffusion models.
-In response to these challenges, we introduce a theoretical framework to characterize memorization in DPMs, which further motivates a novel data extraction method for unconditional DPMs.
-3 Proposed Theory
-In this section, we introduce a memorization metric and provide a theoretical explanation for the universality of data memorization in both conditional and unconditional diffusion models.
-3.1 Memorization Metric
-Intuitively, the memorization of fixed training data points (i.e., pointwise memorization) can be quantified by the degree of overlap between the generated distribution and the distributions centered at each training data point.
-Given a generative model and training dataset, we propose the following memorization metric to quantify the memorization of the training data points in the model.
-Definition 1 (Pointwise Memorization)
-Given a generative model $p_{\theta}$ with parameters $\theta$ and training dataset ${\mathcal{D}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, the degree of memorization in $p_{\theta}$ of ${\mathcal{D}}$ is defined as:
-${\mathcal{M}}_{point}({\mathcal{D}};\theta)=\sum_{{\bm{x}}_{i}\in{\mathcal{D}}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}},$ (1)
-where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ is the $i$-th training sample, $N$ is the total number of training samples, $p_{\theta}({\bm{x}})$ represents the probability density function (PDF) of the generated samples, and $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the probability distribution centered at training sample ${\bm{x}}_{i}$ within a small radius $\epsilon$.
-A straightforward choice for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the Dirac delta function centered at training data point ${\bm{x}}_{i}$: $q({\bm{x}},{\bm{x}}_{i},\epsilon)=\delta({\bm{x}}-{\bm{x}}_{i})$.
-However, this would make equation (1) uncomputable as the Dirac delta function is zero beyond the $\epsilon$-neighborhood.
-Alternatively, we could use the Gaussian distribution with a covariance matrix $\epsilon\textit{I}$ (I is the identity matrix) for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$:
-$q({\bm{x}},{\bm{x}}_{i},\epsilon)=\frac{1}{\sqrt{(2\pi\epsilon)^{d}}}\exp\left\{-\frac{1}{2\epsilon}({\bm{x}}-{\bm{x}}_{i})^{\top}({\bm{x}}-{\bm{x}}_{i})\right\}.$ (2)
-Note that in equation (1), a smaller value of ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ indicates more overlap between the two distributions and thus more memorization.
-As $\epsilon$ tends to 00, the measured memorization becomes more accurate, with the limit characterizing the intrinsic memorization capability of model $p_{\theta}$.
-Semantic Memorization
-As a metric, ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ should be monotonous with the actual semantic memorization effect, which refers to the model’s tendency to reproduce unique features of the training samples rather than generating near-duplicate examples.
-Here, we define semantic memorization in the latent space as it distills the unique semantic information of each training sample.
-Definition 2 (Semantic Memorization)
-Let ${\mathcal{D}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$ be the dataset, $p_{\theta}(x)$ be the PDF of the generated samples by model $p_{\theta}$, and $\bm{z}$ be the learned latent code for data sample $\bm{x}$.
-The semantic memorization of model $p_{\theta}$ on ${\mathcal{D}}$ is:
-[…]
-where $\bm{z_{i}}$ denotes the ground truth latent code of $\bm{x_{i}}$.
-${\mathcal{M}}_{semantic}\left({\mathcal{D}};\theta\right)$ measures semantic memorization because it evaluates how well the learned representations ${\bm{z}}$ of training samples align with their ground truth latent codes, capturing the underlying structure of the data rather than simply memorizing the exact duplicates of the training data points.
-In practice, the ground truth latent codes are known but can be approximated by an independent encoder.
-The following theorem formulates the relationship between pointwise and semantic memorization.
-Theorem 1
-Pointwise memorization ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ is monotonic to semantic memorization, formally:
-$\frac{\partial{\mathcal{M}}_{point}({\mathcal{D}};\theta)}{\partial{\mathcal{M}}_{semantic}({\mathcal{D}};\theta)}=\frac{1}{2\epsilon}+\frac{\operatorname{Tr}(\Sigma_{p_{\theta}}^{-1})}{2}>0$ (4)
-where $\Sigma_{p_{\theta}}$ is the covariance matrix of the learned latent distribution of the training data.
-We first build the relationship map ${\mathcal{M}}_{point}$ into latent space, and then break down equation (4) into separate components and derive each component separately.
-So, the equation (4) can be simplified to $\frac{1}{2\epsilon}$.
-The detailed proof can be found in Appendix A.2.
-Note that the memorization effect studied in this work refers specifically to the pointwise memorization ${\mathcal{M}}_{point}$.
-3.2 Theoretical Framework
-Based on pointwise memorization, here we present a theoretically framework that explains why conditional DPMs memorize more data.
-Our theoretical framework is based on the concept of informative labels, which refers to information that can differentiate subsets of data samples.
-We first give a formal definition of informative labels and then prove their two key properties: 1) they facilitate tighter clustering of samples around their respective means, and 2) they reduce variance in the latent representations.
-Building upon the two properties, we theoretically show that conditional DPMs memorize more data.
-Informative Labels
-The concept of informative labels has been previously discussed as class labels Gu2023OnMI.
-In this work, we introduce a more general definition that encompasses both class labels and random labels as special cases.
-We define an informative label as follows.
-Definition 3 (Informative Label)
-Let ${\mathcal{Y}}=\{y_{i},y_{2},\cdots,y_{C}\}$ be the label set for training dataset ${\mathcal{D}}$ with $C$ unique labels.
-$y_{i}$ is the associated label with ${\bm{x}}_{i}$, and ${\mathcal{D}}_{y=c}=\{{\bm{x}}_{i}:{\bm{x}}_{i}\in\mathcal{D},y_{i}=c\}$ is the subset of training samples shared the same label $y=c$.
-A label $y=c$ is an informative label if
-$|{\mathcal{D}}_{y=c}|<|{\mathcal{D}}|.$ (5)
-Here, the labels are not limited to the conventional class labels; they can also be text captions, features, or cluster information that can be used to group the training samples into subsets.
-The above definition states that an information label should be able to differentiate a subset of samples from others.
-An extreme case is that all samples have the same label; in this case, the label is not informative.
-According to our definition, class-wise and random labels are special cases of informative labels.
-Informative labels can be either explicit like class/random labels and text captions, or implicit like silent features or clusters.
-Next, we will explore the correlation between informative labels and the clustering effect in the latent space of a generative model with an encoder and decoder.
-In diffusion models, the encoder represents the forward diffusion process, while the decoder represents the reverse process.
-Notably, the latent space can be defined at earlier timesteps according to ho2020denoising.
-Memorization in Conditional DPMs
-Informative labels cause a clustering effect in the latent space of DPMs by providing contextual information that allows the encoder to better differentiate between data samples.
-When data points are associated with informative labels, the encoder can map them to a latent distribution that accurately reflects their shared characteristics.
-This results in tighter clusters in the latent space as samples with the same informative label become more concentrated around their respective means.
-Consequently, the latent representations of these samples exhibit reduced variance, leading to a more structured and organized latent space.
-We prove these two properties under the assumption of model convergence.
-Suppose we have an encoder $f_{\theta_{E}}({\bm{x}})$ and a decoder $f_{\theta_{D}}({\bm{z}})$.
-The encoder $f_{\theta_{E}}({\bm{x}})$ maps data samples ${\bm{x}}\in{\mathcal{D}}$ to the latent distribution ${\bm{z}}$, which follows a normal distribution ${\mathcal{N}}(\mathbf{{\bm{\mu}}},\mathbf{\Sigma})$: $p_{\theta}({\bm{z}})={\mathcal{N}}(\mathbf{{\bm{\mu}}},\mathbf{\Sigma})$.
-For ${\bm{x}}_{i}\in{\mathcal{D}}_{y=c}$, the encoder maps ${\bm{x}}_{i}$ to a latent distribution ${\bm{z}}_{c}$ subject to ${\mathcal{N}}(\mathbf{{\bm{\mu}}_{c}},\mathbf{\Sigma_{c}})$.
-The decoder $f_{\theta_{D}}({\bm{z}})$ maps ${\bm{z}}$ back to the original data samples ${\bm{x}}$.
-$y_{i}$ is the label of data sample ${\bm{x}}_{i}$.
-Training a generative $p_{\theta}$ is to optimize the following:
-$\min_{\theta}-\sum_{{{\bm{x}}}_{\mathrm{i}}\in\mathrm{{\mathcal{D}}}}{\log{p}_{\theta}\left({{\bm{x}}_{i}}|y_{i}\right)}.$ (6)
-Assumption 1
-Given suffix training on ${\mathcal{D}}$, the generative model $p_{\theta}$ converges to an optimal solution for equation (6) : $\theta^{*}=\arg\min_{\theta}-\sum_{{\bm{x}}_{\mathrm{i}}\in\mathcal{D}}\log p_{\theta}({\bm{x}}_{i}|y_{i})$.
-Based on the above assumption, we can derive the following Proposition 1, with detailed proof deferred to Appendix A.3
-Proposition 1
-Let ${\bm{z}}$ be the latent space of generative model $p_{\theta}$ conditioned on an informative label $y=c$, the latent representations learned by $p_{\theta}$ under Assumption 1 satisfy:
-$\sum_{{\bm{z}}_{i}\in\mathcal{D}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})\leq\sum_{{\bm{z}}_{i}\in\mathcal{D}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}),$ (7)
-$\|\mathbf{\Sigma}_{c}\|_{*}\leq\|\mathbf{\Sigma}\|_{*},$ (8)
-where $\Sigma$ is the covariance matrix, $\Sigma_{c}$ is the covariance matrix conditioned on informative label $y=c$, $\bm{\mu}_{c}$ denotes the mean of the latent representations ${\bm{z}}$ conditioned on $y=c$ and $\bm{\mu}$ denotes the mean on the overall dataset ${\mathcal{D}}$.
-Proposition 1 describes two properties of the learned latent space driven by informative labels: 1) tighter clustering as defined in equation (7) and 2) reduced variance as defined in equation (8).
-Tighter clustering means that the data samples associated with the same informative label are more closely clustered, which allows the model to more effectively capture and memorize the specific features and patterns relevant to those labels.
-This proximity in the latent space enhances the model’s ability to recall memorized samples during generation, as the representations are organized around distinct means.
-Additionally, reduced variance in these clustered representations leads to greater stability, ensuring that the model can consistently reproduce memorized outputs from the clustered latent codes.
-Based on this understanding, we formalize the relationship between informative labels and memorization in conditional DPMs via the following theorem.
-Theorem 2
-A generative model $p_{\theta}$ incurs a higher degree of pointwise memorization when conditioned on informative labels $y=c$, mathematically expressed as:
-$\lim_{\epsilon\to 0}\frac{{\mathcal{M}}_{point}({\mathcal{D}}_{y=c},\theta_{y=c})}{{\mathcal{M}}_{point}({\mathcal{D}}_{y=c},\theta)}\leq 1$ (9)
-where $\theta_{y=c}$ denotes the parameters of the model when trained on dataset ${\mathcal{D}}_{y=c}$.
-The proof can be found in Appendix A.4.
-Theorem 2 states that any form of information labels can incur more memorization in DPMs, including conventional class labels and random labels.
-As shown in our proof, the informative nature of a label reduces the entropy (or uncertainty) of the data distribution conditioned on that label, leading to a more focused and memorable data representation.
-This explains the two empirical observations made in existing works Gu2023OnMI: 1) conditional DPMs are more prone to memorization and 2) even random labels can lead to more memorization.
-It also explains the findings that unconditional models do not replicate data and that text conditioning increases memorization Somepalli2022DiffusionAO; chen2024towards.
-Memorization in Unconditional DPMs
-According to Theorem 2, one could leverage informative labels to extract training data from conditional DPMs.
-Intuitively, the Text captions and class labels commonly used to train conditional DPMs are valid informative labels, which we call explicit informative labels.
-However, unconditional DPMs do not have explicit informative labels and thus are more difficult to extract training data from.
-Nevertheless, our theory indicates that the learned representation clusters by an unconditional DPM can also serve as a type of informative labels, which we call implict information labels.
-It means that if we can formulate the clustering information in the training data, we could construct implicit informative labels to help extract training data from unconditional DPMs.
-This motivates us to propose a new data extraction method SIDE for unconditional DPMs in the next section.
+[25] mitigated memorization by changing the attention score or masking out summary tokens in the cross-attention layer.
+Wen et al.[38] achieved this by minimizing the magnitude during inference.
+So far, the motivated data extraction attack or defend methods are primarily for conditional DPMs [5, 37], and studies have shown that extracting training data from unconditional DPMs can be much more challenging than that on conditional DPMs [13].
+Though challenging, investigating data extraction methods in unconditional DPMs is still necessary.
+This is because 1) Unconditional DPMs are key parts in conditional models like Stable Diffusion, which adapt them for conditional generation, and 2) Exploring this approach enhances our understanding of memorization in DPMs.
+In response to these challenges, we propose a novel data extraction method called SIDE for unconditional DPMs and providing theoretical understanding of SIDE.
 3 Proposed SIDE Method
-In this section, we will construct implicit informative labels for unconditional DPMs, convert the implicit labels into explicit ones, and then leverage the explicit labels to extract training data.
+3.1 Intuition
+Previous studies [13, 32] have shown that conditional DPMs are more prone to memorization due to their reliance on explicit conditioning labels, such as class labels or text conditions, which we refer to as explicit informative labels or explicit labels.
+In contrast, unconditional DPMs do not have access to these explicit labels.
+Nevertheless, they still learn to form representation clusters from the training data, even though these clusters are not directly specified in the data itself [8].
+We name these representation clusters as implicit informative labels or implicit labels.
+By formalizing the clustering patterns within the training data, it may be possible to construct implicit labels and make the unconditional DPMs conditional on these implicit labels to extract training data from unconditional DPMs.
+In the following section, we will construct implicit informative labels for unconditional DPMs, convert the implicit labels into explicit ones, and then leverage the explicit labels to extract training data.
 3.2 Constructing Implicit Informative Labels
@@
 The classifier can be a normal classifier trained on the same dataset as the diffusion model.
-When such a classifier is not available, random labels or representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can also be used as the implicit labels, according to our theoretical analysis in Section 3.2.
-We assume an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process defined as:
-$\mathrm{d}{\bm{x}}=\left[f({\bm{x}},t)-g(t)^{2}\left(\nabla_{\bm{x}}\log p_{\theta}^{t}({\bm{x}})+\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\right)\right]\mathrm{d}t+g(t)\mathrm{d}w,$ (10)
+When such a classifier is not available, random labels or representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can also be used as the implicit labels.
+We assume an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process:
+$\mathrm{d}{\bm{x}}=$ $\Big{[}f({\bm{x}},t)-g(t)^{2}\Big{(}\nabla_{\bm{x}}\log p_{\theta}^{t}({\bm{x}})+$
+$\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\Big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w$ (1)
 where $x$ represents the state vector, $f(x,t)$ denotes the drift coefficient, $g(t)$ is the diffusion coefficient, $\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$ denotes the gradient of model $p_{\theta}$ given ${\bm{x}}$ at time $t$, and $\mathrm{d}w$ corresponds to the increment of the Wiener process.
-We can use the classifier that generates the implicit labels to approximate the gradient in equation (10).
-However, a known challenge associated with neural network classifiers is their tendency towards miscalibration guo2017calibration.
+We can use the classifier that generates the implicit labels to approximate the gradient in Eq.
+(3.2).
+However, a known challenge associated with neural network classifiers is their tendency towards miscalibration [14].
 Specifically, the classifier could be overconfident or underconfident about its predictions.
 To mitigate the potential impact of miscalibration on the sampling process, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output on the diffusion path using power prior as follows:
-$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right).$ (11)
+$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right)$ (2)
 Then, we have:
-$\mathrm{d}{\bm{x}}=\left[f({\bm{x}},t)-g(t)^{2}\left(\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\right)\right]\mathrm{d}t+g(t)\mathrm{d}w.$ (12)
+$\mathrm{d}{\bm{x}}=$ $\Big{[}f({\bm{x}},t)-g(t)^{2}\big{(}\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$
+$+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w$ (3)
 […]
@@
 3.3 Time-dependent Classifier
-In equation (151), the classifier is denoted by $p_{\theta}^{t}(y|{\bm{x}})$ with $t$ being the timestep, implying its time-dependent nature.
+In Eq.
+(3.2), the classifier is denoted by $p_{\theta}^{t}(y|{\bm{x}})$ with $t$ being the timestep, implying its time-dependent nature.
 However, we do not have a time-dependent classifier but only a time-independent classifier.
@@
 TDKD equips the classifier with time-dependent guidance during sampling.
-It operates in two steps: first, the network architecture is adjusted to accommodate time-dependent inputs, and the structure of the time-dependent module and modification are illustrated in Appendix C; second, a synthetic dataset and pseudo labels are created to facilitate knowledge distillation from the normal classifier to its time-dependent counterpart.
-As the original training dataset is unknown, we employ the target DPM to generate a synthetic dataset, following the generative data augmentation techniques chen2023data; chen2024comprehensive.
+It operates in two steps: first, the network architecture is adjusted to accommodate time-dependent inputs, and the structure of the time-dependent module and modification are illustrated in Appendix C; second, a synthetic dataset and pseudo labels are created to distil knowledge from the normal classifier to its time-dependent counterpart.
+As the original training dataset is unknown, we employ the target DPM to generate a synthetic dataset, following the generative data augmentation techniques [10, 9].
 We then use the normal classifier trained on the original dataset to generate pseudo labels for the synthetic dataset.
 Finally, we train a time-dependent classifier on the labelled synthetic dataset.
-The objective of this training is defined as following:
-$\mathcal{L}_{distil}=D_{KL}\left(p_{\theta}(y_{I}|{\bm{x}}),p_{\theta}^{t}(y_{I}|{\bm{x}}_{t})\right).$ (13)
+The objective of this training is defined as follows:
+$\mathcal{L}_{distill}=D_{KL}\left(p_{\theta}(y_{I}|{\bm{x}}),p_{\theta}^{t}(y_{I}|{\bm{x}}_{t})\right)$ (4)
 Overall Procedure of SIDE
@@
 Assume we condition on the label $y=c$.
-First, we pick a set of values for $\lambda$, called $\mathcal{S}_{\lambda}$, to use in the SIDE attack.
+Firstly, we pick a set of values for $\lambda$, called $\mathcal{S}_{\lambda}$, to use in the SIDE attack.
 Then, we gather $N_{G}$ data samples for each value of $\lambda$ from the set.
-During each sampling timestep $t$, we compute the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ ($CE(\cdot)$ is the cross-entropy loss), then we use the gradient and the target DPM to reverse the diffusion process.
-Third, we compute the similarity score for each generated image.
-Lastly, we evaluate the attack performance using evaluation metrics.
+For each sampling timestep $t$, we compute the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ ($CE(\cdot)$ denotes the cross-entropy loss) and use the gradient and the target DPM to reverse the diffusion process.
+Next, we compute the similarity score for each generated image.
+Finally, we evaluate the attack performance using evaluation metrics.
 Defer to Appendix E for SIDE’s pseudocode.
 4 Experiments
-In this section, we first describe our experimental setup, introduce the performance metrics, and then present the main evaluation results of our SIDE method.
+In this section, we first describe our experimental setup, introduce the performance metrics and then present the main evaluation results of our SIDE method.
 We empirically verify that the memorized images are not from the classifier.
@@
 4.1 Experimental Setup
-We evaluate the effectiveness of our method on three datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) na2022unrestricted which consists of 5478 images, 2) a subset of the CelebA (CelebA-25000) liu2015faceattributes which contains 25,000 images and 2) CIFAR-10 containing 50,000 images.
+We evaluate the effectiveness of our method on three datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) [23] which consists of 5478 images, 2) a subset of the CelebA (CelebA-25000) [21] which contains 25,000 images and 2) CIFAR-10 containing 50,000 images.
 All the images are normalized to [-1,1].
-We use the AdamW optimizer loshchilov2018decoupled with a learning rate of 1e-4 to train the time-dependent classifier.
-We train denoising DPMs with a discrete denoising scheduler (DDIM song2021denoising) using the HuggingFace implementation diffusers.
+We use the AdamW optimizer [22] with a learning rate of 1e-4 to train the time-dependent classifier.
+We train denoising DPMs with a discrete denoising scheduler (DDIM [33]) using the HuggingFace implementation [36].
 All DPMs are trained with batch size 64.
 We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, respectively.
-We use ResNet34 He2015DeepRL as the normal classifier.
+We use ResNet34 [15] as the normal classifier.
 4.2 Performance Metrics
 Determining whether a generated image is a memorized copy of a training image is difficult, as $L_{p}$ distances in the pixel space are ineffective.
-Previous research addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection Somepalli2022DiffusionAO; Gu2023OnMI.
+Previous research addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection [31, 13].
 However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization.
-Here, we propose two new memorization scores to solve these issues: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS metrics).
+Here, we propose two new memorization scores to solve these issues: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS).
 The AMS and UMS metrics are defined as follows:
-$AMS\left({{\mathcal{D}}}_{{\text{gen}}},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{\text{gen}}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)}}{N_{G}},$ (14)
-$UMS\left({{\mathcal{D}}}_{{\text{gen}}},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{\text{gen}}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)|}{N_{G}},$ (15)
-where ${{\mathcal{D}}}_{\text{gen}}$ is the generated dataset, ${\mathcal{D}}_{\text{train}}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring.
-$\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria.
-$\phi({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria.
-Mathematically, we can represent $\mathcal{F}(x_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)=\mathbb{1}[\max_{x_{j}\in{\mathcal{D}}_{\text{train}}}\gamma(x_{i},x_{j})\geq\alpha\;\&\;\gamma(x_{i},x_{j})\leq\beta]$ $\varphi(x_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)=\{j:x_{j}\in{\mathcal{D}}_{\text{train}},\gamma(x_{i},x_{j})\geq\alpha\;\&\;\gamma(x_{i},x_{j})\leq\beta\}$.
-$\gamma$ represents the similarity/distance function.
-For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$ following carlini2023extracting, while for high-resolution datasets, we use the SSCD score as $\gamma$.
+$AMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{1},\alpha,\beta\right)}}{N_{G}}$ (5)
+$UMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)|}{N_{G}}$ (6)
+where ${{\mathcal{D}}}_{1}$ is the generated dataset, ${\mathcal{D}}_{2}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring.
+$\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria.
+$\phi({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria.
+Mathematically, we can represent:
+$\mathcal{F}(x_{i},{\mathcal{D}}_{2},\alpha,\beta)=\mathds{1}\Big{[}\max_{x_{j}\in{\mathcal{D}}_{2}}\gamma(x_{i},x_{j})\in[\alpha,\beta]\Big{]}$ (7)
+$\phi(x_{i},{\mathcal{D}}_{2},\alpha,\beta)=\{j:x_{j}\in{\mathcal{D}}_{2},\gamma(x_{i},x_{j})\in[\alpha,\beta]\}$ (8)
+where $\gamma$ represents the similarity/distance function.
+For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$ following [5], while for high-resolution datasets, we use the SSCD score as $\gamma$.
 In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity.
-The thresholds for the normalized $L_{2}$ ditance carlini2023extracting are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
+(examples of different thresholds are shown in Figure 3) The thresholds for the normalized $L_{2}$ ditance [5] are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
 The AMS averages the similarity scores across generated images, ensuring that memorized images are not overlooked.
@@
 By further categorizing the two scores into three levels—low, mid, and high—we obtain more comprehensive measurements for different types of memorization.
-While carlini2023extracting; chen2024towards introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is essential for assessing different types of copyright infringement, such as character or style copying lee2023talkin; sag2023copyright; sobel2023elements.
-Additionally, the UMS considers the number of generated images $N_{G}$, which was overlooked in carlini2023extracting.
+While [5, 7] introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is essential for assessing different types of copyright infringement, such as character or style copying [20, 27, 29].
+Additionally, the UMS considers the number of generated images $N_{G}$, which was overlooked in [5].
 The significance of $N_{G}$ lies in its non-linear impact on the UMS (see Appendix B for proof), indicating that UMS scores should not be compared across different values of $N_{G}$.
@@
 This classification highlights varying degrees of semantic resemblance among the image pairs.
-5.3 Main Results
-We compare SIDE with a random baseline and a variant of SIDE that substitutes the time-dependent classifier with a standard (time-independent) classifier.
-Here, "TD" refers to the time-dependent classifier trained using our proposed TDKD method, "TI" denotes the time-independent classifier, and "OL" indicates training with the original dataset labels.
-The "Random" baseline generates images directly using the target unconditional DPM, as described in carlini2023extracting.
-We average the results across various values of $\lambda$ (defined in equation (151)), ranging from 5 to 9, with a detailed analysis provided in Section 5.3.
-It is important to note that $\lambda=0$ corresponds to the "Random" baseline.
-For each $\lambda$, including $\lambda=0$, we generate 50,000 images for CelebA and 10,000 images for CIFAR-10 to validate our theoretical analysis and the proposed SIDE method.
-This effort results in one of the largest generated image datasets to date for studying the memorization of DPMs.
-Table 1: The AMS (%) and UMS (%) results at low (top), mid (middle), and high (bottom) levels.
-‘Random’ denotes the baseline that generates images directly using the target unconditional DPM, while OL-TI is a variant of SIDE which is not trained using TDKD.
+Table 1: The AMS (%) and UMS (%) results at low, mid, and high levels.
+‘Random’ denotes the baseline that generates images directly using the target unconditional DPM[5], while OL-TI is a variant of SIDE which is not trained using TDKD.
 Dataset Method Low Similarity Mid Similarity High Similarity
@@
 SIDE (Ours) 5.325 2.053 2.495 0.860 1.770 0.560
+4.3 Main Results
+We compare SIDE with the Random baseline proposed by Carlini et al.[5] and a variant of SIDE that substitutes the time-dependent classifier with a standard (time-independent) classifier.
+Here, ”TD” refers to the time-dependent classifier trained using our proposed TDKD method, ”TI” denotes the time-independent classifier, and ”OL” indicates training with original dataset labels.
+The ”Random” baseline generates images directly using the target unconditional DPM, as described in [5].
+We average the results across various values of $\lambda$ (defined in Eq.
+(3.2)), ranging from 5 to 9, with a detailed analysis provided in Section 4.3.
+It is important to note that $\lambda=0$ corresponds to the ”Random” baseline.
+For each $\lambda$, including $\lambda=0$, we generate 50,000 images for CelebA and 10,000 images for CIFAR-10 to validate the proposed SIDE method.
+This effort results in one of the most enormous generated image datasets for studying the memorization of DPMs.
+Table 2: This table presents the results of fitting a linear model to the relationship between the number of classes and both AMS and UMS.
+Relationship Coefficient (×$\times 10^{-5}$) Intercept $R^{2}$ Correlation Coefficient
+#Class vs. AMS 7.4 (positive) 0.115 0.637 0.80
+#Class and UMS 6.1 (positive) 0.090 0.483 0.70
+Figure 4: Hyper-parameter ($\lambda$) analysis on CelebA-HQ-FI.
+For high similarity, the best $\lambda$ for AMS and UMS are 16 and 13.
+For other similarity levels, the best $\lambda$ for AMS and UMS is 13.
 Effectiveness of SIDE
-The AMS and UMS results for the three datasets are presented in Table 1.
+The AMS and UMS results for the three datasets are presented in Table 1 (results of 95-th percentile of SSCD are listed in Appendix F).
 As shown, SIDE is highly effective in extracting memorized data across all similarity levels, particularly on the CelebA-25000 dataset, a task previously deemed unfeasible due to its large scale.
@@
 Effectiveness of TDKD
-As shown in Table  1, time-independent classifiers perform significantly worse than their time-dependent counterparts on the two high-resolution CelebA datasets, achieving only about 10% of the effectiveness of classifiers trained using our TDKD method.
+As shown in Table 1, time-independent classifiers perform significantly worse than their time-dependent counterparts on the two high-resolution CelebA datasets, achieving only about 10% of the effectiveness of classifiers trained using our TDKD method.
 This discrepancy arises because time-independent classifiers can provide accurate gradients only at the final timestep, whereas time-dependent classifiers deliver accurate gradients at each timestep.
@@
 There might be a concern that the extracted images may originate from the classifier rather than the target DPM.
-We argue that even if a time-dependent classifier could disclose its training images, these images would still originate from the target DPM, as they can also be extracted with the time-dependent classifiers.
-Furthermore, Table  1 demonstrates that employing a simple classifier reduces memorization compared to the baseline extraction method, indicating that the memorized images do not originate from the classifier.
-Figure 4: Hyper-parameter ($\lambda$) analysis on CelebA-HQ-FI.
-For high similarity, the best $\lambda$ for AMS and UMS are 16 and 13.
-For other similarity levels, the best $\lambda$ for AMS and UMS is 13.
+We argue that even if a time-dependent classifier could disclose its training images, these images would still originate from the target DPM because the classifier is distilled on the synthetic dataset generated by the target DPM.
+Furthermore, Table  1 demonstrates that employing a simple classifier reduces memorization compared to the baseline extraction method, indicating that the memorized images do not originate from the simple classifier as well.
 Impact of the Classifier on SIDE
@@
 Here, we conduct experiments to explore the relationship between the number of classes (of the classifier) and the extraction performance at a low similarity level, using 1,200 images per class with $\lambda=5$ on the CelebA-HQ-FI dataset.
-As shown in Table  2 and Figure 7 (Appendix), there exists a strong positive correlation: as the number of classes increases, both AMS and UMS improve.
+As shown in Table  2 and Figure 5, there exists a strong positive correlation: as the number of classes increases, both AMS and UMS improve.
 In summary, increasing number of classes positively affects both AMS and UMS, with a stronger impact on AMS.
 The UMS values reported here differ significantly from Table  1.
-This is because here, we only generated 1,200 images per class, whereas 50,000 images per class in the previous experiment.
+This is because we only generated 1,200 images per class here, whereas 50,000 images per class in the previous experiment.
 Increasing the number of classes improves AMS more than UMS because a higher class count enables the classifier to better differentiate fine details, leading to more accurate matches at low similarity levels.
 UMS is less affected since it relies more on the diversity across images, which is constrained by the smaller number of images per class in this experiment.
-Table 2: This table presents the results of fitting a linear model to the relationship between the number of classes and both AMS and UMS.
-Relationship Coefficient (×$\times 10^{-5}$) Intercept $R^{2}$ Correlation Coefficient
-Class and AMS 7.4 0.115 0.637 0.80
-Class and UMS 6.1 0.090 0.483 0.70
+Figure 5: Scatter plots showing the relationship between the number of classes and two performance metrics, AMS (left) and UMS (right).
+The fitted regression lines demonstrate a positive correlation in both cases.
 Hyper-parameter Analysis
@@
 As shown in Figure 4, the memorization score increases at first, reaching its highest, then decreases as $\lambda$ increases.
-This can be understood from sampling SDE Equation (151).
+This can be understood from sampling SDE Eq.
+(3.2).
 Starting from 0, the diffusion models are unconditional.
@@
 Consequently, the memorization score decreases.
+5 Theoretical Understanding for Memorization in Conditional Generative Models
+In this section, we introduce a metric to quantify the memorization capacity of generative models and provide a theoretical explanation for the data memorization in conditional diffusion models and why SIDE is effective.
+5.1 Memorization Metric
+Intuitively, the memorization of fixed training data points (i.e., point-wise memorization) can be quantified by the degree of overlap between the generated distribution and the distributions centered at each training data point.
+Given a generative model and training dataset, we propose the following memorization metric to quantify the memorization of the training data points in the model.
+Definition 1 (Point-wise Memorization).
+Given a generative model $p_{\theta}$ with parameters $\theta$ and training dataset ${\mathcal{D}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, degree of memorization in $p_{\theta}$ of ${\mathcal{D}}$ is defined as:
+${\mathcal{M}}_{p}({\mathcal{D}};p_{\theta})=\frac{1}{|{\mathcal{D}}|}\sum_{{\bm{x}}_{i}\in{\mathcal{D}}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}}$ (9)
+where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ denotes the $i$-th training sample, $N$ is the total number of training samples, $p_{\theta}({\bm{x}})$ represents the probability density function (PDF) of the generated samples, and $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the probability distribution centered at training sample ${\bm{x}}_{i}$ within a small radius $\epsilon$.
+A straightforward choice for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the Dirac delta function centered at training data point ${\bm{x}}_{i}$: $q({\bm{x}},{\bm{x}}_{i},\epsilon)=\delta({\bm{x}}-{\bm{x}}_{i})$.
+However, this would make Eq.
+(9) uncomputable as the Dirac delta function is zero beyond the $\epsilon$-neighborhood.
+Alternatively, we could let $q({\bm{x}},{\bm{x}}_{i},\epsilon)\sim{\mathcal{N}}(x_{i};\epsilon I)$.
+Note that in Eq.
+(9), a smaller value of ${\mathcal{M}}_{p}({\mathcal{D}};\theta)$ indicates more overlap between the two distributions and thus more memorization.
+As $\epsilon$ tends to 00, the measured memorization becomes more accurate, with the limit characterizing the memorization capability of $p_{\theta}$.
+5.2 Theoretical Analysis
+Based on the point-wise memorization metric, we perform a theoretical understanding that explains why conditional generative models achieve better memorization capacity.
+Our analysis is based on the concept of informative labels, which refers to information that can partition a dataset into various disjoint subsets.
+And we demonstrate that provided with an informative label, the conditional generative models achieve stronger memorization capacity.
+Informative Labels
+The concept of informative labels has been previously discussed as class labels [13].
+In this work, we introduce a more general definition that encompasses both class labels and random labels as special cases.
+We formally define an informative label as follows.
+Definition 2 (Informative Label).
+Let ${\mathcal{Y}}$ be an attribute of the data taking values in $\{y_{i}\}_{i=1}^{C}$.
+We refer to ${\mathcal{Y}}$ as an informative label if we can partition the dataset into several mutually disjoint subsets $\{{\mathcal{D}}_{i}\}_{i=1}^{C}$ based on the values of the attribute ${\mathcal{Y}}$.
+Here, the informative labels are not limited to the conventional class labels; they can also be text captions, features, or cluster information that can be used to group the training samples into subsets.
+The above definition states that an information label should be able to differentiate a subset of samples from others.
+An extreme case is that all samples have the same label; in this case, the label is not informative.
+According to our definition, class-wise and random labels are special cases of informative labels.
+Informative labels can be either explicit like class/random labels and text captions, or implicit like salient features or clusters.
+Main Results
+Here we introduce our main theorem on the memorization mechanism of conditional generative models and explain why SIDE is effective.
+We use ${\mathcal{D}}_{i}$ to denote the subset consisting of data with informative label ${\mathcal{Y}}=y_{i}$.
+In addition, we denote the distribution of data in the original dataset ${\mathcal{D}}$ by $p$.
+The distribution has expectation $\mu$ and covariance matrix $\Sigma$.
+Similarly, for each attribute $y_{i}$, we use $p_{i}$ to denote the distribution of the data with this attribute, and use $\mu_{i}$ and $\Sigma_{i}$ to denote its expectation and covariance matrix.
+Theorem 1.
+If a generative model $p_{\theta_{i}}$ matches the target distribution $p_{i}$ well for each $i$, i.e.
+$TV(p_{i},p_{\theta_{i}})=0$, and the data distribution for some informative label $y_{c}$ satisfies $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$, we have with probability 1:
+$\lim_{|{\mathcal{D}}_{c}|\to\infty}\lim_{\epsilon\to 0}\frac{{\mathcal{M}}_{p}({\mathcal{D}}_{c},p_{\theta_{c}})}{{\mathcal{M}}_{p}({\mathcal{D}}_{c},p_{\theta})}\leq 1$ (10)
+where $p_{\theta_{i}}$ denotes the distribution of generated data conditioned on ${\mathcal{Y}}=y_{i}$.
+The proof can be found in Appendix A.
+Theorem 1 states that any form of information labels can enhance memorization capacity in conditional generative models, including conventional class labels and random labels.
+The condition $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$ is reasonable because it ensures that the spread (variance) of the data within each class is not excessively large.
+$\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{c}}{\left({\bm{x}}_{i}-\mu_{c}\right)^{T}\left({\bm{x}}_{i}-\mu_{c}\right)}}{|{\mathcal{D}}_{c}|}\leq||\mu-\mu_{c}||_{2}^{2}$ (11)
+To understand this, we look at the derivation Eq.
+(11) from $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$.
+For a well-separated class, this squared distance should provide an upper bound on how spread out the class data can be.
+If $tr(\Sigma_{c})$ were large relative to $||\mu-\mu_{c}||_{2}^{2}$, it would imply that data samples represented by informative label $c$ are highly spread out, making it harder for the class labelled with $c$ to be well-formed and distinct.
+Thus, the condition ensures that the data samples form tight clusters around their means, making it reasonable for the class to be well-defined and easily detachable.
+This also explains why SIDE works.
+Chen et al.
+[8] found that unconditional DPMs learn clustered representations from training data.
+The data samples within each class should not be widely dispersed for effective clustering.
+Then, the samples should satisfy Eq.
+11.
+So, although these clustered representations are implicit, we could extract training data from unconditional DPMs if we can convert them from implicit informative labels into explicit labels.
+When conditioning on the explicit labels, the DPMs will exhibit more memorization effects indicated by Theorem 1, which makes room for data extraction attacks.
+6 Limitations:
+SIDE has certain limitations that warrant further research.
+SIDE mainly uses human-labelled data to extract data, which is scarce and limits its practical use.
+Future work could investigate hybrid approaches for generating implicit labels, such as employing unsupervised clustering or utilizing pretrained models to label synthetic datasets.
 7 Conclusion
-In this paper, we introduced a pointwise memorization metric to quantify memorization effects in DPMs.
-We provided a theoretical analysis of conditional memorization, offering a generalized definition of informative labels and clarifying that random labels can also be informative.
-We distinguish between explicit labels and implicit labels and propose a novel method, SIDE, to extract training data from unconditional diffusion models by constructing a surrogate condition.
-The key to this approach is training a time-dependent classifier using our TDKD technique.
-We empirically validate SIDE on subsets of the CelebA and CIFAR-10 datasets with two new memorization scores: AMS and UMS.
-We aim for our work to enhance the understanding of memorization mechanisms in diffusion models and inspire further methods to mitigate memorization.
+In this paper, we distinguish the difference between explicit labels and implicit labels and propose a novel method, SIDE, to extract training data from unconditional DPMs by constructing surrogate conditions.
+The core of SIDE lies in training a time-dependent classifier using TDKD technique.
+We empirically validate effectiveness of SIDE on subsets of CelebA and CIFAR-10 datasets with two new memorization scores: AMS and UMS.
+In addition, we introduce a point-wise memorization metric to quantify memorization effects in DPMs for theoretical analysis.
+Finally, we introduce theoretical understanding of memorization in conditional generative models, offering a generalized definition of informative labels, clarifying random labels can also be informative and explaining why SIDE works.
+Our work aims to deepen understanding of memorization in DPMs and inspire further methods to mitigate memorization.
+Supplementary Material
 Broader Impacts
-Our work proposes a training data extraction method targeted at diffusion models and the effectiveness of our proposed method has far exceeded random methods.
-This may be exploited by potential malicious users to attack existing generative diffusion models.
-By effectively avoiding the introduction of additional or unnecessary informative label information into the training process, the risk of a diffusion model being attacked by data extraction can be greatly reduced.
-Therefore, the contribution of our work to the community outweighs the risks it poses, especially considering that the scale of our extraction method has not reached the same scale of state-of-the-art diffusion models.
-Moreover, the theoretical and experimental understandings established in this work can help the community build a deeper understanding of the memorization mechanism of diffusion models, and inspire better defense methods.
+Our work introduces a memorization metric that not only quantifies memorization effects in diffusion models but also extends to deep learning models more broadly.
+This contribution is significant in enhancing our understanding of when and how models memorize training data, which is critical for addressing concerns about data privacy and model robustness.
+By providing a theoretical framework for conditional memorization, we pave the way for developing effective memorization mitigation strategies tailored for diffusion models.
+These advancements can lead to the design of more secure and trustworthy AI systems, reducing the risk of potential data leakage while fostering greater accountability in the use of generative technologies.
+Ultimately, our findings aim to empower researchers and practitioners to create models that better respect privacy, thus benefiting the wider AI community.
+Appendix A Proof of Theorem1
+By the assumption $TV(p_{i},p_{\theta_{i}})=0$ in the theorem, we can replace $p_{\theta_{c}}$ with $p_{c}$ and $p_{\theta}$ with $p$ to simplify the notation without compromising correctness.
+Firstly, we compute $\mathcal{M}_{point}(\mathcal{D}_{c};p_{c})$ here:
 […]
-Appendix A Proofs
-A.1 Preliminaries
-If $p(x)$ and $q(x)$ are normal distributions:
-$p(\bm{x})=\frac{1}{\sqrt{(2\pi)^{d}\operatorname{det}(\bm{\Sigma}_{p})}}\exp\left\{-\frac{1}{2}(\bm{x}-\bm{\mu_{p}})^{\top}\bm{\Sigma}_{p}^{-1}(\bm{x}-\bm{\mu}_{p})\right\}$ (16)
-$q(\bm{x})=\frac{1}{\sqrt{(2\pi)^{d}\operatorname{det}(\bm{\Sigma}_{q})}}\exp\left\{-\frac{1}{2}(\bm{x}-\bm{\mu}_{q})^{\top}\bm{\Sigma}_{q}^{-1}(\bm{x}-\bm{\mu}_{q})\right\}$ (17)
-Then we have:
-$\mathbb{E}_{\bm{x}\sim p(\bm{x})}\left[\left(\bm{x}-\bm{\mu}_{q}\right)^{\top}\mathbf{\Sigma}_{q}^{-1}\left({x}-\bm{\mu}_{q}\right)\right]$ (18)
-$=\mathrm{Tr}\left(\mathbf{\Sigma}_{q}^{-1}\mathbf{\Sigma}_{p}\right)+\left(\bm{\mu}_{p}-\bm{\mu}_{q}\right)^{\top}\mathbf{\Sigma}_{q}^{-1}\left(\bm{\mu}_{p}-\bm{\mu}_{q}\right)$ (19)
-$\mathbb{E}_{\bm{x}\sim q\left(\bm{x}\right)}\left[\left(\bm{x}-\bm{\mu}_{q}\right)^{\top}\mathbf{\Sigma}_{q}^{-1}\left(\bm{x}-\bm{\mu}_{q}\right)\right]=d$ (21)
-The entropy of $p(x)$:
-$H_{p}\left(\bm{x}\right)=\mathbb{E}_{\bm{x}\sim p\left(\bm{x}\right)}[-\log p(x)]=\frac{n}{2}(1+\log 2\pi)+\frac{1}{2}\log\det\mathrm{(}\bm{\Sigma}_{p})$ (22)
-The KL divergence between the two distributions is:
-$D_{KL}(p(\bm{x})\|q(\bm{x}))$ (23)
-$=\frac{1}{2}\left[\left(\bm{\mu}_{p}-\bm{\mu}_{q}\right)^{\top}\bm{\Sigma}_{q}^{-1}\left(\bm{\mu}_{p}-\bm{\mu}_{q}\right)-\log\operatorname{det}\left(\bm{\Sigma}_{q}^{-1}\bm{\Sigma}_{p}\right)+\operatorname{Tr}\left(\bm{\Sigma}_{q}^{-1}\bm{\Sigma}_{p}\right)-d\right]$ (24)
-A.2 Proof for theorem 1
-We begin by assuming that we have an encoder $f_{\theta_{E}}({\bm{x}})$ and a decoder $f_{\theta_{D}}({\bm{z}})$.
-The encoder $f_{\theta_{E}}({\bm{x}})$ maps the data samples ${\bm{x}}$ into the latent distribution ${\bm{z}}$, which is modeled by a normal distribution $N\left(\bm{\mu},\bm{\Sigma}\right)$, where ${\bm{z}}\in\mathbb{R}^{d}$ is the latent space of dimension $d$.
-The decoder $f_{\theta_{D}}({\bm{z}})$ maps the latent variables ${\bm{z}}$ back to the original data samples ${\bm{x}}$.
-This structure forms the basis of many variational autoencoder (VAE) frameworks, where we aim to optimize the relationship between ${\bm{x}}$ and ${\bm{z}}$ through probabilistic modeling.
-Transformation of Probability Distributions
-Based on the transformation of probability density functions (PDF) and the method of change of variables for multiple integrals, we can express the likelihoods as follows:
-1. Conditional Probability of ${\bm{x}}$ given $y=c$:
-$p_{\theta}\left({\bm{x}}\mid y=c\right)$ $=p_{\theta}\left({\bm{z}}\mid y=c\right)\det\left(\frac{\partial{\bm{z}}}{\partial{\bm{x}}}\right)$ (25)
-$=p_{\theta}\left({\bm{z}}\mid y=c\right)\det\left(\frac{\partial f_{\theta_{E}}({\bm{x}})}{\partial{\bm{x}}}\right)$ (26)
-Here, $p_{\theta}\left({\bm{x}}\mid y=c\right)$ is the probability of the data ${\bm{x}}$ conditioned on the label $y=c$, which depends on the latent variable ${\bm{z}}$.
-The term $\det\left(\frac{\partial f_{\theta_{E}}({\bm{x}})}{\partial{\bm{x}}}\right)$ is the determinant of the Jacobian matrix of the encoder function $f_{\theta_{E}}({\bm{x}})$, which accounts for the change of variables from ${\bm{x}}$ to ${\bm{z}}$.
-2. Marginal Probability of ${\bm{x}}$:
-$p_{\theta}\left({\bm{x}}\right)$ $=p_{\theta}\left({\bm{z}}\right)\det\left(\frac{\partial{\bm{z}}}{\partial{\bm{x}}}\right)$ (27)
-$=p_{\theta}\left({\bm{z}}\right)\det\left(\frac{\partial f_{\theta_{E}}({\bm{x}})}{\partial{\bm{x}}}\right)$ (28)
-This expression captures the marginal distribution of ${\bm{x}}$, which is obtained by marginalizing over the latent variable ${\bm{z}}$.
-3. Volume Element Change:
-$d{\bm{x}}$ $=\det\left(\frac{\partial{\bm{x}}}{\partial{\bm{z}}}\right)d{\bm{z}}$ (29)
-$=\det\left(\frac{\partial f_{\theta_{D}}({\bm{z}})}{\partial{\bm{z}}}\right)d{\bm{z}}$ (30)
-$=\det\left(\frac{\partial{\bm{x}}}{\partial f_{\theta_{E}}({\bm{x}})}\right)d{\bm{z}}$ (31)
-This represents the volume element transformation between the latent space ${\bm{z}}$ and the data space ${\bm{x}}$.
-The determinant of the Jacobian of the decoder $f_{\theta_{D}}({\bm{z}})$ relates the volume elements in ${\bm{z}}$ and ${\bm{x}}$.
-Objective Function and Change of Variables
-Given the above transformations, we can rewrite the memorization objective $\mathcal{M}_{point}(\mathcal{D};\theta)$ as follows:
-1. Original Memorization Objective:
-$\mathcal{M}_{point}(\mathcal{D};\theta)$ $=\sum_{\bm{x}_{i}\in\mathcal{D}}\int p_{\theta}(\bm{x})\log\frac{p_{\theta}(\bm{x})}{q(\bm{x},\bm{x}_{i},\epsilon)}\,d\bm{x}$ (32)
-This measures the difference between the model distribution $p_{\theta}(\bm{x})$ and a perturbed distribution $q(\bm{x},\bm{x}_{i},\epsilon)$ at each point $\bm{x}_{i}$ in the dataset $\mathcal{D}$.
-2. Transformed Memorization Objective in Latent Space:
-$\mathcal{M}_{semantic}(\mathcal{D};\theta)$ $=\sum_{\bm{z}_{i}\in\mathcal{D}}\int p_{\theta}(\bm{z})\log\frac{p_{\theta}(\bm{z})}{q(\bm{z},\bm{z}_{i},\epsilon)}\,d\bm{z}$ (33)
-By applying the change of variables, we transform the objective into the latent space, where the same logic applies, but now the integration is over the latent variables $\bm{z}$ instead of the data space $\bm{x}$.
-Monotonicity Derivation
-We now derive the monotonicity of the memorization objective with respect to the content-related part $\mathcal{M}_{semantic}(\mathcal{D};\theta)$.
-Specifically, we want to show that the objective increases monotonically with $\mathcal{M}_{semantic}$:
-1. Partial Derivative of the Objective:
-$\frac{\partial\mathcal{M}(\mathcal{D};\theta)}{\partial\mathcal{M}_{semantic}(\mathcal{D};\theta)}$ $=\frac{\partial}{\partial\mathcal{M}_{semantic}(\mathcal{D};\theta)}\left(\sum_{\bm{x}_{i}\in\mathcal{D}}\int p_{\theta}(\bm{x})\log\frac{p_{\theta}(\bm{x})}{q(\bm{x},\bm{x}_{i},\epsilon)}\,d\bm{z}\right)$ (34)
-2. Expanding the Derivatives:
-[…]
-3. Applying the Chain Rule:
-[…]
-4. Evaluating Each Term:
-[…]
-5. Final Form:
-$=\frac{1}{2\epsilon}+\frac{\operatorname{Tr}(\Sigma_{p_{\theta}}^{-1})}{2}$ (38)
-A.2.1 Derive $\frac{\partial\mathcal{H}(p_{\theta})}{\partial\mathcal{M}_{semantic}}$
-We are given:
-  • A multivariate normal distribution $p_{\theta}(\bm{z})=\mathcal{N}(\bm{\mu},\Sigma_{p_{\theta}})$, where $\bm{\mu}$ is the mean vector and $\Sigma_{p_{\theta}}$ is the covariance matrix.
-  • The entropy of this distribution is:
-    $H(p_{\theta})=\frac{1}{2}\ln\left((2\pi e)^{d}|\Sigma_{p_{\theta}}|\right)=\frac{d}{2}\ln(2\pi e)+\frac{1}{2}\ln|\Sigma_{p_{\theta}}|$ (39)
-  • The sum expression:
-    ${\mathcal{M}}_{semantic}=\sum_{\bm{z}_{i}\in\mathcal{D}}\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]=|\mathcal{D}|\operatorname{Tr}(\Sigma_{p_{\theta}})+\sum_{i=1}^{|\mathcal{D}|}\|\bm{\mu}-\bm{z}_{i}\|^{2}$ (40)
-    where $|\mathcal{D}|$ is the number of data points in $\mathcal{D}$.
-Our goal is to find $\frac{\partial H(p_{\theta})}{\partial{\mathcal{M}}_{semantic}}$.
-Entropy of a Multivariate Normal Distribution
-The entropy of a multivariate normal distribution $p_{\theta}(\bm{z})=\mathcal{N}(\bm{\mu},\Sigma_{p_{\theta}})$ is given by:
-[…]
-$=\frac{1}{2}\ln\left((2\pi e)^{d}|\Sigma_{p_{\theta}}|\right)$ (42)
-Where:
-  • $d$ is the dimensionality of the vector $\bm{z}$.
-  • $|\Sigma_{p_{\theta}}|$ denotes the determinant of the covariance matrix $\Sigma_{p_{\theta}}$.
-Thus, we can write:
-$H(p_{\theta})=\frac{d}{2}\ln(2\pi e)+\frac{1}{2}\ln|\Sigma_{p_{\theta}}|$ (43)
-Computing the Sum ${\mathcal{M}}_{semantic}$
-We have:
-${\mathcal{M}}_{semantic}=\sum_{\bm{z}_{i}\in\mathcal{D}}\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]$ (44)
-First, compute $\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]$ for each $\bm{z}_{i}$:
-Expanding the Squared Norm
-$\|\bm{z}-\bm{z}_{i}\|^{2}$ $=(\bm{z}-\bm{z}_{i})^{\top}(\bm{z}-\bm{z}_{i})$ (45)
-$=\bm{z}^{\top}\bm{z}-2\bm{z}^{\top}\bm{z}_{i}+\bm{z}_{i}^{\top}\bm{z}_{i}$ (46)
-Taking the Expectation
-$\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]$ $=\mathbb{E}_{p_{\theta}}\left[\bm{z}^{\top}\bm{z}\right]-2\bm{z}_{i}^{\top}\mathbb{E}_{p_{\theta}}\left[\bm{z}\right]+\bm{z}_{i}^{\top}\bm{z}_{i}$ (47)
-$=\operatorname{Tr}\left(\mathbb{E}_{p_{\theta}}\left[\bm{z}\bm{z}^{\top}\right]\right)-2\bm{z}_{i}^{\top}\bm{\mu}+\|\bm{z}_{i}\|^{2}$ (48)
-Computing the Expectations
-  • First Term: The second moment of $\bm{z}$:
-    $\mathbb{E}_{p_{\theta}}\left[\bm{z}\bm{z}^{\top}\right]=\Sigma_{p_{\theta}}+\bm{\mu}\bm{\mu}^{\top}$ (49)
-    Therefore:
-    $\operatorname{Tr}\left(\mathbb{E}_{p_{\theta}}\left[\bm{z}\bm{z}^{\top}\right]\right)=\operatorname{Tr}(\Sigma_{p_{\theta}})+\operatorname{Tr}(\bm{\mu}\bm{\mu}^{\top})=\operatorname{Tr}(\Sigma_{p_{\theta}})+\|\bm{\mu}\|^{2}$ (50)
-  • Second Term: The mean of $\bm{z}$:
-    $\mathbb{E}_{p_{\theta}}\left[\bm{z}\right]=\bm{\mu}$ (51)
-  • Third Term: Constant term involving $\bm{z}_{i}$:
-    $\|\bm{z}_{i}\|^{2}=\bm{z}_{i}^{\top}\bm{z}_{i}$ (52)
-Combining the Terms
-Substitute back into the expectation:
-$\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]$ $=\left(\operatorname{Tr}(\Sigma_{p_{\theta}})+\|\bm{\mu}\|^{2}\right)-2\bm{z}_{i}^{\top}\bm{\mu}+\|\bm{z}_{i}\|^{2}$ (53)
-$=\operatorname{Tr}(\Sigma_{p_{\theta}})+\left(\|\bm{\mu}\|^{2}-2\bm{\mu}^{\top}\bm{z}_{i}+\|\bm{z}_{i}\|^{2}\right)$ (54)
-$=\operatorname{Tr}(\Sigma_{p_{\theta}})+\|\bm{\mu}-\bm{z}_{i}\|^{2}$ (55)
-Thus, for each $\bm{z}_{i}$:
-$\mathbb{E}_{p_{\theta}}\left[\|\bm{z}-\bm{z}_{i}\|^{2}\right]=\operatorname{Tr}(\Sigma_{p_{\theta}})+\|\bm{\mu}-\bm{z}_{i}\|^{2}$ (56)
-Summing Over All Data Points
-The total sum ${\mathcal{M}}_{semantic}$ becomes:
-${\mathcal{M}}_{semantic}$ $=\sum_{i=1}^{|\mathcal{D}|}\left(\operatorname{Tr}(\Sigma_{p_{\theta}})+\|\bm{\mu}-\bm{z}_{i}\|^{2}\right)$ (57)
-$=|\mathcal{D}|\operatorname{Tr}(\Sigma_{p_{\theta}})+\sum_{i=1}^{|\mathcal{D}|}\|\bm{\mu}-\bm{z}_{i}\|^{2}$ (58)
-Let’s define the sample variance $\text{Var}_{\text{data}}$:
-$\text{Var}_{\text{data}}=\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\|\bm{\mu}-\bm{z}_{i}\|^{2}$ (59)
-Then, ${\mathcal{M}}_{semantic}$ can be expressed as:
-${\mathcal{M}}_{semantic}=|\mathcal{D}|\operatorname{Tr}(\Sigma_{p_{\theta}})+|\mathcal{D}|\text{Var}_{\text{data}}=|\mathcal{D}|\left(\operatorname{Tr}(\Sigma_{p_{\theta}})+\text{Var}_{\text{data}}\right)$ (60)
-Attempting to Relate $H(p_{\theta})$ and ${\mathcal{M}}_{semantic}$
-Our challenge is to express $H(p_{\theta})$ as a function of ${\mathcal{M}}_{semantic}$ so that we can compute $\frac{\partial H(p_{\theta})}{\partial{\mathcal{M}}_{semantic}}$.
-However, we face a difficulty:
-  • The entropy $H(p_{\theta})$ depends on $\ln|\Sigma_{p_{\theta}}|$.
-  • The sum ${\mathcal{M}}_{semantic}$ depends on $\operatorname{Tr}(\Sigma_{p_{\theta}})$.
-For a general covariance matrix $\Sigma_{p_{\theta}}$, there is no direct algebraic relationship between $\operatorname{Tr}(\Sigma_{p_{\theta}})$ and $\ln|\Sigma_{p_{\theta}}|$.
-Therefore, we need to explore an alternative method.
-Computing Derivatives with Respect to $\Sigma_{p_{\theta}}$
-The entropy $H(p_{\theta})$ is given by:
-$H(p_{\theta})=\frac{d}{2}\ln(2\pi e)+\frac{1}{2}\ln|\Sigma_{p_{\theta}}|$ (61)
-To find the derivative of $H(p_{\theta})$ with respect to $\Sigma_{p_{\theta}}$, we proceed as follows:
-$\frac{\partial H(p_{\theta})}{\partial\Sigma_{p_{\theta}}}$ $=\frac{\partial}{\partial\Sigma_{p_{\theta}}}\left(\frac{d}{2}\ln(2\pi e)+\frac{1}{2}\ln|\Sigma_{p_{\theta}}|\right)$ (62)
-$=\frac{1}{2}\frac{\partial}{\partial\Sigma_{p_{\theta}}}\ln|\Sigma_{p_{\theta}}|$ (63)
-$=\frac{1}{2}\Sigma_{p_{\theta}}^{-1}$ (64)
-Recall that:
-${\mathcal{M}}_{semantic}=|\mathcal{D}|\left(\operatorname{Tr}(\Sigma_{p_{\theta}})+\text{Var}_{\text{data}}\right)$ (65)
-Since $\text{Var}_{\text{data}}$ does not depend on $\Sigma_{p_{\theta}}$, the derivative of ${\mathcal{M}}_{semantic}$ with respect to $\Sigma_{p_{\theta}}$ is:
-$\frac{\partial{\mathcal{M}}_{semantic}}{\partial\Sigma_{p_{\theta}}}=|\mathcal{D}|\frac{\partial}{\partial\Sigma_{p_{\theta}}}\operatorname{Tr}(\Sigma_{p_{\theta}})$ (66)
-Computing $\frac{\partial H(p_{\theta})}{\partial{\mathcal{M}}_{semantic}}$ Using the Chain Rule
-Using the chain rule for derivatives:
-$\frac{\partial H(p_{\theta})}{\partial{\mathcal{M}}_{semantic}}=\operatorname{Tr}\left(\frac{\partial H(p_{\theta})}{\partial\Sigma_{p_{\theta}}}\cdot\left(\frac{\partial{\mathcal{M}}_{semantic}}{\partial\Sigma_{p_{\theta}}}\right)^{-1}\right)$ (67)
-Computing $\left(\frac{\partial{\mathcal{M}}_{semantic}}{\partial\Sigma_{p_{\theta}}}\right)^{-1}$
-Since $\frac{\partial{\mathcal{M}}_{semantic}}{\partial\Sigma_{p_{\theta}}}=|\mathcal{D}|I$, its inverse is:
-$\left(\frac{\partial{\mathcal{M}}_{semantic}}{\partial\Sigma_{p_{\theta}}}\right)^{-1}=\frac{1}{|\mathcal{D}|}I$ (68)
-Combine together, we can derive
-$\frac{\partial H(p_{\theta})}{\partial{\mathcal{M}}_{semantic}}=\frac{\operatorname{Tr}(\Sigma_{p_{\theta}}^{-1})}{2|\mathcal{D}|}$ (69)
-Thus, we conclude that the memorization objective increases monotonically with respect to the memorization metric.
-A.3 Proof for Proposition 1
-A.3.1 Proof for equation 8
-Covariance Definitions:
-$\Sigma$ $=\text{Cov}(Z),\quad\Sigma_{c}=\text{Cov}(Z\mid Y)$ (70)
-Here:
-  • $\Sigma$ is the overall covariance matrix of $Z$,
-  • $\Sigma_{c}$ is the conditional covariance matrix of $Z$ given $Y$.
-The goal is to prove that the trace of the conditional covariance matrix is less than or equal to the trace of the overall covariance matrix, i.e.,
-$\text{Tr}(\Sigma_{c})\leq\text{Tr}(\Sigma)$ (71)
-for all realizations of $Y$.
-Proof of Trace Inequality
-$\Sigma$ $=\text{Cov}(Z),\quad\Sigma_{c}=\text{Cov}(Z\mid Y)$ (72)
-Here:
-  • $\Sigma$ is the overall covariance matrix of the random vector $Z\in\mathbb{R}^{n}$,
-  • $\Sigma_{c}$ is the conditional covariance matrix of $Z$ given $Y$.
-Goal Prove that:
-$\text{Tr}(\Sigma_{c})\leq\text{Tr}(\Sigma)$ (73)
-for all realizations of $Y$.
-We begin by expressing the overall covariance matrix $\Sigma$ in terms of the conditional covariance matrix $\Sigma_{c}$ and the covariance of the conditional expectation of $Z$ given $Y$.
-$\Sigma$ $=\text{Cov}(Z)$ (74)
-$=\mathbb{E}\left[(Z-\mathbb{E}[Z])(Z-\mathbb{E}[Z])^{T}\right]$ (75)
-$=\mathbb{E}\left[\left(Z-\mathbb{E}[Z\mid Y]+\mathbb{E}[Z\mid Y]-\mathbb{E}[Z]\right)\left(Z-\mathbb{E}[Z\mid Y]+\mathbb{E}[Z\mid Y]-\mathbb{E}[Z]\right)^{T}\right]$ (76)
-Expanding the Product Inside the Expectation
-We expand the product inside the expectation:
-$\Sigma$ $=\mathbb{E}\Bigg{[}\underbrace{(Z-\mathbb{E}[Z\mid Y])(Z-\mathbb{E}[Z\mid Y])^{T}}_{\text{Term 1}}$
-$\quad+\underbrace{(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}}_{\text{Term 2}}$
-$\quad+\underbrace{(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])(Z-\mathbb{E}[Z\mid Y])^{T}}_{\text{Term 3}}$
-$\quad+\underbrace{(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}}_{\text{Term 4}}\Bigg{]}$ (77)
-Analyzing Each Term
-Term 1:
-$\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(Z-\mathbb{E}[Z\mid Y])^{T}\right]=\text{Cov}(Z\mid Y)=\Sigma_{c}$ (78)
-Term 2:
-$\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\right]$ (79)
-To evaluate Term 2, we condition on $Y$:
-$\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\right]=\mathbb{E}\left[\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\mid Y\right]\right]$ (80)
-Inside the inner expectation, $\mathbb{E}[Z\mid Y]-\mathbb{E}[Z]$ is treated as a constant with respect to $Z$, so:
-$\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\mid Y\right]$ (81)
-$=\left(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z]\right)\mathbb{E}\left[Z-\mathbb{E}[Z\mid Y]\mid Y\right]$ (82)
-$=\left(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z]\right)\cdot 0=0$ (83)
-Thus:
-$\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\right]=0$ (84)
-Term 3:
-$\mathbb{E}\left[(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])(Z-\mathbb{E}[Z\mid Y])^{T}\right]$ (85)
-Similarly, we condition on $Y$:
-$\mathbb{E}\left[(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])(Z-\mathbb{E}[Z\mid Y])^{T}\right]$ (86)
-$=\mathbb{E}\left[(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])\mathbb{E}\left[(Z-\mathbb{E}[Z\mid Y])^{T}\mid Y\right]\right]$ (87)
-$=\mathbb{E}\left[(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])\cdot 0^{T}\right]$ (88)
-$=0$ (89)
-Term 4:
-$\mathbb{E}\left[(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])(\mathbb{E}[Z\mid Y]-\mathbb{E}[Z])^{T}\right]=\text{Cov}\left(\mathbb{E}[Z\mid Y]\right)=\text{Cov}(\mathbb{E}[Z\mid Y])$ (90)
-Combining All Terms
-Putting all the terms together:
-$\Sigma$ $=\Sigma_{c}+0+0+\text{Cov}(\mathbb{E}[Z\mid Y])$ (91)
-$=\Sigma_{c}+\text{Cov}(\mathbb{E}[Z\mid Y])$ (92)
-Taking the Trace
-Taking the trace on both sides of the covariance decomposition:
-$\text{Tr}(\Sigma)$ $=\text{Tr}\left(\Sigma_{c}\right)+\text{Tr}\left(\text{Cov}(Z\mid Y)\right)$ (93)
-$=\left[\text{Tr}(\Sigma_{c})\right]+\text{Tr}\left(\text{Cov}(\mathbb{E}[Z\mid Y])\right)$ (94)
-Since the trace of a covariance matrix is non-negative:
-$\text{Tr}\left(\text{Cov}(\mathbb{E}[Z\mid Y])\right)\geq 0$ (95)
-Thus:
-$\text{Tr}(\Sigma_{c})\leq\text{Tr}(\Sigma)$ (96)
-Conclusion
-We have proven that the trace of the conditional covariance matrix $\Sigma_{c}$ is less than or equal to the trace of the overall covariance matrix $\Sigma$. Formally,
-$\text{Tr}(\Sigma_{c})\leq\text{Tr}(\Sigma)$ (97)
-This result is a multivariate generalization of the variance decomposition, showing that the expected conditional variability of $Z$ given $Y$ does not exceed the overall variability of $Z$.
-A.3.2 Proof for equation 7
-Centering and Decomposing
-By the definition of variance, we can express the sum of squared distances for the conditional mean $\bm{\mu}_{c}$ and overall mean $\bm{\mu}$:
-$\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu})$ (98)
-$\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu})$ $=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}\left(({\bm{z}}_{i}-\bm{\mu}_{c}+\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c}+\bm{\mu}_{c}-\bm{\mu})\right)$ (99)
-$=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}\Big{[}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})$
-$\quad+({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$
-$\quad+(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})$
-$\quad+(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})\Big{]}$ (100)
-$=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})$
-$\quad+\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$
-$\quad+\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})$
-$\quad+\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$ (101)
-$=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})$
-$\quad+(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}1$ (102)
-$=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})+|{\mathcal{D}}_{y=c}|(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$ (103)
-Expanding this expression, we have:
-$=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}\left(({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})+2({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})+(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})\right).$ (104)
-Step 2: Simplifying the Inequality
-The term $(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$ is a constant for the conditional samples, and the term $2({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu})$ sums to zero when averaged over the samples in ${\mathcal{D}}_{y=c}$ due to the definition of $\bm{\mu}_{c}$.
-Thus, we have:
-$\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu})=\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})+|{\mathcal{D}}_{y=c}|(\bm{\mu}_{c}-\bm{\mu})^{\mathrm{T}}(\bm{\mu}_{c}-\bm{\mu}),$ (105)
-where $|{\mathcal{D}}_{y=c}|$ is the number of samples with label $y=c$.
-Step 3: Establishing the Inequality
-Since the variance (sum of squared distances) from the conditional mean will always be less than or equal to the variance from the overall mean, we conclude:
-$\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu}_{c})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}_{c})\leq\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}}({\bm{z}}_{i}-\bm{\mu})^{\mathrm{T}}({\bm{z}}_{i}-\bm{\mu}).$ (106)
-This establishes that the latent space representation conditioned on an informative label exhibits reduced variance, confirming our initial claim.
-Generalized Variance Comparison
-In this section, we examine a broader comparison of variance that does not restrict the analysis to the subset ${\mathcal{D}}_{y=c}$ but rather considers the entire dataset ${\mathcal{D}}$.
-Namely, we prove that:
-$\sum_{\bm{z}_{i}\in{\mathcal{D}}_{y=c}}{(}\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})\leq\sum_{\bm{z}_{i}\in{\mathcal{D}}}{(}\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (107)
-For a set of data points $D$, the sum of squared deviations from the mean is given by:
-$\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (108)
-Derivation:
-First, let’s write out the sums of squared deviations for the unconditional and conditional cases.
-Unconditional Sum of Squared Deviations:
-$\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (109)
-where $D$ represents the entire dataset of $\bm{z}$ values.
-Conditional Sum of Squared Deviations:
-$\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})$ (110)
-where $D_{y=c}$ represents the subset of data points $\bm{z}_{i}$ where $y=c$.
-Covariance Matrix and Sum of Squared Deviations:
-The covariance matrix can be related to the sum of squared deviations. For the unconditional case:
-$\Sigma=\frac{1}{n}\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})(\bm{z}_{i}-\bm{\mu})^{T}$ (111)
-Taking the trace on both sides:
-$\text{tr}(\Sigma)=\frac{1}{n}\text{tr}\left(\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})(\bm{z}_{i}-\bm{\mu})^{T}\right)$ (112)
-Since the trace of a sum is the sum of the traces:
-$\text{tr}(\Sigma)=\frac{1}{n}\sum_{\bm{z}_{i}\in D}\text{tr}\left((\bm{z}_{i}-\bm{\mu})(\bm{z}_{i}-\bm{\mu})^{T}\right)$ (113)
-The trace of the outer product of a vector with itself is the sum of squared elements of the vector:
-$\text{tr}\left((\bm{z}_{i}-\bm{\mu})(\bm{z}_{i}-\bm{\mu})^{T}\right)=(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (114)
-Therefore:
-$\text{tr}(\Sigma)=\frac{1}{n}\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (115)
-Similarly, for the conditional case:
-$\Sigma_{c}=\frac{1}{n_{c}}\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})(\bm{z}_{i}-\bm{\mu}_{c})^{T}$ (116)
-Taking the trace:
-$\text{tr}(\Sigma_{c})=\frac{1}{n_{c}}\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})$ (117)
-Inequality of Traces:
-Given that conditioning on $y=c$ provides information about $\bm{z}$, it generally reduces the variance of $\bm{z}$. Mathematically, this can be expressed as:
-$\text{tr}(\Sigma_{c})\leq\text{tr}(\Sigma)$ (118)
-In terms of sums of squared deviations:
-$\frac{1}{n_{c}}\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})\leq\frac{1}{n}\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (119)
-Multiplying both sides by their respective sample sizes $n_{c}$ and $n$:
-$\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})\leq\frac{n_{c}}{n}\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (120)
-Since $n_{c}\leq n$, this further simplifies to:
-$\sum_{\bm{z}_{i}\in D_{y=c}}(\bm{z}_{i}-\bm{\mu}_{c})^{T}(\bm{z}_{i}-\bm{\mu}_{c})\leq\sum_{\bm{z}_{i}\in D}(\bm{z}_{i}-\bm{\mu})^{T}(\bm{z}_{i}-\bm{\mu})$ (121)
-A.4 Proof for theorem 2
-This section will detail the proof for the theorem 2.
-$\lim_{\epsilon\to 0}\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{y=c}}{\int{p_{\theta}}}({\bm{x}}|y=c)\log\frac{p_{\theta}({\bm{x}}|y=c)}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,dx}{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{y=c}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}}}\leq 1$ (122)
-Define ${\mathcal{D}}_{y=c}^{z}=\{{\bm{z}}_{i}:f_{\theta_{E}}({\bm{x}}_{i})\in{\mathcal{D}}_{y=c}\}$
-${\mathcal{D}}_{y=c}=\{{\bm{x}}_{i}:{\bm{x}}_{i}\in\mathcal{D},y_{i}=c\}$
-By using the change of variable theorem, the (122) becomes:
-$\lim_{\epsilon\to 0}\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{y=c}}{\int{p_{\theta}}}({\bm{x}}|y=c)\log\frac{p_{\theta}({\bm{x}}|y=c)}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,dx}{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{y=c}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}}}\leq 1$ (123)
-$\Rightarrow\,\,\lim_{\epsilon\to 0}\frac{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\int{p_{\theta}}}({\bm{z}}|y=c)\log\frac{p_{\theta}({\bm{z}}|y=c)}{q({\bm{z}},{\bm{z}}_{i},\epsilon)}\,d{\bm{z}}}{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\int{p_{\theta}}}({\bm{z}})\log\frac{p_{\theta}({\bm{z}})}{q({\bm{z}},{\bm{z}}_{i},\epsilon)}\,d{\bm{z}}}\leq 1$ (124)
-Since $p_{\theta}({\bm{z}})=N\left(\bm{\mu},\bm{\Sigma}\right)$, it is reasonable to assume that its conditional distribution is also a normal distribution. Then:
-$p_{\theta}\left({\bm{z}}|y=c\right)=N\left(\bm{\mu}_{c},\bm{\Sigma}_{c}\right)$ (125)
-where $\bm{\mu}_{c}\in\mathbb{R}^{d},\bm{\Sigma}_{c}\in\mathbb{R}^{d\times d}$. Moreover, because $p_{\theta}\left({\bm{z}}|y=c\right)$ depends on label $c$, we can derive the following:
-$\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\left({\bm{z}}_{i}-\bm{\mu}_{c}\right)^{\mathrm{T}}\left({\bm{z}}_{i}-\bm{\mu}_{c}\right)}\leq\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\left(\bm{z}_{i}-\bm{\mu}\right)^{\mathrm{T}}\left(\bm{z}_{i}-\bm{\mu}\right)}$ (126)
-where $\forall{\bm{z}}_{i}$, $f_{\theta_{D}}({\bm{z}}_{i})\in y_{c}$.
-Intuitively, (126) means that the latent code of each training sample conditioned on the label $y=c$ is more centered around the learned latent space of distribution $p_{\theta}({\bm{z}}|y=c)$ than around the distribution $p_{\theta}({\bm{z}})$.
-We now look into the KL divergence:
-$\int{p_{\theta}}(\bm{z}|y=c)\log\frac{p_{\theta}(\bm{z}|y=c)}{q(\bm{z};\bm{z}_{i})}\,d\bm{z}$ (127)
-[…]
-$=-\frac{d}{2}\left(1+\log 2\pi\right)-\frac{1}{2}\log\det\left(\bm{\Sigma}_{c}\right)+\mathbb{E}_{\bm{z}\sim p_{\theta}(\bm{z}|y=c)}\left(-\log q\left(\bm{z};\bm{z}_{i}\right)\right)$ (129)
-$=\frac{1}{2}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}_{c}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)}{\epsilon}-\log\frac{\det\left(\mathbf{\Sigma}_{c}\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\mathbf{\Sigma}_{c}\right)}{\epsilon}-d\right]$ (130)
-We use the SVD decomposition to decompose $\bm{\Sigma}_{c}$:
-$\bm{\Sigma}_{c}=U_{c}\varLambda_{c}U_{c}^{\mathrm{T}}$ (131)
-And:
-$\log\det\mathbf{\Sigma}_{c}=\log\det U_{c}\varLambda U_{c}^{\mathrm{T}}=\log|U_{c}||\varLambda_{c}||U_{c}^{\mathrm{T}}|=\log|\varLambda_{c}|$ (132)
-$\mathrm{Tr}\left(\mathbf{\Sigma}_{c}\right)=\mathrm{Tr}\left(U_{c}\varLambda_{c}U_{c}^{\mathrm{T}}\right)=\mathrm{Tr}\left(\varLambda_{c}U_{c}U_{c}^{\mathrm{T}}\right)=\mathrm{Tr}\left(\varLambda_{c}\right)$ (133)
-Thus, (130) simplifies to:
-$\frac{1}{2}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}_{c}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)}{\epsilon}-\log\frac{\det\left(\varLambda_{c}\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\varLambda_{c}\right)}{\epsilon}-d\right]$ (134)
-Similarly:
-$\int{p_{\theta}}(\bm{z})\log\frac{p_{\theta}(\bm{z})}{q(\bm{z};\bm{z}_{i})}\,d\bm{z}$ (136)
-$=\frac{1}{2}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}\right)}{\epsilon}-\log\frac{\det\left(\mathbf{\Sigma}_{c}\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\mathbf{\Sigma}_{c}\right)}{\epsilon}-d\right]$ (137)
-$=\frac{1}{2}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}\right)}{\epsilon}-\log\frac{\det\left(\varLambda\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\varLambda\right)}{\epsilon}-d\right]$ (138)
-where
-$\bm{\Sigma}=U\varLambda U^{\mathrm{T}}$ (139)
-According to the equation (8), the nuclear norm of the two covariance matrices differs. Specifically:
-$\|\mathbf{\Sigma}_{c}\|_{*}\leq\|\mathbf{\Sigma}\|_{*}$ (140)
-Thus, according to the definition of the nuclear norm, we have:
-$\mathrm{Tr}\left(\varLambda_{c}\right)\leq\mathrm{Tr}\left(\varLambda\right)$ (141)
-Therefore:
-$\lim_{\epsilon\to 0}\frac{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\int{p_{\theta}}({\bm{z}}|y=c)\log\frac{p_{\theta}({\bm{z}}|y=c)}{q({\bm{z}},{\bm{z}}_{i},\epsilon)}\,d{\bm{z}}}}{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}{\int{p_{\theta}}({\bm{z}})\log\frac{p_{\theta}({\bm{z}})}{q({\bm{z}},{\bm{z}}_{i},\epsilon)}\,d{\bm{z}}}}$ (142)
-$\Rightarrow\lim_{\epsilon\to 0}\frac{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}_{c}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)}{\epsilon}-\log\frac{\det\left(\varLambda_{c}\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\varLambda_{c}\right)}{\epsilon}-d\right]}{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left[\frac{\left(\bm{z}_{i}-\bm{\mu}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}\right)}{\epsilon}-\log\frac{\det\left(\varLambda\right)}{\epsilon^{d}}+\frac{\mathrm{Tr}\left(\varLambda\right)}{\epsilon}-d\right]}$ (143)
-Using L’Hospital’s rule:
-$\lim_{\epsilon\rightarrow 0}\frac{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left[-\frac{\left(\bm{z}_{i}-\bm{\mu}_{c}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)}{\epsilon^{2}}+\frac{d}{\epsilon}-\frac{\mathrm{Tr}\left(\varLambda_{c}\right)}{\epsilon^{2}}\right]}{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left[-\frac{\left(\bm{z}_{i}-\bm{\mu}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}\right)}{\epsilon^{2}}+\frac{d}{\epsilon}-\frac{\mathrm{Tr}\left(\varLambda\right)}{\epsilon^{2}}\right]}$ (144)
-$=\frac{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}_{c}\right)+\mathrm{Tr}\left(\varLambda_{c}\right)}{\sum_{{\bm{z}}_{i}\in{\mathcal{D}}_{y=c}^{z}}\left(\bm{z}_{i}-\bm{\mu}\right)^{\top}\left(\bm{z}_{i}-\bm{\mu}\right)+\mathrm{Tr}\left(\varLambda\right)}$ (145)
-$\leq 1$ (146)
-Appendix B Explanation of the $N_{G}$’s importance in fair comparison
+In the following, we will omit the expression ”almost surely”. Similarly, for $\mathcal{M}_{p}(\mathcal{D}_{c};p)$:
+$\begin{split}\mathcal{M}_{p}(\mathcal{D}_{c};p)&=o(\frac{1}{\epsilon})+\frac{1}{2\epsilon|\mathcal{D}_{c}|}\sum_{i=1}^{|\mathcal{D}_{c}|}(tr(\Sigma)+(\mu-z_{i})^{T}(\mu-z_{i}))\\ &\rightarrow o(\frac{1}{\epsilon})+\frac{1}{2\epsilon}(tr(\Sigma)+tr(\Sigma_{c})+||\mu-\mu_{c}||^{2}_{2})\end{split}$ (13)
+By definition, $p(x)=\sum_{k}\pi_{k}p_{k}(x)$ ,where $\pi_{k}$ denotes the probability that the informative label $\mathcal{Y}$ equals $y_{k}$.
+$\begin{split}\Sigma=\underset{x\sim p(x)}{\mathrm{Cov}}(x)&=\mathbb{E}_{p}[xx^{T}]-\mathbb{E}_{p}[x]\mathbb{E}_{p}[x]^{T}\\ &=\sum_{k}\pi_{k}\mathbb{E}_{p_{k}}[zz^{T}]-\mu\mu^{T}\\ &=\sum_{k}\pi_{k}(\mu_{k}\mu_{k}^{T}+\Sigma_{k})-\mu\mu^{T}\\ &=\mathbb{E}_{\pi}[\Sigma_{k}]+\mathbb{E}_{\pi}[(\mu-\mu_{k})(\mu-\mu_{k})^{T}]\end{split}$ (14)
+Finally, we have
+$\begin{split}\lim_{|\mathcal{D}|_{c}\rightarrow\infty}\lim_{\epsilon\rightarrow 0}\frac{\mathcal{M}_{p}(\mathcal{D}_{c};p_{c})}{\mathcal{M}_{p}(\mathcal{D}_{c};p)}&=\frac{2tr(\Sigma_{c})}{tr(\Sigma_{c})+\mathbb{E}_{\pi}[tr(\Sigma_{k})]+\mathbb{E}_{\pi}[||\mu-\mu_{k}||^{2}_{2}]+||\mu-\mu_{c}||^{2}_{2}}\end{split}$ (15)
+Here we use $\mathbb{E}_{\pi}[tr(\Sigma_{k})]$ and $\mathbb{E}_{\pi}[||\mu-\mu_{k}||_{2}^{2}]$ to denote $\sum_{k}\pi_{k}tr(\Sigma_{k})$ and $\sum_{k}\pi_{k}||\mu-\mu_{k}||_{2}^{2}$.
+Clearly, by Jensen’s inequality we have:
+$\mathbb{E}_{\pi}[\lim_{|\mathcal{D}_{c}|\rightarrow\infty}\lim_{\epsilon\rightarrow 0}\frac{\mathcal{M}_{p}(\mathcal{D}_{c};p_{c})}{\mathcal{M}_{p}(\mathcal{D}_{c};p)}]\leq 1$ (16)
+The equality holds only when $\mu_{k}$ and $tr(\Sigma_{k})$ are the same for all $k$. And under the assumption $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$, we have:
+$\lim_{|\mathcal{D}_{c}|\rightarrow\infty}\lim_{\epsilon\rightarrow 0}\frac{\mathcal{M}_{p}(\mathcal{D}_{c};p_{c})}{\mathcal{M}_{p}(\mathcal{D}_{c};p)}\leq 1$ (17)
+Hence we have completed the proof of Theorem1. ∎
+Appendix B Explanation UMS’s importance in fair comparison
 By highlighting the significance of $N_{G}$, UMS allows a fair comparison among extraction methods.
@@
 Comparing UMS across different $N_{G}$ values is flawed because varying $N_{G}$ leads to different outcomes, underscoring the need for consistent $N_{G}$ values to ensure fair comparisons.
-In carlini2023extracting, the significance of $N_{G}$ was overlooked, as they only reported the number of uniquely extracted images.
-Our theoretical analysis accurately aligns with the behavior observed in the experimental data in Fig 5.
+In [5], the significance of $N_{G}$ was overlooked, as they only reported the number of uniquely extracted images.
+Our theoretical analysis accurately aligns with behavior observed in the experimental data in Fig 6.
+Appendix C Refinement ResNet block (Fig. 7)
 Figure 6: Validation of $N_{G}$ ’s significance
-Appendix C Refinement ResNet block
 Figure 7: Refinement ResNet block with time-dependent module integration.
@@
 Appendix D Results on Classifier Choice
-Figure 7: Scatter plots showing the relationship between the number of classes and two performance metrics, AMS (left) and UMS (right).
-The fitted regression lines demonstrate a positive correlation in both cases.
 The analysis of classifier performance across varying numbers of classes reveals interesting patterns for both AMS and UMS.
@@
 The following Algorithm 1 outlines the detailed steps of the SIDE method for extracting training data from unconditional diffusion models.
-This algorithm combines the construction of implicit informative labels, the training of a time-dependent classifier, and the conditional generation process to effectively extract valuable training samples.
+This algorithm combines the construction of implicit informative labels, the training of a time-dependent classifier, and the conditional generation process to extract valuable training samples effectively.
 Algorithm 1 SIDE Method for Extracting Training Data from Unconditional Diffusion Models
@@
 19:     for $i=1$ to $N_{G}$ do
-20:         Sample $\bm{x}_{0}$ conditioned on $y=c$ using equation (151)
+20:         Sample $\bm{x}_{0}$ conditioned on $y=c$ using Eq. (22)
 21:         Compute gradient of cross-entropy loss:
@@
 Appendix F SSCD-related metrics
-In our evaluation of the proposed method, it is essential to include a discussion of the 95th percentile SSCD metric alongside our newly introduced metrics, AMS (Average Matching Similarity) and UMS (Unconditional Matching Similarity).
+In our evaluation of the proposed method, it is essential to include a discussion of the 95th percentile SSCD metric alongside our newly introduced metrics, AMS and UMS.
 While 95th percentile SSCD metric has its limitations, it still serves as a useful reference point for assessing the relative similarity between extracted images and the original training dataset.
 As shown in Table 3, we provide a comprehensive comparison of performance across different datasets.
-In addition, we report detailed metrics in the context of our experiments on the CelebA-HQ-FI dataset, which further illustrates the effectiveness of our approach.
-These results underscore the significance of incorporating SSCD-related scores to complement AMS and UMS in providing a more nuanced understanding of the similarities between generated and training images.
+In addition, we report detailed metrics in the context of our experiments on the CelebA-HQ-FI dataset, further illustrating our approach’s effectiveness.
 Table 3: Performance Comparison on CelebA-HQ-FI and CelebA-25000 Datasets
 Dataset Method Top 0.01% Top 0.05% Top 0.1% Top 0.5% Top 1.0% Top 5.0%
-CelebA-HQ-FI Uncond 0.656 0.624 0.604 0.544 0.518 0.463
+CelebA-HQ-FI Uncond 0.636 0.604 0.584 0.524 0.498 0.433
 SIDE(ours) 0.680 0.639 0.618 0.567 0.544 0.485
-CelebA-25000 Uncond 0.565 0.538 0.525 0.491 0.475 0.434
+CelebA-25000 Uncond 0.545 0.518 0.505 0.471 0.455 0.404
 SIDE(ours) 0.585 0.554 0.539 0.506 0.491 0.450
-Table 4: Generate Training Epoch: 3000 Dataset: CelebA-HQ-FI Generate Nums Per $\lambda$: 50000. The AMS and UMS is measured on Mid Similarity
-AMS(%) UMS(%) Top 0.1% Top 0.5% Top 1.0% Top 5.0% Top 10.0%
-$\lambda$
-0 0.596 0.328 0.604 0.544 0.518 0.463 0.440
-1 0.588 0.312 0.596 0.540 0.517 0.463 0.440
-2 0.640 0.350 0.591 0.541 0.518 0.465 0.441
-3 0.764 0.386 0.594 0.549 0.525 0.470 0.446
-4 0.850 0.390 0.604 0.553 0.529 0.473 0.448
-5 0.952 0.436 0.596 0.551 0.530 0.476 0.451
-6 1.092 0.414 0.611 0.560 0.536 0.480 0.454
-7 1.110 0.446 0.607 0.562 0.539 0.482 0.457
-8 1.148 0.444 0.618 0.566 0.542 0.484 0.458
-9 1.274 0.478 0.615 0.567 0.544 0.485 0.459
-10 1.338 0.444 0.613 0.569 0.546 0.487 0.461
-11 1.292 0.454 0.604 0.562 0.541 0.486 0.460
-12 1.262 0.406 0.617 0.567 0.544 0.486 0.460
-13 1.390 0.432 0.617 0.569 0.546 0.489 0.462
-14 1.232 0.384 0.613 0.567 0.544 0.485 0.459
-15 1.516 0.462 0.616 0.570 0.548 0.490 0.463
-16 1.280 0.390 0.612 0.566 0.543 0.487 0.461
-17 1.282 0.386 0.605 0.561 0.541 0.486 0.460
-18 1.330 0.374 0.616 0.569 0.545 0.488 0.461
-19 1.204 0.354 0.612 0.564 0.541 0.485 0.460
-20 1.178 0.358 0.603 0.559 0.538 0.483 0.458
-21 1.172 0.342 0.617 0.566 0.542 0.484 0.459
-22 1.208 0.368 0.602 0.560 0.539 0.485 0.459
-23 1.286 0.302 0.607 0.561 0.540 0.485 0.459
-24 1.244 0.352 0.597 0.558 0.538 0.484 0.458
-25 1.198 0.340 0.599 0.560 0.538 0.483 0.458
-26 1.220 0.338 0.601 0.559 0.539 0.483 0.458
-27 1.128 0.320 0.608 0.561 0.538 0.483 0.457
-28 1.102 0.314 0.604 0.556 0.534 0.481 0.456
-29 1.034 0.290 0.595 0.556 0.534 0.481 0.456
-30 1.026 0.326 0.602 0.557 0.535 0.480 0.455
-31 1.020 0.268 0.591 0.551 0.531 0.479 0.455
-32 1.054 0.282 0.593 0.551 0.531 0.479 0.455
-33 1.106 0.306 0.600 0.555 0.535 0.481 0.456
-34 1.062 0.288 0.582 0.547 0.529 0.479 0.454
-35 0.922 0.266 0.587 0.547 0.527 0.477 0.453
-36 0.874 0.260 0.585 0.545 0.525 0.477 0.453
-37 0.964 0.258 0.589 0.549 0.528 0.477 0.452
-38 0.888 0.246 0.582 0.543 0.524 0.475 0.452
-39 0.940 0.274 0.587 0.548 0.528 0.476 0.452
-40 0.808 0.234 0.587 0.544 0.524 0.474 0.451
-41 0.870 0.252 0.582 0.543 0.524 0.476 0.452
-42 0.872 0.238 0.584 0.543 0.523 0.475 0.451
-43 0.856 0.244 0.584 0.545 0.525 0.475 0.451
-44 0.796 0.212 0.578 0.540 0.521 0.473 0.449
-45 0.770 0.242 0.580 0.538 0.519 0.472 0.449
-46 0.774 0.218 0.580 0.540 0.521 0.472 0.448
-47 0.754 0.214 0.581 0.542 0.521 0.471 0.448
-48 0.716 0.218 0.572 0.536 0.518 0.471 0.448
-49 0.694 0.216 0.570 0.533 0.515 0.469 0.446
-50 0.728 0.204 0.576 0.535 0.518 0.471 0.447
+Appendix G Extended SIDE for Text-to-Image Diffusion Models
+We extend the SIDE method for a text-to-image diffusion model in a fine-tuning scenario.
+Instead of using a classifier as a surrogate condition to extract training data, we use random strings as surrogate conditions.
+The attack setting is simple and practical: the attacker can access the fine-tuned model and modify the text inputs in the text-to-image pairs.
+Notably, the attacker does not have access to the training data.
+We name this method as Extended SIDE.
+G.1 Attack Setting
+The attack setting for the Extended SIDE method involves two primary components: the victim’s fine-tuning process and the attacker’s access to the fine-tuned model.
+Below, we provide a detailed description of each component.
+  • Victim Fine-Tuning:
+    The victim fine-tunes a text-to-image diffusion model, denoted as $p_{\theta}(\bm{x}\mid s)$, using a dataset of image-text pairs $\mathcal{D}_{train}=\{(s_{i},\bm{x}_{i})\}_{i=1}^{N}$.
+In this context:
+    • $s_{i}\in\mathcal{S}_{s}$ are random strings crafted by the attacker to serve as surrogate text prompts.
+    • $\bm{x}_{i}\in\mathbb{R}^{d}$ represents the corresponding images associated with each random string.
+    The fine-tuning objective is to optimize the model parameters $\theta$ to minimize the loss function over the training data:
+    $\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(s,\bm{x})\sim\mathcal{D}_{train}}\left[\mathcal{L}(p_{\theta}(\bm{x}\mid s),\bm{x})\right]$ (28)
+    where $\mathcal{L}$ denotes the loss function, such as mean squared error (MSE) or cross-entropy loss.
+The use of random strings as surrogate conditions aims to induce memorization within the model, causing it to overfit specific image-text associations during the fine-tuning process.
+  • Attacker’s Access:
+    After the fine-tuning phase, the attacker gains access to the fine-tuned model $p_{\theta}(\bm{x}\mid s)$. This access allows the attacker to perform the following actions:
+    • Input Manipulation: The attacker can input any text prompt $s\in\mathcal{S}_{s}$ into the model to generate corresponding images:
+      $\bm{x}\sim p_{\theta}(\bm{x}\mid s)$ (29)
+    • Data Extraction: By systematically querying the model with the same set of random strings $\mathcal{S}_{s}$, the attacker can generate multiple images $\{\bm{x}_{i,j}\}_{j=1}^{N_{G}}$ for each string $s_{i}$.
+The goal is to identify images that closely resemble the memorized training samples:
+      $\bm{x}_{i,j}\sim p_{\theta}(\bm{x}\mid s_{i}),\quad j=1,\dots,N_{G}$ (30)
+    • Memorization Detection: The attacker analyzes the consistency and similarity among the generated images for each random string $s_{i}$.
+A low variance $\sigma_{s_{i}}^{2}$ among the samples indicates potential memorization:
+      $\sigma_{s_{i}}^{2}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i,j}-\overline{\bm{x}}_{s_{i}}\|^{2}$ (31)
+      where $\overline{\bm{x}}_{s_{i}}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i,j}$ is the mean image for string $s_{i}$.
+If $\sigma_{s_{i}}^{2}$ falls below a predefined threshold $\tau$, the attacker considers $\overline{\bm{x}}_{s_{i}}$ as a potential memorized training sample.
+  • Assumptions:
+    • Knowledge of Surrogate Strings: The attacker is aware of the set of random strings $\mathcal{S}_{s}$ used during the victim’s fine-tuning process.
+    • Model Query Capability: The attacker can perform a large number of queries to the fine-tuned model with each random string $s_{i}$ to generate sufficient samples for variance analysis.
+    • No Access to Training Data: The attacker does not have direct access to the victim’s training dataset $\mathcal{D}_{train}$.
+  • Attack Goals:
+    • Data Reconstruction: To reconstruct or approximate the training images $\bm{x}_{i}$ by exploiting the model’s memorization of the random string-image associations.
+    • Privacy Breach: To demonstrate the potential privacy risks associated with fine-tuned diffusion models by extracting sensitive or proprietary training data.
+G.2 Extended SIDE Method
+The Extended SIDE method leverages random strings as surrogate conditions to directly extract training images from a fine-tuned diffusion model.
+Unlike the original SIDE method, which relies on a classifier to guide the extraction process, this approach simplifies the attack by using random strings to induce and exploit memorization within the model.
+Algorithm 2 Extended SIDE Method for Extracting Training Data Using Random Strings
+1:Input:
+  • Fine-tuned diffusion model $p_{\theta}(\bm{x}\mid s)$
+  • Set of random strings $\mathcal{S}_{s}$
+  • Number of samples per string $N_{G}$
+  • Variance threshold $\tau$
+2:Output: Extracted data $\mathcal{D}_{\text{extracted}}$
+3:Initialize $\mathcal{D}_{\text{extracted}}\leftarrow\emptyset$
+4:for each $s\in\mathcal{S}_{s}$ do
+5:     Generate samples $\{\bm{x}_{i}^{(j)}\}_{j=1}^{N_{G}}$ where $\bm{x}_{i}^{(j)}\sim p_{\theta}(\bm{x}\mid s)$
+6:     Compute mean image:
+$\overline{\bm{x}}_{s}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i}^{(j)}$ (32)
+7:     Compute sample variance:
+$\sigma_{s}^{2}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i}^{(j)}-\overline{\bm{x}}_{s}\|^{2}$ (33)
+8:     if $\sigma_{s}^{2}<\tau$ then
+9:         $\mathcal{D}_{\text{extracted}}\leftarrow\mathcal{D}_{\text{extracted}}\cup\{\overline{\bm{x}}_{s}\}$
+10:     end if
+11:end for
+12:return $\mathcal{D}_{\text{extracted}}$
+The Extended SIDE method operates on the principle that random strings, when used as surrogate conditions during fine-tuning, become tightly coupled with specific training images.
+This follows the results of our theoretical understanding in Theorem 1.
+The attacker exploits this behavior to extract training data by performing the following steps:
+  1. 1.
+    Selection of Random Strings: The attacker identifies or generates a set of random strings $\mathcal{S}_{s}$ that were used during the victim’s fine-tuning process.
+These strings act as unique keys that the model associates with specific images.
+  2. 2.
+    Image Generation: For each random string $s\in\mathcal{S}_{s}$, the attacker generates $N_{G}$ images by querying the fine-tuned model:
+    $\bm{x}_{i}^{(j)}\sim p_{\theta}(\bm{x}\mid s),\quad j=1,\dots,N_{G}$ (34)
+    The repetition of image generation for the same string allows the attacker to assess the consistency of the outputs.
+  3. 3.
+    Consistency Analysis: The attacker computes the mean image $\overline{\bm{x}}_{s}$ and the sample variance $\sigma_{s}^{2}$ for the generated images associated with each string $s$:
+    $\overline{\bm{x}}_{s}$ $=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i}^{(j)}$ (35)
+    $\sigma_{s}^{2}$ $=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i}^{(j)}-\overline{\bm{x}}_{s}\|^{2}$ (36)
+    A low variance $\sigma_{s}^{2}<\tau$ indicates that the generated images are highly similar, suggesting that the model has memorized the association between the string $s$ and a specific training image $\bm{x}_{s}$.
+  4. 4.
+    Data Extraction: If the variance $\sigma_{s}^{2}$ is below the threshold $\tau$, the attacker considers $\overline{\bm{x}}_{s}$ as an extracted training image and adds it to the dataset $\mathcal{D}_{\text{extracted}}$:
+    $\mathcal{D}_{\text{extracted}}\leftarrow\mathcal{D}_{\text{extracted}}\cup\{\overline{\bm{x}}_{s}\}$ (37)
+This method effectively leverages the deterministic mapping between random strings and training images induced by memorization, allowing the attacker to reconstruct the training dataset without direct access.
+G.3 Algorithmic Implementation
+The Extended SIDE method is formalized in Algorithm 2.
+The algorithm outlines the steps the attacker takes to generate images, compute the variance, and extract the training data based on the consistency of generated images.
+Since the attacker cannot compute similarity to the training data, they detect memorization by analyzing the diversity of generated images conditioned on the same random string $s$.
+The attacker identifies memorization by detecting low variance in the images generated from the same random string $s$.
+A low variance $\sigma_{s}^{2}$ indicates that the model consistently generates similar images for $s$, suggesting that it has memorized specific training data associated with $s$.
+The Extended SIDE is shown in Algorithm 2
+G.4 Experiments
+To verify the effectiveness of the Extended SIDE, we conduct experiments on the subset of the Laion-2B dataset, finetuning the Stable Diffusion v1.5 using the LORA[18] method.
+We use the following metrics to evaluate the attack results: mean SSCD (M-SSCD), AMS (mid-similarity) and LIPIPS.
+The results are shown in Tab.
+4.
+We use the generated images of original prompts as the Baseline method.
+The extracted examples are shown in Figure 8.
+As the results show, our method could effectively extract training data from fine-tuned DPMs.
+Table 4: Performance Comparison for Extended SIDE method
+Dataset Method M-SSCD AMS LIPIPS
+Subset of Laion-2B Baseline 0.028 0.000 0.892
+Extended SIDE(ours) 0.467 0.672 0.809
+As presented in Tab. 4, the Extended SIDE method outperforms the baseline across all evaluation metrics:
+  • M-SSCD: The SIDE method achieves an M-SSCD of 0.467 compared to the baseline’s 0.028.
+This significant increase indicates that SIDE extracts images with higher structural similarity to the training data, demonstrating effective memorization.
+  • AMS: SIDE records an AMS of 0.672, markedly higher than the baseline’s 0.000.
+This suggests that SIDE successfully captures a substantial number of moderately similar images, enhancing the diversity and relevance of the extracted data.
+  • LIPIPS: SIDE exhibits a lower LIPIPS of 0.809 compared to the baseline’s 0.892.
+The difference is less pronounced.
+This is because the primary purpose of LIPIPS is to evaluate perceptual differences between images, rather than structural or training-specific similarity.
+It measures how humans perceive image similarity based on deep features from pre-trained networks, making it effective for detecting perceptual distortions or quality changes but not for identifying whether an image was directly memorized from the training data.
+Visual examples Figure 8 showcases examples of extracted images using the SIDE method alongside the corresponding training images.
+The visual comparison underscores the high degree of similarity achieved by SIDE, affirming its efficacy in extracting memorized training data.
+Figure 8: Examples of Extracted Images using SIDE Method.
+Top Row: Corresponding training images from the Laion-2B subset.
+Bottom Row: Images generated by querying the fine-tuned Stable Diffusion model with random strings using the SIDE method.
+The high visual similarity demonstrates the method’s effectiveness in extracting training data.
+The Extended SIDE method effectively extracts training data from fine-tuned DPMs by exploiting the memorization induced by random strings during the fine-tuning process.
+The quantitative results demonstrate that SIDE significantly enhances the extracted images’ structural and mid-range similarity compared to the baseline.
+This indicates that SIDE can reliably reproduce specific training images by leveraging the unique associations between random strings and their corresponding images.

### v6

@@
 Abstract
-As diffusion probabilistic models (DPMs) are being employed as mainstream models for Generative Artificial Intelligence (GenAI), the study of their memorization has attracted growing attention.
-Existing works in this field aim to establish an understanding of whether or to what extent DPMs learn via memorization.
-Such an understanding is crucial for identifying potential risks of data leakage and copyright infringement in diffusion models and, more importantly, for trustworthy application of GenAI.
-Existing works revealed that conditional DPMs are more prone to memorize training data than unconditional DPMs.
-And most data extraction methods developed so far target conditional DPMs.
-Although unconditional DPMs are less prone to data extraction, further investigation into these attacks remains essential since they serve as the foundation for conditional models like Stable Diffusion, and exploring these attacks will enhance our understanding of memorization in DPMs.
-In this work, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) that leverages a time-dependent classifier trained on generated data as surrogate conditions to extract training data from unconditional DPMs.
-Empirical results demonstrate that it can extract training data in challenging scenarios where previous methods fail, and it is, on average, over 50% more effective across different scales of the CelebA dataset.
-Furthermore, we provide a theoretical understanding of memorization in both conditional and unconditional DPMs and why SIDE is effective.
+As diffusion probabilistic models (DPMs) become the mainstream framework for Generative AI (GenAI), understanding their memorization behavior is critical for assessing risks of data leakage, copyright infringement, and ensuring trustworthy deployment.
+Previous studies indicate that conditional DPMs are more prone to memorization compared to unconditional DPMs, with the majority of extraction techniques focusing on conditional models.
+However, studying unconditional DPMs remains crucial, as they serve as the foundation for advanced models such as Stable Diffusion.
+Understanding their memorization behavior can provide valuable insights into the broader characteristics of DPMs.
+In this work, we propose Surrogate condItional Data Extraction (SIDE), a novel data extraction method for unconditional DPMs.
+SIDE employs a time-dependent classifier trained on generated data as surrogate conditions to extract training samples.
+Empirical results show that SIDE extracts training data effectively in challenging scenarios where existing methods fail, with an average performance gain of over 50% on the CelebA dataset across different scales.
+We also provide a unified theoretical framework explaining memorization in both conditional and unconditional DPMs and why SIDE works.
+Our findings contribute to a deeper understanding of memorization in GenAI and provide practical guidance for mitigating associated risks.
 1 Introduction
-The diffusion probabilistic models (DPMs) [16, 30, 34] is a family of powerful generative models that learn the distribution of a dataset by first gradually destroying the structure of the data through an iterative forward diffusion process and then restoring the data structure via a reverse diffusion process.
-Due to their outstanding capability in capturing data distribution, DPMs have become the foundation models for many pioneering Generative Artificial Intelligence (GenAI) products such as Stable Diffusion [26], DALL-E 3 [2], and Sora [3].
-Despite the widespread adoption of DPMs, a potential risk they face is data memorization, i.e., the risk of memorizing a certain proportion of the raw training samples.
-This could result in the generation of memorized (rather than new) samples via direct copying, causing data leakage, privacy breaches or copyright infringement, as highlighted in the literature [31, 32, 1, 11].
-A current case argues that Stable Diffusion is a 21st-century collage tool, remixing the copyrighted creations of countless artists whose works were included in the training data[4].
-Furthermore, data memorization also gives rise to data extraction attacks, which is one type of privacy attacks that attempt to extract the training data from a well-trained model.
-Notably, recent work by [5] demonstrated the feasibility of extracting training data samples from DPMs like Stable Diffusion [26], revealing the potential privacy and copyright dangers associated with them.
-Several works have investigated the data memorization phenomenon in diffusion models.
-For example, it has been observed that there exists a strong correlation between training data memorization and conditional DPMs, i.e., conditional-trained DPMs are more prone to memorization [32].
-Gu et al.
-[13] investigates influential factors on memorization behaviors via comprehensive experiments.
-They found that conditioning on random-labeled data can significantly trigger memorization, and unconditional models memorize much less training data.
-Though these works indicate that unconditional diffusion models are much less prone to data extraction attacks, investigating such attacks for unconditional diffusion models is still necessary.
-This is because 1) unconditional DPMs are the foundations of conditional generative models like Stable Diffusion, which adapts pre-trained unconditional DPMs as the essential part for conditional generation or combines them with conditional training approaches [26, 39]; 2) exploring data extraction attacks on unconditional DPMs enhances understanding of memorization mechanisms in both conditional and unconditional DPMs.
-In this work, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) to extract training data from unconditional DPMs.
-This attack method uses a classifier trained on the same or similar training data as a surrogate condition for unconditional DPMs.
-SIDE can extract training data effectively by making the classifier time-dependent (to the diffusion sampling process).
-We empirically verify the effectiveness of SIDE on CIFAR-10 and different scales of the CelebA dataset (attack results in Figure 1).
-Additionally, to offer a theoretical analysis, we introduce a memorization metric to quantify the degree of memorization in DPMs by measuring the overlap between the generated and training data in a point-wise manner.
-Based on this metric, we explain why conditional DPMs are more likely to memorize training data, why random labelling can increase this memorization and why the SIDE method is effective for extracting training data in unconditional DPMs.
+Diffusion probabilistic models (DPMs) [21, 38, 42] are a family of powerful generative models that learn data distributions by progressively corrupting data through a forward diffusion process and then reconstructing it via a reverse diffusion process.
+Due to their exceptional capability to capture complex data distributions, DPMs have become foundation models for many pioneering Generative Artificial Intelligence (GenAI) products, including Stable Diffusion [33], DALL-E 3 [4], and Sora [5].
+However, the growing adoption of DPMs raises concerns about data memorization — the risk of memorizing raw training samples.
+This may result in the generation of duplicated rather than novel content, increasing the risk of data leakage, privacy violations, and copyright infringement.
+[39, 40, 2, 13].
+For instance, Stable Diffusion has been criticized as a “21st-century collage tool” that remixes copyrighted works from artists whose data was used for training [6].
+Moreover, data memorization enables data extraction attacks, a type of privacy attack aimed at recovering training data from trained models.
+Recent work by [7] demonstrated the feasibility of extracting training data from DPMs like Stable Diffusion [33], highlighting significant privacy and copyright risks.
+Existing studies have explored data memorization in DPMs, revealing that conditional DPMs exhibit a higher tendency to memorize training data than unconditional DPMs [40].
+For example, [18] found that conditioning on random-labeled data significantly increases memorization, while unconditional models exhibit much lower memorization rates.
+Despite these findings, investigating data extraction attacks on unconditional DPMs remains crucial for two reasons: (1) unconditional DPMs form the foundation of conditional models like Stable Diffusion, which adapt pre-trained unconditional DPMs for conditional generation [33, 48]; and (2) understanding memorization in unconditional DPMs provides insights into the broader memorization mechanisms of DPMs.
+In this work, we propose Surrogate condItional Data Extraction (SIDE), a novel method for extracting training data from unconditional DPMs.
+SIDE employs a time-dependent classifier trained on generated data as surrogate conditions to effectively extract training samples.
+We empirically validate SIDE on CIFAR-10, multiple scales of the CelebA and ImageNet, demonstrating its superiority over baseline methods by achieving over 50% higher effectiveness on average.
+Examples of extracted images are shown in Figure 1.
+Additionally, we introduce a divergence to measure memorization in DPMs and present a theoretical analysis explaining: (1) why conditional DPMs memorize more data (even labelled with random labels), (2) why SIDE is effective for extracting training data.
 In summary, our main contributions are:
-  • We propose a novel training data extraction method SIDE that leverages the implicit labels learned by a time-dependent classifier to extract training data from unconditional DPMs.
-  • We introduce a novel metric to measure the degree of point-wise memorization in DPMs and present a theoretical analysis that explains 1) why conditional DPMs are more prone to memorization, 2) why random labels can lead to more memorization, and 3) why the SIDE method is effective for extracting training data in unconditional DPMs.
-  • We evaluate the effectiveness of SIDE on CIFAR-10 and various scales of CelebA datasets and show that, on average, it can outperform the baseline method by more than 50%.
-Figure 1: A few examples of the extracted images from a DDPM trained on a subset of the CelebA dataset using our SIDE method. Top: training images; bottom: extracted images.
+  • We propose SIDE, a novel method that leverages implicit labels provided by a time-dependent classifier to extract training data from unconditional DPMs.
+  • We introduce a memorization divergence measure and provide a theoretical analysis that explains the impact of conditioning in DPMs and the effectiveness of SIDE.
+  • We empirically validate the effectiveness of SIDE on CIFAR-10, CelebA and ImageNet, demonstrating it outperforms baseline methods by over 50% on average.
+Figure 1: Examples of training images (top) and corresponding extracted images (bottom) from a DDPM trained on a subset of CelebA using our SIDE method.
 2 Related Work
-Diffusion Probabilistic Models  DPMs [30] such as Stable Diffusion [26], DALL-E 3 [2], Sora [3], Runway [26], and Imagen [28] have achieved state-of-the-art performance in image/video generation across a wide range of benchmarks [12].
-These models can be viewed from two perspectives.
-The first is score matching [34], where diffusion models learn the gradient of the image distribution [35].
-The second perspective involves denoising DPMs [16], which add Gaussian noise at various time steps to clean images and train models to denoise them.
-To conditionally sample from diffusion models, [12] utilizes a classifier to guide the denoising process at each sampling step.
-Additionally, [17] introduces classifier-free guidance for conditional data sampling using DPMs.
-Memorization in Diffusion Models  Early exploration of memorization in large models was focused on language models [6, 19], which has motivated more in-depth research on DPMs [32, 13, 24].
-Notably, a recent research on image-generation DPMs: [31] found that 0.5-2% of generated images duplicate training samples, a result concurrently reported by [5] in broader experiments for both conditional and unconditional diffusion models.
-Further studies [32] and [13] linked memorization to model conditioning, showing that conditional models are more prone to memorization.
-Several works have proposed methods to detect and mitigate memorization to address this issue.
-Wen et al.
-[38] proposed a method to detect whether a prompt can trigger memorization by inspecting the magnitude of text-conditional predictions.
-It delivers high accuracy even at the first generation step, with a single generation per prompt to detect whether a prompt can trigger memorization.
-Ren et al.
-[25] proposed two metrics based on findings on cross-attention in DPMs to detect memorization.
-To mitigate memorization, Chen et al.[7] proposed anti-memorization guidance to mitigate memorization during the sampling process.
-Instead of using guidance, Ren et al.
-[25] mitigated memorization by changing the attention score or masking out summary tokens in the cross-attention layer.
-Wen et al.[38] achieved this by minimizing the magnitude during inference.
-So far, the motivated data extraction attack or defend methods are primarily for conditional DPMs [5, 37], and studies have shown that extracting training data from unconditional DPMs can be much more challenging than that on conditional DPMs [13].
-Though challenging, investigating data extraction methods in unconditional DPMs is still necessary.
-This is because 1) Unconditional DPMs are key parts in conditional models like Stable Diffusion, which adapt them for conditional generation, and 2) Exploring this approach enhances our understanding of memorization in DPMs.
-In response to these challenges, we propose a novel data extraction method called SIDE for unconditional DPMs and providing theoretical understanding of SIDE.
-3 Proposed SIDE Method
-3.1 Intuition
-Previous studies [13, 32] have shown that conditional DPMs are more prone to memorization due to their reliance on explicit conditioning labels, such as class labels or text conditions, which we refer to as explicit informative labels or explicit labels.
-In contrast, unconditional DPMs do not have access to these explicit labels.
-Nevertheless, they still learn to form representation clusters from the training data, even though these clusters are not directly specified in the data itself [8].
-We name these representation clusters as implicit informative labels or implicit labels.
-By formalizing the clustering patterns within the training data, it may be possible to construct implicit labels and make the unconditional DPMs conditional on these implicit labels to extract training data from unconditional DPMs.
-In the following section, we will construct implicit informative labels for unconditional DPMs, convert the implicit labels into explicit ones, and then leverage the explicit labels to extract training data.
-3.2 Constructing Implicit Informative Labels
-Intuitively, one could use a classifier to generate (predict) implicit labels $y_{I}$ during the sampling process of the diffusion model.
-The classifier can be a normal classifier trained on the same dataset as the diffusion model.
-When such a classifier is not available, random labels or representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can also be used as the implicit labels.
-We assume an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process:
+Diffusion Probabilistic Models
+DPMs [38] have achieved state-of-the-art performance in image and video generation, as exemplified by models such as Stable Diffusion [33], DALL-E 3 [4], Sora [5], Runway [33] and Imagen [35].
+These models excel on various benchmarks [15].
+DPMs can be interpreted from two perspectives: 1) score matching [42], where model learns the gradient of data distribution [43], and (2) denoising diffusion [21], where Gaussian noise is added to clean images over multiple time steps, and the model is trained to reverse this process.
+For conditional sampling, [15] introduced classifier guidance to steer the denoising process, while [22] proposed classifier-free guidance, enabling conditional generation without explicit classifiers.
+Memorization in Diffusion Models
+Early research on memorization primarily focused on language models [8, 25], which later inspired subsequent studies on DPMs [40, 18, 31, 36, 16, 1, 16, 27, 47, 3, 17].
+Notably, [39] found that 0.5-2% of generated images duplicate training samples, a result corroborated by [7] through more extensive experiments on both conditional and unconditional DPMs.
+Further studies [40, 18] linked memorization to model conditioning, showing that conditional DPMs are more prone to memorization.
+To address memorization, several methods have been proposed for detection and mitigation.
+For example, [46] introduced a method to detect memorization-triggering prompts by analyzing the magnitude of text-conditional predictions, achieving high accuracy with minimal computational overhead.
+[32] proposed metrics based on cross-attention patterns in DPMs to identify memorization.
+On the mitigation side, [9] developed anti-memorization guidance to reduce memorization during sampling, while [32] modified attention scores or masked summary tokens in the cross-attention layer.
+[46] minimized memorization by controlling prediction magnitudes during inference.
+Despite these advances, the current research mostly focuses on conditional DPMs [7, 45], as extracting data from unconditional DPMs has been found to be extremely more challenging [18].
+However, investigating unconditional DPMs remains critical because 1) they form the foundation of conditional models like Stable Diffusion, which adapt them for conditional generation, and 2) understanding memorization in unconditional DPMs provides insights into the broader behavior of DPMs.
+In response to these challenges, we propose SIDE, a novel data extraction method for unconditional DPMs, and provide a unified theoretical framework to explain memorization in both conditional and unconditional DPMs.
+3 Surrogate Conditional Data Extraction
+Threat Model
+In this work, we focus on unconditional DPMs and assume that the attacker has access to the target DPM as well as a classifier trained on the same or similar dataset.
+The attacker’s objective is to extract the original training samples from the target DPM.
+In cases where the classifier is unavailable, the attacker can leverage generated data to train one, following a standard training procedure, potentially with certain necessary annotations.
+It is important to note that the classifier does not leak training data, as we will empirically demonstrate in Sec.
+5.3.
+Its sole purpose, in this context, is to generate implicit labels, which can alternatively be achieved through clustering techniques.
+3.1 Intuition of SIDE
+Previous research [18, 40] have shown that conditional DPMs are more susceptible to memorization due to their reliance on explicit labels (e.g., class labels or text conditions).
+In contrast, unconditional DPMs lack access to these explicit labels.
+However, they still learn to form representation clusters from the training data, even though these clusters are not explicitly defined [10].
+We term these representation clusters as implicit labels.
+By formalizing the clustering patterns within the training data, it becomes possible to construct implicit labels and condition unconditional DPMs on these labels to facilitate training data extraction.
+In the following, we detail the construction of implicit labels for unconditional DPMs, their conversion into explicit labels, and the subsequent use of these explicit labels for training data extraction.
+3.2 Constructing Implicit Labels
+Intuitively, implicit labels can be generated during the sampling process of the diffusion model using a classifier.
+This classifier can be trained on the same/similar dataset as the diffusion model.
+When such a classifier is unavailable, representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can serve as implicit labels.
+We assume that an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process described as:
 $\mathrm{d}{\bm{x}}=$ $\Big{[}f({\bm{x}},t)-g(t)^{2}\Big{(}\nabla_{\bm{x}}\log p_{\theta}^{t}({\bm{x}})+$
-$\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\Big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w$ (1)
-where $x$ represents the state vector, $f(x,t)$ denotes the drift coefficient, $g(t)$ is the diffusion coefficient, $\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$ denotes the gradient of model $p_{\theta}$ given ${\bm{x}}$ at time $t$, and $\mathrm{d}w$ corresponds to the increment of the Wiener process.
-We can use the classifier that generates the implicit labels to approximate the gradient in Eq.
-(3.2).
-However, a known challenge associated with neural network classifiers is their tendency towards miscalibration [14].
-Specifically, the classifier could be overconfident or underconfident about its predictions.
-To mitigate the potential impact of miscalibration on the sampling process, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output on the diffusion path using power prior as follows:
-$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right)$ (2)
-Then, we have:
+$\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\Big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w,$ (1)
+where ${\bm{x}}$ represents the state vector, $f(x,t)$ denotes the drift coefficient, $g(t)$ is the diffusion coefficient, $\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$ denotes the gradient of model $p_{\theta}$ given ${\bm{x}}$ at time $t$, and $\mathrm{d}w$ corresponds to the increment of the Wiener process.
+To approximate the gradient in the sampling process, we use the classifier that generates the implicit labels.
+However, neural network classifiers are prone to miscalibration [19], which can affect the sampling process.
+To address this, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output using a power prior:
+$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right).$ (2)
+This leads to a classifier-conditioned sampling process:
 $\mathrm{d}{\bm{x}}=$ $\Big{[}f({\bm{x}},t)-g(t)^{2}\big{(}\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$
-$+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w$ (3)
-[…]
-Figure 2: An illustration of our proposed Time-Dependent Knowledge Distillation (TDKD) method that trains a time-dependent classifier on a pseudo-labeled synthetic dataset.
-3.3 Time-dependent Classifier
+$+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w.$ (3)
+This approach differs from [15], which introduced classifier guidance for improving sample quality under the assumption that with $Z$ being a constant.
+However, this assumption only holds when $\lambda=1$, as $Z$ is explicitly dependent on the ${\bm{x}}_{t}$ (the $t$-th step of sampling image ${\bm{x}}$ ) when $\lambda\neq 1$.
+Our derivation solves this issue by redefining $p^{t}_{\theta}\left(y|{\bm{x}}\right)$ using the power prior.
+Figure 2: An illustration of our proposed time-dependent classifier training method on a pseudo-labeled synthetic dataset.
+3.3 Training Time-dependent Classifier
 In Eq.
-(3.2), the classifier is denoted by $p_{\theta}^{t}(y|{\bm{x}})$ with $t$ being the timestep, implying its time-dependent nature.
-However, we do not have a time-dependent classifier but only a time-independent classifier.
-To address this problem, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier.
+(3.2) above, the classifier is represented as $p_{\theta}^{t}(y|{\bm{x}})$, where $t$ denotes the timestep, indicating its time-dependent nature.
+However, we only have access to a time-dependent classifier.
+To address this limitation, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier based on synthetic data generated at different sampling steps.
 The distillation process is illustrated in Figure 2.
-TDKD equips the classifier with time-dependent guidance during sampling.
-It operates in two steps: first, the network architecture is adjusted to accommodate time-dependent inputs, and the structure of the time-dependent module and modification are illustrated in Appendix C; second, a synthetic dataset and pseudo labels are created to distil knowledge from the normal classifier to its time-dependent counterpart.
-As the original training dataset is unknown, we employ the target DPM to generate a synthetic dataset, following the generative data augmentation techniques [10, 9].
-We then use the normal classifier trained on the original dataset to generate pseudo labels for the synthetic dataset.
-Finally, we train a time-dependent classifier on the labelled synthetic dataset.
-The objective of this training is defined as follows:
+TDKD enables the classifier to provide time-dependent guidance during sampling through a two-step process:
+Step 1: The network architecture is modified to incorporate time-dependent inputs. The structure of the time-dependent module and the specific modifications are detailed in Appendix B.
+Step 2: A synthetic dataset with pseudo labels is generated to distill knowledge from the time-independent classifier to its time-dependent counterpart.
+At Step 2, we can leverage the target DPM to generate a synthetic dataset, following established generative data augmentation techniques [12, 11].
+Pseudo labels for the synthetic dataset are then generated using the time-independent classifier trained on the original dataset.
+Finally, the time-dependent classifier is trained on the labeled synthetic dataset.
+The training objective is defined as:
 $\mathcal{L}_{distill}=D_{KL}\left(p_{\theta}(y_{I}|{\bm{x}}),p_{\theta}^{t}(y_{I}|{\bm{x}}_{t})\right)$ (4)
-Overall Procedure of SIDE
-With the trained time-dependent classifier $p_{\theta}^{t}(y|{\bm{x}}_{t})$ and the target DPM, our SIDE extracts training data from the model following a conditional generation process.
-Assume we condition on the label $y=c$.
-Firstly, we pick a set of values for $\lambda$, called $\mathcal{S}_{\lambda}$, to use in the SIDE attack.
-Then, we gather $N_{G}$ data samples for each value of $\lambda$ from the set.
-For each sampling timestep $t$, we compute the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ ($CE(\cdot)$ denotes the cross-entropy loss) and use the gradient and the target DPM to reverse the diffusion process.
-Next, we compute the similarity score for each generated image.
-Finally, we evaluate the attack performance using evaluation metrics.
-Defer to Appendix E for SIDE’s pseudocode.
+3.4 Overall Procedure of SIDE
+With the trained time-dependent classifier $p_{\theta}^{t}(y|{\bm{x}}_{t})$ and the target DPM, our SIDE method extracts training data from the model through a conditional generation process, as described in Algorithm 1.
+Given the label $y=c$, the complete procedure of SIDE is as follows.
+It begins by selecting a set of values for $\lambda$, denoted as $\mathcal{S}_{\lambda}$ (Line 1).
+For each $\lambda$, the target DPM is required to generate $N_{G}$ data samples (Lines 18 - 27).
+Then, at each sampling timestep $t$, the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ (Line 22) is computed, where $CE(\cdot)$ is the cross-entropy loss.
+This gradient, together with the target DPM, is then used to guide the reverse diffusion process.
+For each generated image, a similarity score is computed against every ground-truth training images to identify the most closely matching original image.
+Using the similarity scores obtained for the extracted images, a performance metric can then be calculated to evaluate the effectiveness of the extraction process.
+Existing methods predominantly rely on $L_{p}$ distance metrics to identify pixel-wise similar samples.
+However, such metrics are insufficient for capturing semantically meaningful similarities.
+In the experiment section, we will introduce a set of performance metrics to facilitate a more comprehensive evaluation of the extraction methods.
+Algorithm 1 SIDE
+1:Input: DPM $p_{\theta}(\bm{x})$, classifier $p_{\theta}(y_{I}|\bm{x})$, $\mathcal{S}_{\lambda}$, $N_{G}$, label $c$, $N_{syn}$, $T$, Denoise Step $DS(\bm{x}_{t},t,\bm{s})$
+2:Output: Extracted data $\mathcal{D}_{ext}$, classifier $p_{\theta}^{t}(y_{I}|\bm{x}_{t})$
+3:Classifier Training:
+4:Step 1: Architecture
+5:Modify $p_{\theta}(y|\bm{x})$ to $p_{\theta}^{t}(y|\bm{x}_{t})$ for time input
+6:Step 2: Distillation
+7:for $i=1$ to $N_{syn}$ do $\triangleright$ Generate synthetic data
+8:    $\bm{x}_{T}\sim\mathcal{N}(0,I)$
+9:    for $t=T$ down to $1$ do
+10:         $\bm{x}_{t-1}\leftarrow DS(\bm{x}_{t},t,s_{\theta}(\bm{x}_{t},t))$
+11:    end for
+12:    $y_{I}^{(i)}\leftarrow p_{\theta}(y_{I}|\bm{x}_{0}^{(i)})$triangleright Generate pseudo-labels
+13:end for
+14:Step 3: Train $p_{\theta}^{t}(y_{I}|\bm{x}_{t})$:
+15:$\mathcal{L}_{distill}=\frac{1}{N_{syn}}\sum_{i=1}^{N_{syn}}D_{KL}(p_{\theta}(y_{I}|\bm{x}_{0}^{(i)}),p_{\theta}^{t}(y_{I}|\bm{x}_{t}^{(i)}))$
+16:Data Extraction:
+17:Initialize $\mathcal{D}_{ext}\leftarrow\emptyset$
+18:for $\lambda\in\mathcal{S}_{\lambda}$ do
+19:    for $i=1$ to $N_{G}$ do
+20:         $\bm{x}_{T}\sim\mathcal{N}(0,I)$
+21:         for $t=T$ down to $1$ do
+22:             $g_{t}\leftarrow\nabla_{\bm{x}_{t}}CE(c,p_{\theta}^{t}(y_{I}|\bm{x}_{t}))$
+23:             $s_{\text{guided}}\leftarrow\nabla_{\bm{x}_{t}}\log p_{\theta}^{t}(\bm{x}_{t})-\lambda g_{t}$
+24:             $\bm{x}_{t-1}\leftarrow DS(\bm{x}_{t},t,s_{\text{guided}})$
+25:         end for
+26:         Append $\bm{x}_{0}$ to $\mathcal{D}_{ext}$
+27:    end for
+28:end for
+29:return $\mathcal{D}_{ext}$, $p_{\theta}^{t}(y_{I}|\bm{x}_{t})$
+4 Theoretical Analysis
+In this section, we first introduce a Kullback-Leibler (KL) divergence-based distribution measure to quantify the degree of memorization in generative models.
+Building on this measure, we provide a theoretical explanation for data memorization in conditional diffusion probabilistic models (DPMs) and elucidate why SIDE is effective for unconditional DPMs.
+4.1 Distributional Memorization Measure
+There are various approaches to measure the memorization effect of a generative model.
+One approach involves comparing each generated sample with the raw training samples in a sample-by-sample manner, as previously mentioned.
+While such measures (e.g., $L_{p}$ distances) are effective for evaluating data extraction performance, they are insufficient for analyzing the overall memorization effect of the model.
+To characterize the model-level memorization effect with respect to the training data distribution and to support our theoretical analysis, we introduce the following distributional memorization measure.
+To assess the model’s memorization ability with respect to the overall data distribution while also accounting for the memorization of each individual data point, it is natural to consider the Kullback-Leibler (KL) divergence between the uniform empirical distribution over the dataset $\mathcal{D}$, defined as $\frac{1}{|\mathcal{D}|}\sum_{x_{i}\in\mathcal{D}}\delta(x_{i})$ with $\delta(\cdot)$ is the Dirac delta function, and the distribution $p$ induced by the model’s generated samples.
+Here, the $\delta(\cdot)$ function enforces a point-wise memorization measure, which quantifies how closely the generated samples align with each individual data point in the original distribution.
+A smaller KL divergence indicates that the model exhibits stronger memorization of the training data points.
+However, for a continuous probability distribution $p$, computing such a KL divergence directly is infeasible.
+To address this, we approximate each Dirac delta function by a normal distribution with a small variance in the following definition.
+Definition 1 (Memorization Divergence).
+Given a generative model $p_{\theta}$ with parameters $\theta$ and training dataset $\mathcal{D}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, the degree of divergence between $p_{\theta}$ and distribution of training dataset is defined as:
+$\mathcal{M}(\mathcal{D};p_{\theta},\epsilon)$ $=D_{\mathrm{KL}}(q_{\epsilon}\|p_{\theta})$ (5)
+with $\text{with }q_{\epsilon}(x)=$ $\frac{1}{N}\sum_{x_{i}\in\mathcal{D}}\mathcal{N}(x|x_{i},\epsilon^{2}I),$
+where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ denotes the $i$-th training sample, $N$ is the total number of training samples, $p_{\theta}({\bm{x}})$ represents the probability density function (PDF) of the generated samples, and $\mathcal{N}(x|x_{i},\epsilon^{2}I)$ is the normal distribution with mean ${\bm{x}}_{i}$ and covariance matrix $\epsilon^{2}I$.
+Note that in Eq.
+(5), a smaller value of ${\mathcal{M}}({\mathcal{D}};p_{\theta},\epsilon)$ indicates greater overlap between the two distributions, reflecting stronger memorization.
+As $\epsilon$ approaches 00, the measured memorization divergence becomes more accurate.
+In fact, the normal distribution $\mathcal{N}(x|x_{i},\epsilon^{2}I)$ can be replaced by any family of continuous distributions $\hat{q}(x|x_{j},\epsilon)$ that satisfies two conditions: 1) symmetric with respect to $x$ and $x_{j}$(i.e., $\hat{q}(x|x_{j},\epsilon)=\hat{q}(x_{j}|x,\epsilon)$) and 2) convergence to $\delta(x_{i})$ in distribution.
+This substitution does not affect the result of Theorem 1 in Sec.
+4.2.
+While concerns may arise regarding the influence of $\epsilon$ on the divergence, this divergence is primarily intended for comparative analysis when $\epsilon$ is sufficiently small.
+When comparing the memorization divergence across different models, the $\epsilon$ will not influence relative differences, as demonstrated in Appendix A.
+4.2 Theoretical Analysis
+Building on the memorization divergence measure, here we provide a theoretical analysis to explain why conditional DPMs exhibit stronger memorization effect.
+Our analysis centers on the concept of informative labels, which represent information capable of partitioning a dataset into multiple disjoint subsets.
+We show that DPMs, conditioned on informative labels, exhibit stronger memorization effects.
+Informative Labels
+The concept of informative labels has been previously discussed as class labels [18].
+In this work, we extend this notion by introducing a more general definition that encompasses includes both class labels and random labels as special cases.
+Formally, we define an informative label as follows.
+Definition 2 (Informative Label).
+Let ${\mathcal{Y}}$ be an attribute of the data that takes values in $\{y_{i}\}_{i=1}^{C}$.
+We define ${\mathcal{Y}}$ as an informative label if it enables the partitioning of the dataset into several mutually disjoint subsets $\{{\mathcal{D}}_{i}\}_{i=1}^{C}$, where each subset corresponds to a distinct value of ${\mathcal{Y}}$.
+In the above definition, informative labels are not restricted to conventional class labels; they can also include text captions, features, or cluster information that can be used to group training samples into subsets.
+The definition emphasizes that an informative label must be capable of distinguishing a subset of samples from others.
+An extreme case occurs when all samples share the same label, rendering the label non-informative.
+According to our definition, both class-wise and random labels are special cases of informative labels.
+Informative labels can be either explicit, such as class labels, random labels, or text captions, or implicit, such as salient features or clusters.
+Next, we present our main theoretical result on the memorization mechanism of conditional DPMs and explain why SIDE is effective.
+Let ${\mathcal{D}}_{i}$ denote the subset consisting of data with the informative label ${\mathcal{Y}}=y_{i}$.
+Additionally, we denote the data distribution in the original dataset ${\mathcal{D}}$ by $p$.
+Similarly, for each attribute $y_{i}$, we use $p_{i}$ to denote the data distribution corresponding to that attribute.
+Theorem 1.
+If a generative model $p_{\theta_{i}}$ matches the target distribution $p_{i}$ almost everywhere for the informative label $y_{i}$, i.e., $TV(p_{i},p_{\theta_{i}})=0$, then with probability 1:
+$\lim_{\epsilon\to 0}\lim_{|{\mathcal{D}}_{i}|\to\infty}\left({\mathcal{M}}({\mathcal{D}}_{i};p_{\theta_{i}},\epsilon)-{\mathcal{M}}({\mathcal{D}}_{i};p_{\theta},\epsilon)\right)\leq 0,$ (6)
+where $TV(\cdot)$ is the total variance distance, $p_{\theta_{i}}$ and $p_{\theta}$ denote the distribution of generated data for model trained on data labeled $y_{i}$ and the whole dataset.
+The equality holds if and only if $TV(p,p_{i})=0$.
+The proof for Theorem 1 is in Appendix A.
+The theorem implies that any form of informative labels, including class labels and random labels, can enhance the memorization effect of conditional generative models.
+Conditional models exhibit stronger memorization by isolating data subsets through explicit informative labels, allowing focused training on each group’s unique patterns, which sharpens their ability to memorize specific training examples.
+SIDE succeeds by uncovering and exploiting the unconditional DPM’s inherent but implicit informative labels, using classifier guidance to concentrate generation on specific memorized features.
+Table 1: The AMS (%) and UMS (%) results at low, mid, and high similarity levels.
+‘Random’ is the baseline method proposed in [7], while OL-TI is SIDE without TDKD training.
+We also compare SIDE using classifiers trained on the original (SIDE-O) vs.
+disjoint auxiliary (SIDE-A) datasets.
+The similar performance between SIDE-O and SIDE-A suggests there is no training data leaking via the classifier.
+Dataset Method Low Similarity Mid Similarity High Similarity
+AMS(%) UMS(%) AMS(%) UMS(%) AMS(%) UMS(%)
+CIFAR-10 Random 2.470 1.770 0.910 0.710 0.510 0.420
+OL-TI 2.460 1.780 0.800 0.680 0.420 0.370
+SIDE-A 5.250 2.020 2.300 0.880 1.620 0.640
+SIDE-O 5.320 2.050 2.490 0.860 1.770 0.560
+CelebA-HQ-FI Random 11.656 2.120 0.596 0.328 0.044 0.040
+OL-TI 2.649 0.744 0.075 0.057 0.005 0.005
+SIDE-A 15.010 2.624 1.310 0.554 0.090 0.082
+SIDE-O 15.172 2.342 1.115 0.444 0.054 0.044
+CelebA-25000 Random 5.000 4.240 0.100 0.100 0.000 0.000
+OL-TI 0.164 0.152 0.000 0.000 0.000 0.000
+SIDE-A 8.712 6.802 0.234 0.234 0.010 0.010
+SIDE-O 8.756 6.940 0.224 0.212 0.012 0.012
+CelebA Random 1.953 1.895 0.000 0.000 0.000 0.000
+OL-TI 1.145 1.125 0.000 0.000 0.000 0.000
+SIDE-A 4.682 4.706 0.098 0.098 0.000 0.000
+SIDE-O 4.883 4.766 0.078 0.078 0.000 0.000
+ImageNet Random 0.000 0.000 0.000 0.000 0.000 0.000
+OL-TI 0.000 0.000 0.000 0.000 0.000 0.000
+SIDE-A 0.152 0.152 0.076 0.076 0.000 0.000
+SIDE-O 0.114 0.114 0.038 0.038 0.000 0.000
+Figure 3: A comparison between the original training images (top row) and extracted images (bottom row) by our SIDE method.
+The matches are classified into three categories: low similarity (SSCD score $<$ 0.5), mid similarity (SSCD score between 0.5 and 0.6), and high similarity (SSCD score $>$ 0.6).
+This highlights varying degrees of semantic resemblance among the image pairs.
 5 Experiments
-In this section, we first describe our experimental setup, introduce the performance metrics and then present the main evaluation results of our SIDE method.
-We empirically verify that the memorized images are not from the classifier.
-We also conduct an ablation study and hyper-parameter analysis to help understand the working mechanism of SIDE.
+In this section, we first introduce the experimental setup and performance metrics, followed by the main evaluation results.
+We also show that the memorized images are not derived from the classifier.
+Additionally, we conduct an ablation study and hyperparameter analysis to provide deeper insights into the working mechanism of SIDE.
 5.1 Experimental Setup
-We evaluate the effectiveness of our method on three datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) [23] which consists of 5478 images, 2) a subset of the CelebA (CelebA-25000) [21] which contains 25,000 images and 2) CIFAR-10 containing 50,000 images.
-All the images are normalized to [-1,1].
+We evaluate the effectiveness of our method on five datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) [30], which consists of 5478 images with a resolution of 128$\times$128; 2) a subset of the CelebA (CelebA-25000) [28], containing 25,000 images with a resolution of 128$\times$128; 3) CIFAR-10, containing 50,000 images with a resolution of 32$\times$32; 4) CelebA [28], which contains 22059 images with a resolution of 128$\times$128; and 5) ImageNet [14], which contains $\approx$ 14 million images with a resolution of 256$\times$256.
+All the images are normalized to the range of [-1,1].
 We use the AdamW optimizer [29] with a learning rate of 1e-4 to train the time-dependent classifier.
@@
 All DPMs are trained with batch size 64.
-We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, respectively.
-We use ResNet34 [15] as the normal classifier.
-4.2 Performance Metrics
-Determining whether a generated image is a memorized copy of a training image is difficult, as $L_{p}$ distances in the pixel space are ineffective.
-Previous research addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection [31, 13].
+We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and CelebA, 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, and 1980K steps on ImageNet, respectively.
+We use ResNet34 [20] as the pseudo-labeler.
+For ImageNet, the DPM is trained on the entire dataset.
+However, due to the scale of ImageNet, we only evaluate on a smaller subset known as ImageNette [23].
+5.2 Image-level Performance Metrics
+As previously mentioned, determining whether an extracted image is a memorized copy of a training image is challenging, and $L_{p}$ distances in pixel space are ineffective when the image is semantically similar but not an exact copy.
+Prior research [39, 18] addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection.
 However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization.
-Here, we propose two new memorization scores to solve these issues: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS).
-The AMS and UMS metrics are defined as follows:
-$AMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{1},\alpha,\beta\right)}}{N_{G}}$ (5)
-$UMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)|}{N_{G}}$ (6)
+To address these issues, we propose two new memorization scores: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS).
+The AMS and UMS scores are defined as follows:
+$AMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)}}{N_{G}}$ (7)
+$UMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)|}{N_{G}},$ (8)
 where ${{\mathcal{D}}}_{1}$ is the generated dataset, ${\mathcal{D}}_{2}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring.
@@
 $\phi({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria.
-Mathematically, we can represent:
+Mathematically, they can be formulated as:
 $\mathcal{F}(x_{i},{\mathcal{D}}_{2},\alpha,\beta)=\mathds{1}\Big{[}\max_{x_{j}\in{\mathcal{D}}_{2}}\gamma(x_{i},x_{j})\in[\alpha,\beta]\Big{]}$ (9)
 $\phi(x_{i},{\mathcal{D}}_{2},\alpha,\beta)=\{j:x_{j}\in{\mathcal{D}}_{2},\gamma(x_{i},x_{j})\in[\alpha,\beta]\}$ (10)
-where $\gamma$ represents the similarity/distance function.
-For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$ following [5], while for high-resolution datasets, we use the SSCD score as $\gamma$.
+where $\gamma$ denotes the similarity/distance function.
+For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$, following [7], while for high-resolution datasets, we use the SSCD score as $\gamma$.
+The AMS averages the similarity scores across extracted images, ensuring that memorized images are not overlooked.
+In contrast, the UMS quantifies distinct memorized images by evaluating unique matches, thereby accounting for the uniqueness of the memorized images.
+By further categorizing the two scores into three levels—low, mid, and high—we obtain more comprehensive measurements for different types of memorization.
 In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity.
-(examples of different thresholds are shown in Figure 3) The thresholds for the normalized $L_{2}$ ditance [5] are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
-The AMS averages the similarity scores across generated images, ensuring that memorized images are not overlooked.
-In contrast, the UMS quantifies distinct memorized instances by evaluating unique matches, thereby accounting for the uniqueness of the memorized images.
-By further categorizing the two scores into three levels—low, mid, and high—we obtain more comprehensive measurements for different types of memorization.
-While [5, 7] introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is essential for assessing different types of copyright infringement, such as character or style copying [20, 27, 29].
-Additionally, the UMS considers the number of generated images $N_{G}$, which was overlooked in [5].
-The significance of $N_{G}$ lies in its non-linear impact on the UMS (see Appendix B for proof), indicating that UMS scores should not be compared across different values of $N_{G}$.
-Figure 3: A comparison between the original training images (top row) and generated images (bottom row) by our SIDE method.
-The matches are classified into three categories: low similarity (SSCD score $<$ 0.5), mid similarity (SSCD score between 0.5 and 0.6), and high similarity (SSCD score $>$ 0.6).
-This classification highlights varying degrees of semantic resemblance among the image pairs.
-Table 1: The AMS (%) and UMS (%) results at low, mid, and high levels.
-‘Random’ denotes the baseline that generates images directly using the target unconditional DPM[5], while OL-TI is a variant of SIDE which is not trained using TDKD.
-Dataset Method Low Similarity Mid Similarity High Similarity
-AMS(%) UMS(%) AMS(%) UMS(%) AMS(%) UMS(%)
-CelebA-HQ-FI Random 11.656 2.120 0.596 0.328 0.044 0.040
-OL-TI 2.649 0.744 0.075 0.057 0.005 0.005
-SIDE (Ours) 15.172 2.342 1.115 0.444 0.054 0.044
-CelebA-25000 Random 5.000 4.240 0.100 0.100 0.000 0.000
-OL-TI 0.164 0.152 0.000 0.000 0.000 0.000
-SIDE (Ours) 8.756 6.940 0.224 0.212 0.012 0.012
-CIFAR-10 Random 2.470 1.770 0.910 0.710 0.510 0.420
-OL-TI 2.460 1.780 0.800 0.680 0.420 0.370
-SIDE (Ours) 5.325 2.053 2.495 0.860 1.770 0.560
+(examples of different thresholds are shown in Figure 3) The thresholds for the normalized $L_{2}$ distance [7] are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
+Relation to Existing Metrics  While [7, 9] introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is crucial for assessing different types of copyright infringement, such as character or style copying [26, 34, 37].
+Additionally, the UMS considers the number of extracted images $N_{G}$, a factor overlooked in [7].
+The significance of $N_{G}$ lies in its non-linear impact on the UMS.
+Let $p_{\gamma}(i)$ be the probability that image $i$ is memorized, and $M$ be the size of the dataset.
+The expected number of uniquely memorized images is given by: $\mathbb{E}[N_{\mathrm{umem}}]=\sum_{i=1}^{M}{1-\left(1-p_{\gamma}\left(i\right)\right)^{N_{G}}}$, revealing that varying $N_{G}$ affects UMS non-linearly.
+This means that we should compare UMS under the same $N_{G}$.
+Note that the AMS and UMS scores are also different from the memorization divergence measure defined in Eq.
+(1), which is a distributional measure rather than an individual measure.
+Table 2: The coefficient and $R^{2}$ results of a fitted linear model between the number of classes vs. AMS/UMS. $R^{2}$ measures how well a regression model fits the data (1 is perfect fitting).
+Relationship Coefficient $R^{2}$
+Class vs. AMS 7.4 ×$\times 10^{-5}$ 0.637
+Class vs. UMS 6.1 ×$\times 10^{-5}$ 0.483
 5.3 Main Results
@@
 Here, ”TD” refers to the time-dependent classifier trained using our proposed TDKD method, ”TI” denotes the time-independent classifier, and ”OL” indicates training with original dataset labels.
-The ”Random” baseline generates images directly using the target unconditional DPM, as described in [5].
+The Random baseline generates images directly using the target unconditional DPM, as described in [7].
 We average the results across various values of $\lambda$ (defined in Eq.
-(3.2)), ranging from 5 to 9, with a detailed analysis provided in Section 4.3.
-It is important to note that $\lambda=0$ corresponds to the ”Random” baseline.
-For each $\lambda$, including $\lambda=0$, we generate 50,000 images for CelebA and 10,000 images for CIFAR-10 to validate the proposed SIDE method.
-This effort results in one of the most enormous generated image datasets for studying the memorization of DPMs.
-Table 2: This table presents the results of fitting a linear model to the relationship between the number of classes and both AMS and UMS.
-Relationship Coefficient (×$\times 10^{-5}$) Intercept $R^{2}$ Correlation Coefficient
-#Class vs. AMS 7.4 (positive) 0.115 0.637 0.80
-#Class and UMS 6.1 (positive) 0.090 0.483 0.70
+(3.2)), ranging from 5 to 10, with a detailed analysis provided in Section 5.3.
+It is important to note that $\lambda=0$ corresponds to the Random baseline.
+To evaluate SIDE, for each $\lambda$, including $\lambda=0$, we generate 51,200 images for CelebA-HQ-FI, 50,000 images for CelebA-25000, 10,000 images for CIFAR-10, 5,120 images for CelebA and 2,560 images for ImageNet.
+The results are presented in Table 4.
 Figure 4: Hyper-parameter ($\lambda$) analysis on CelebA-HQ-FI.
@@
 For other similarity levels, the best $\lambda$ for AMS and UMS is 13.
+Figure 5: The relationship between the number of classes and AMS/UMS. The fitted regression lines demonstrate a positive correlation in both metrics.
 Effectiveness of SIDE
-The AMS and UMS results for the three datasets are presented in Table 1 (results of 95-th percentile of SSCD are listed in Appendix F).
-As shown, SIDE is highly effective in extracting memorized data across all similarity levels, particularly on the CelebA-25000 dataset, a task previously deemed unfeasible due to its large scale.
-On CelebA-HQ-FI, SIDE increases mid-level AMS by 87% to 1.115% and UMS by 37% to 0.444%, with an average improvement of 20% across other levels.
-For the CelebA-25000 dataset, SIDE dramatically enhances AMS and UMS, achieving increases of 75% and 63% for low similarity and 124% and 112% for mid similarity.
-In the high similarity, SIDE excels at extracting memorized data.
-On CIFAR-10, SIDE also outperforms the baselines across all similarity levels.
-For low similarity, it achieves an AMS of 5.325%, more than double Random’s 2.470%, and a UMS of 2.05% compared to 1.780%.
-For mid similarity, SIDE reaches 2.495% AMS and 0.860% UMS, significantly higher than Random’s 0.910% AMS and 0.710% UMS, respectively.
-In the high similarity category, SIDE achieves 1.770% AMS and 0.560% UMS, well ahead of Random’s 0.510% AMS and 0.420% UMS.
+Table 4 summarizes the AMS and UMS results for the five datasets (results of 95th percentile SSCD are in Appendix C).
+SIDE shows strong performance in extracting training data across various datasets and similarity levels.
+Data extraction of unconditional DPM on large-scale datasets was considered impossible.
+But SIDE successfully extracts from these datasets.
+On CIFAR-10, SIDE outperforms baselines at all similarity levels.
+For low similarity, it achieves 5.320% AMS (vs.
+Random’s 2.470%) and 2.05% UMS (vs.
+1.780%).
+For mid similarity, SIDE reaches 2.490% AMS and 0.860% UMS, surpassing Random’s 0.910% AMS and 0.710% UMS.
+At high similarity, SIDE achieves 1.770% AMS and 0.560% UMS, outperforming Random’s 0.510% AMS and 0.420% UMS.
+On CelebA-HQ-FI, SIDE boosts mid-level AMS by 87% to 1.115% and UMS by 37% to 0.444%, with a 20% average improvement across other levels.
+SIDE could extract 900 high, 11,500 mid, and 151,720 low similarity images from 1 million samples.
+On CelebA-25000, SIDE improves AMS and UMS by 75% and 63% for low similarity, and 124% and 112% for mid similarity, excelling at high-similarity memorized data extraction.
+For CelebA and ImageNet, SIDE successfully extracts training data where baselines fail.
 Effectiveness of TDKD
 As shown in Table 4, time-independent classifiers perform significantly worse than their time-dependent counterparts on the two high-resolution CelebA datasets, achieving only about 10% of the effectiveness of classifiers trained using our TDKD method.
-This discrepancy arises because time-independent classifiers can provide accurate gradients only at the final timestep, whereas time-dependent classifiers deliver accurate gradients at each timestep.
+This discrepancy arises because time-independent classifiers produce accurate gradients only at the final timestep, whereas time-dependent classifiers deliver accurate gradients for each timestep.
 However, the performance gap narrows on the CIFAR-10 dataset due to its simplicity, consisting of only 10 classes.
-Thus, the gradients produced by the time-independent classifier are less prone to inaccuracy, which mitigates the limitations of the time-independent classifier.
-Memorized Images Are Not From the Classifiers
-There might be a concern that the extracted images may originate from the classifier rather than the target DPM.
-We argue that even if a time-dependent classifier could disclose its training images, these images would still originate from the target DPM because the classifier is distilled on the synthetic dataset generated by the target DPM.
-Furthermore, Table  1 demonstrates that employing a simple classifier reduces memorization compared to the baseline extraction method, indicating that the memorized images do not originate from the simple classifier as well.
+As a result, the gradients produced by the time-independent classifier are less susceptible to inaccuracies, thereby mitigating its inherent limitations.
+No Training Data Leakage From the Classifier
+A potential concern is that the extracted images might reflect artifacts introduced by the classifier rather than the target DPM itself.
+However, we argue that even if a time-dependent classifier were to expose its training images, those images would inherently originate from the training set of the target DPM, as the classifier is distilled on synthetic data generated by the DPM.
+Furthermore, as shown in Table 4, using a simple classifier reduces memorization compared to the baseline extraction method, indicating that the extracted memorized images do not primarily originate from the classifier.
+To further confirm our conjecture, we compare SIDE using classifiers trained on two different datasets: SIDE-A (an auxiliary dataset, disjoint from the DPM’s training data) and SIDE-O (the original training dataset).
+As shown in Table 4, SIDE-A and SIDE-O achieve similar results.
+This verifies that the classifier does not leak training data information, as SIDE only leverages it to label synthetic data rather than generate data samples.
+Figure 6: Validation of $N_{G}$ ’s significance
 Impact of the Classifier on SIDE
 The classifier used to train SIDE is associated with a certain number of classes.
-Here, we conduct experiments to explore the relationship between the number of classes (of the classifier) and the extraction performance at a low similarity level, using 1,200 images per class with $\lambda=5$ on the CelebA-HQ-FI dataset.
-As shown in Table  2 and Figure 5, there exists a strong positive correlation: as the number of classes increases, both AMS and UMS improve.
-In summary, increasing number of classes positively affects both AMS and UMS, with a stronger impact on AMS.
-The UMS values reported here differ significantly from Table  1.
-This is because we only generated 1,200 images per class here, whereas 50,000 images per class in the previous experiment.
-Increasing the number of classes improves AMS more than UMS because a higher class count enables the classifier to better differentiate fine details, leading to more accurate matches at low similarity levels.
-UMS is less affected since it relies more on the diversity across images, which is constrained by the smaller number of images per class in this experiment.
-Figure 5: Scatter plots showing the relationship between the number of classes and two performance metrics, AMS (left) and UMS (right).
-The fitted regression lines demonstrate a positive correlation in both cases.
+Here, we explore the relationship between the number of classes (of the classifier) and the extraction performance at a low similarity level, using 1,200 images per class with $\lambda=5$ on the CelebA-HQ-FI dataset.
+As shown in Table 2 and Figure 5, the number of classes exhibits a strong positive correlation with both AMS and UMS.
+Increasing the number of classes enhances both metrics, with a more pronounced effect on AMS.
+The UMS values here differ significantly from those in Table 4 due to the smaller sample size per class (1,200 images) in this experiment.
+The stronger improvement in AMS arises because the classifier captures fine-grained details more effectively with more classes, leading to more accurate matches at lower similarity levels.
+In contrast, UMS is less affected, as it depends more on overall image diversity, which is constrained by the smaller sample size per class in this setting.
+Importance of $N_{G}$ in UMS
+The number of uniquely memorized samples in a dataset of size $M$ can be formulated as $\sum_{i=1}^{M}{1-\left(1-k_{i}^{N_{G}}\right)}$, where $k_{i}$ denotes the probability the $i$-th sample is extracted per trial.
+To empirically verify the importance of $N_{G}$ , we generate 1 million samples using a DPM trained on CelebA-HQ-FI.
+As shown in Figure 6, the theoretical and empirical results align closely, confirming that $N_{G}$ non-linearly influences UMS.
 Hyper-parameter Analysis
@@
 As $\lambda$ increases, the diffusion models become conditional, and according to Theorem 1, the memorization effect will be triggered.
-However, when $\lambda$ is too large, the generated images will overfit the classifier’s decision boundaries, leading to a low diversity and ignoring the data distribution.
+However, when $\lambda$ becomes excessively large, the generated images will overfit the classifier’s decision boundaries, resulting in reduced diversity and a failure to accurately reflect the underlying data distribution.
 Consequently, the memorization score decreases.
-5 Theoretical Understanding for Memorization in Conditional Generative Models
-In this section, we introduce a metric to quantify the memorization capacity of generative models and provide a theoretical explanation for the data memorization in conditional diffusion models and why SIDE is effective.
-5.1 Memorization Metric
-Intuitively, the memorization of fixed training data points (i.e., point-wise memorization) can be quantified by the degree of overlap between the generated distribution and the distributions centered at each training data point.
-Given a generative model and training dataset, we propose the following memorization metric to quantify the memorization of the training data points in the model.
-Definition 1 (Point-wise Memorization).
-Given a generative model $p_{\theta}$ with parameters $\theta$ and training dataset ${\mathcal{D}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, degree of memorization in $p_{\theta}$ of ${\mathcal{D}}$ is defined as:
-${\mathcal{M}}_{p}({\mathcal{D}};p_{\theta})=\frac{1}{|{\mathcal{D}}|}\sum_{{\bm{x}}_{i}\in{\mathcal{D}}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}}$ (9)
-where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ denotes the $i$-th training sample, $N$ is the total number of training samples, $p_{\theta}({\bm{x}})$ represents the probability density function (PDF) of the generated samples, and $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the probability distribution centered at training sample ${\bm{x}}_{i}$ within a small radius $\epsilon$.
-A straightforward choice for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the Dirac delta function centered at training data point ${\bm{x}}_{i}$: $q({\bm{x}},{\bm{x}}_{i},\epsilon)=\delta({\bm{x}}-{\bm{x}}_{i})$.
-However, this would make Eq.
-(9) uncomputable as the Dirac delta function is zero beyond the $\epsilon$-neighborhood.
-Alternatively, we could let $q({\bm{x}},{\bm{x}}_{i},\epsilon)\sim{\mathcal{N}}(x_{i};\epsilon I)$.
-Note that in Eq.
-(9), a smaller value of ${\mathcal{M}}_{p}({\mathcal{D}};\theta)$ indicates more overlap between the two distributions and thus more memorization.
-As $\epsilon$ tends to 00, the measured memorization becomes more accurate, with the limit characterizing the memorization capability of $p_{\theta}$.
-5.2 Theoretical Analysis
-Based on the point-wise memorization metric, we perform a theoretical understanding that explains why conditional generative models achieve better memorization capacity.
-Our analysis is based on the concept of informative labels, which refers to information that can partition a dataset into various disjoint subsets.
-And we demonstrate that provided with an informative label, the conditional generative models achieve stronger memorization capacity.
-Informative Labels
-The concept of informative labels has been previously discussed as class labels [13].
-In this work, we introduce a more general definition that encompasses both class labels and random labels as special cases.
-We formally define an informative label as follows.
-Definition 2 (Informative Label).
-Let ${\mathcal{Y}}$ be an attribute of the data taking values in $\{y_{i}\}_{i=1}^{C}$.
-We refer to ${\mathcal{Y}}$ as an informative label if we can partition the dataset into several mutually disjoint subsets $\{{\mathcal{D}}_{i}\}_{i=1}^{C}$ based on the values of the attribute ${\mathcal{Y}}$.
-Here, the informative labels are not limited to the conventional class labels; they can also be text captions, features, or cluster information that can be used to group the training samples into subsets.
-The above definition states that an information label should be able to differentiate a subset of samples from others.
-An extreme case is that all samples have the same label; in this case, the label is not informative.
-According to our definition, class-wise and random labels are special cases of informative labels.
-Informative labels can be either explicit like class/random labels and text captions, or implicit like salient features or clusters.
-Main Results
-Here we introduce our main theorem on the memorization mechanism of conditional generative models and explain why SIDE is effective.
-We use ${\mathcal{D}}_{i}$ to denote the subset consisting of data with informative label ${\mathcal{Y}}=y_{i}$.
-In addition, we denote the distribution of data in the original dataset ${\mathcal{D}}$ by $p$.
-The distribution has expectation $\mu$ and covariance matrix $\Sigma$.
-Similarly, for each attribute $y_{i}$, we use $p_{i}$ to denote the distribution of the data with this attribute, and use $\mu_{i}$ and $\Sigma_{i}$ to denote its expectation and covariance matrix.
-Theorem 1.
-If a generative model $p_{\theta_{i}}$ matches the target distribution $p_{i}$ well for each $i$, i.e.
-$TV(p_{i},p_{\theta_{i}})=0$, and the data distribution for some informative label $y_{c}$ satisfies $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$, we have with probability 1:
-$\lim_{|{\mathcal{D}}_{c}|\to\infty}\lim_{\epsilon\to 0}\frac{{\mathcal{M}}_{p}({\mathcal{D}}_{c},p_{\theta_{c}})}{{\mathcal{M}}_{p}({\mathcal{D}}_{c},p_{\theta})}\leq 1$ (10)
-where $p_{\theta_{i}}$ denotes the distribution of generated data conditioned on ${\mathcal{Y}}=y_{i}$.
-The proof can be found in Appendix A.
-Theorem 1 states that any form of information labels can enhance memorization capacity in conditional generative models, including conventional class labels and random labels.
-The condition $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$ is reasonable because it ensures that the spread (variance) of the data within each class is not excessively large.
-$\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{c}}{\left({\bm{x}}_{i}-\mu_{c}\right)^{T}\left({\bm{x}}_{i}-\mu_{c}\right)}}{|{\mathcal{D}}_{c}|}\leq||\mu-\mu_{c}||_{2}^{2}$ (11)
-To understand this, we look at the derivation Eq.
-(11) from $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$.
-For a well-separated class, this squared distance should provide an upper bound on how spread out the class data can be.
-If $tr(\Sigma_{c})$ were large relative to $||\mu-\mu_{c}||_{2}^{2}$, it would imply that data samples represented by informative label $c$ are highly spread out, making it harder for the class labelled with $c$ to be well-formed and distinct.
-Thus, the condition ensures that the data samples form tight clusters around their means, making it reasonable for the class to be well-defined and easily detachable.
-This also explains why SIDE works.
-Chen et al.
-[8] found that unconditional DPMs learn clustered representations from training data.
-The data samples within each class should not be widely dispersed for effective clustering.
-Then, the samples should satisfy Eq.
-11.
-So, although these clustered representations are implicit, we could extract training data from unconditional DPMs if we can convert them from implicit informative labels into explicit labels.
-When conditioning on the explicit labels, the DPMs will exhibit more memorization effects indicated by Theorem 1, which makes room for data extraction attacks.
-6 Limitations:
-SIDE has certain limitations that warrant further research.
-SIDE mainly uses human-labelled data to extract data, which is scarce and limits its practical use.
-Future work could investigate hybrid approaches for generating implicit labels, such as employing unsupervised clustering or utilizing pretrained models to label synthetic datasets.
+Limitation
+While showing promising results on unconditional DPMs, our SIDE still relies on a reliable pseudo-labeler, such as a classifier trained on the auxiliary data or an accurate clustering method.
+We plan to reduce this dependency by exploring pre-trained image encoders in future work.
+Additionally, SIDE requires access to the original training data for performance evaluation, which may not be feasible for commercial models.
+Future research could focus on developing evaluation methods that do not depend on the original training data.
 6 Conclusion
-In this paper, we distinguish the difference between explicit labels and implicit labels and propose a novel method, SIDE, to extract training data from unconditional DPMs by constructing surrogate conditions.
-The core of SIDE lies in training a time-dependent classifier using TDKD technique.
-We empirically validate effectiveness of SIDE on subsets of CelebA and CIFAR-10 datasets with two new memorization scores: AMS and UMS.
-In addition, we introduce a point-wise memorization metric to quantify memorization effects in DPMs for theoretical analysis.
-Finally, we introduce theoretical understanding of memorization in conditional generative models, offering a generalized definition of informative labels, clarifying random labels can also be informative and explaining why SIDE works.
-Our work aims to deepen understanding of memorization in DPMs and inspire further methods to mitigate memorization.
+In this paper, we proposed a novel method SIDE for extracting training data from unconditional diffusion probabilistic models (DPMs) by constructing surrogate conditions.
+At its core, SIDE trains a time-dependent classifier to create these surrogate conditions.
+We empirically validated the effectiveness of SIDE on datasets of varying scales, including CelebA, ImageNet, and CIFAR-10.
+Additionally, we provide a theoretical analysis to explain the role of information labels in facilitating memorization in DPMs.
+Our work highlights the potential for memorization in unconditional DPMs and demonstrates that such memorized images can be extracted by constructing surrogate conditions.
+Our study advances the understanding of memorization mechanisms in DPMs and motivates the development of more effective memorization mitigation strategies, such as regularizing deep representations to be less informative.
 Supplementary Material
 Broader Impacts
-Our work introduces a memorization metric that not only quantifies memorization effects in diffusion models but also extends to deep learning models more broadly.
+Our work not only quantifies memorization effects in diffusion models, but also extends to deep generative models.
 This contribution is significant in enhancing our understanding of when and how models memorize training data, which is critical for addressing concerns about data privacy and model robustness.
@@
 Appendix A Proof of Theorem1
-By the assumption $TV(p_{i},p_{\theta_{i}})=0$ in the theorem, we can replace $p_{\theta_{c}}$ with $p_{c}$ and $p_{\theta}$ with $p$ to simplify the notation without compromising correctness.
-Firstly, we compute $\mathcal{M}_{point}(\mathcal{D}_{c};p_{c})$ here:
+By the assumption $TV(p_{i},p_{\theta_{i}})=0$ , we can replace $p_{\theta_{i}}$ with $p_{i}$ and $p_{\theta}$ with $p$ to simplify the notation without compromising correctness. Hence, we have:
+$\mathcal{M}(\mathcal{D}_{i};p_{\theta_{i}},\epsilon)-\mathcal{M}(\mathcal{D}_{i};p_{\theta},\epsilon)=\frac{1}{N}\sum_{x_{k}\in\mathcal{D}_{i}}\int\mathcal{N}(x|x_{k},\epsilon^{2}I)\log{\frac{p(x)}{p_{i}(x)}}dx$ (11)
+By the Strong Law of Large Numbers, as $N\to\infty$, with probability 1:
+$\frac{1}{N}\sum_{x_{k}\in\mathcal{D}_{i}}\mathcal{N}(x|x_{k},\epsilon^{2}I)$ $\to\mathbb{E}_{y\sim p_{i}}\left[\frac{1}{(2\pi\epsilon^{2})^{d/2}}\exp{(-\frac{||y-x||^{2}}{2\epsilon^{2}})}\right]$ (12)
 […]
-In the following, we will omit the expression ”almost surely”. Similarly, for $\mathcal{M}_{p}(\mathcal{D}_{c};p)$:
-$\begin{split}\mathcal{M}_{p}(\mathcal{D}_{c};p)&=o(\frac{1}{\epsilon})+\frac{1}{2\epsilon|\mathcal{D}_{c}|}\sum_{i=1}^{|\mathcal{D}_{c}|}(tr(\Sigma)+(\mu-z_{i})^{T}(\mu-z_{i}))\\ &\rightarrow o(\frac{1}{\epsilon})+\frac{1}{2\epsilon}(tr(\Sigma)+tr(\Sigma_{c})+||\mu-\mu_{c}||^{2}_{2})\end{split}$ (13)
-By definition, $p(x)=\sum_{k}\pi_{k}p_{k}(x)$ ,where $\pi_{k}$ denotes the probability that the informative label $\mathcal{Y}$ equals $y_{k}$.
-$\begin{split}\Sigma=\underset{x\sim p(x)}{\mathrm{Cov}}(x)&=\mathbb{E}_{p}[xx^{T}]-\mathbb{E}_{p}[x]\mathbb{E}_{p}[x]^{T}\\ &=\sum_{k}\pi_{k}\mathbb{E}_{p_{k}}[zz^{T}]-\mu\mu^{T}\\ &=\sum_{k}\pi_{k}(\mu_{k}\mu_{k}^{T}+\Sigma_{k})-\mu\mu^{T}\\ &=\mathbb{E}_{\pi}[\Sigma_{k}]+\mathbb{E}_{\pi}[(\mu-\mu_{k})(\mu-\mu_{k})^{T}]\end{split}$ (14)
-Finally, we have
-$\begin{split}\lim_{|\mathcal{D}|_{c}\rightarrow\infty}\lim_{\epsilon\rightarrow 0}\frac{\mathcal{M}_{p}(\mathcal{D}_{c};p_{c})}{\mathcal{M}_{p}(\mathcal{D}_{c};p)}&=\frac{2tr(\Sigma_{c})}{tr(\Sigma_{c})+\mathbb{E}_{\pi}[tr(\Sigma_{k})]+\mathbb{E}_{\pi}[||\mu-\mu_{k}||^{2}_{2}]+||\mu-\mu_{c}||^{2}_{2}}\end{split}$ (15)
-Here we use $\mathbb{E}_{\pi}[tr(\Sigma_{k})]$ and $\mathbb{E}_{\pi}[||\mu-\mu_{k}||_{2}^{2}]$ to denote $\sum_{k}\pi_{k}tr(\Sigma_{k})$ and $\sum_{k}\pi_{k}||\mu-\mu_{k}||_{2}^{2}$.
-Clearly, by Jensen’s inequality we have:
-$\mathbb{E}_{\pi}[\lim_{|\mathcal{D}_{c}|\rightarrow\infty}\lim_{\epsilon\rightarrow 0}\frac{\mathcal{M}_{p}(\mathcal{D}_{c};p_{c})}{\mathcal{M}_{p}(\mathcal{D}_{c};p)}]\leq 1$ (16)
-The equality holds only when $\mu_{k}$ and $tr(\Sigma_{k})$ are the same for all $k$. And under the assumption $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$, we have:
-$\lim_{|\mathcal{D}_{c}|\rightarrow\infty}\lim_{\epsilon\rightarrow 0}\frac{\mathcal{M}_{p}(\mathcal{D}_{c};p_{c})}{\mathcal{M}_{p}(\mathcal{D}_{c};p)}\leq 1$ (17)
-Hence we have completed the proof of Theorem1. ∎
-Appendix B Explanation UMS’s importance in fair comparison
-By highlighting the significance of $N_{G}$, UMS allows a fair comparison among extraction methods.
-We explain this both theoretically and experimentally.
-Let $N_{\mathrm{umem}}$ denote the number of uniquely memorized images, $M$ the dataset size, and $p_{\gamma}(i)$ the probability that image $i$ is memorized.
-To find $\mathbb{E}(N_{\mathrm{umem}})$, we introduce a new variable $I_{i}$, representing the generation of image $i$ in $N_{G}$ trials:
-$p(I_{i})=1-\left(1-p_{\gamma}(i)\right)^{N_{G}}$ (18)
-Using the linearity of expectation, we derive: $\mathbb{E}\left(N_{\mathrm{mem}}\right)$
-$\mathbb{E}\left(N_{\mathrm{umem}}\right)$ (19)
-$=\mathbb{E}\left(I_{1}\right)+\mathbb{E}\left(I_{2}\right)+\cdots+\mathbb{E}\left(I_{M}\right)$ (20)
-$=\sum_{i=1}^{M}{1-\left(1-p_{\gamma}\left(i\right)\right)^{N_{G}}}$ (21)
-The importance of $N_{G}$ lies in its impact on the non-linear expectation of uniquely memorized images.
-Comparing UMS across different $N_{G}$ values is flawed because varying $N_{G}$ leads to different outcomes, underscoring the need for consistent $N_{G}$ values to ensure fair comparisons.
-In [5], the significance of $N_{G}$ was overlooked, as they only reported the number of uniquely extracted images.
-Our theoretical analysis accurately aligns with behavior observed in the experimental data in Fig 6.
-Appendix C Refinement ResNet block (Fig. 7)
-Figure 6: Validation of $N_{G}$ ’s significance
+[…]
+Now, we show term $L\to 0$ as $\epsilon\to 0$.
+By the continuity of $p_{i}$, for any $\eta>0$, select $r>0$ such that $|p_{i}(y)-p_{i}(x)|<\eta,\ \ \forall\ ||y-x||<r$.
+Then we can decomposite $L$ into two parts:
+$|L|$ $\leq\frac{1}{(2\pi\epsilon^{2})^{d/2}}\left(\int_{||y-x||<r}+\int_{||y-x||\geq r}\right)|p_{i}(y)-p_{i}(x)|\exp{(-\frac{||y-x||^{2}}{2\epsilon^{2}})}dy$ (15)
+[…]
+$\to\eta\text{ \ \ as }\epsilon\to 0$ (17)
+Due to the arbitrariness of $\eta$, $L\to 0$ as $\epsilon\to 0$, we obtain
+$\lim_{\epsilon\to 0}\lim_{|{\mathcal{D}}_{i}|\to\infty}({\mathcal{M}}({\mathcal{D}}_{i};p_{\theta_{i}},\epsilon)-{\mathcal{M}}({\mathcal{D}}_{i};p_{\theta},\epsilon))=\int p_{i}(x)\log{\frac{p(x)}{p_{i}(x)}}dx=-D_{\mathrm{KL}}(p_{i}||p)\leq 0$ (18)
+Appendix B Refinement ResNet block (Figure 7)
 Figure 7: Refinement ResNet block with time-dependent module integration.
@@
 Image $x_{BN}$ is the image processed after the first Batch Normalization Layer.
-The integration of the time module directly after batch normalization within the network architecture is a reasonable design choice rooted in the functionality of batch normalization itself.
+This section elaborates on our time module’s design principles and architectural rationale, which strategically integrates temporal dynamics into normalized feature spaces through a post-batch normalization framework.The integration of the time module directly after batch normalization within the network architecture is a reasonable design choice rooted in the functionality of batch normalization itself.
 Batch normalization standardizes the inputs to the network layer, stabilizing the learning process by reducing internal covariate shifts.
-By positioning the time module immediately after this normalization process, the model can introduce time-dependent adaptations to the already stabilized features.
+The model can introduce time-dependent adaptations to the already stabilized features by positioning the time module immediately after this normalisation process.
 This placement ensures that the temporal adjustments are applied to a normalized feature space, thereby enhancing the model’s ability to learn temporal dynamics effectively.
@@
 It allows the network to maintain a streamlined architecture while still gaining the necessary capacity to handle time-varying inputs.
-Appendix D Results on Classifier Choice
-The analysis of classifier performance across varying numbers of classes reveals interesting patterns for both AMS and UMS.
-As shown in Figure 5, the scatter plots highlight a positive relationship between the number of classes and the performance metrics.
-For AMS, the regression line suggests a stronger relationship (R² = 0.637), implying that as the number of classes increases, the AMS metric improves with a moderately strong association.
-In contrast, the relationship between the number of classes and UMS, while still positive, exhibits a slightly weaker connection (R² = 0.483).
-These results suggest that classifier performance, particularly as measured by AMS, benefits more significantly from an increase in the number of classes compared to UMS.
-The shaded regions in the plots represent the 95% confidence intervals, indicating the range of uncertainty around the fitted regression lines.
-Overall, the findings imply that the choice of classifier could have a notable impact on AMS , with a less pronounced but still meaningful effect on UMS.
-Appendix E Pseudocode for SIDE method
-The following Algorithm 1 outlines the detailed steps of the SIDE method for extracting training data from unconditional diffusion models.
-This algorithm combines the construction of implicit informative labels, the training of a time-dependent classifier, and the conditional generation process to extract valuable training samples effectively.
-Algorithm 1 SIDE Method for Extracting Training Data from Unconditional Diffusion Models
-1:Input:
-  • Unconditional Diffusion Probabilistic Model (DPM) $p_{\theta}(\bm{x})=\mathcal{N}(\bm{\mu},\Sigma_{p_{\theta}})$
-  • Pre-trained Classifier $p_{\theta}(y_{I}\mid\bm{x})$
-  • Hyperparameter set $\mathcal{S}_{\lambda}$
-  • Number of generated samples per $\lambda$, $N_{G}$
-  • Target label $y=c$
-2:Output: Extracted training data $\mathcal{D}_{extracted}$
-3:Construct Implicit Informative Labels
-4:Input: Unconditional DPM $p_{\theta}(\bm{x})$, Classifier $p_{\theta}(y_{I}\mid\bm{x})$
-5:Define sampling process with implicit labels:
-$\mathrm{d}\bm{x}$ $=\left[f(\bm{x},t)-g(t)^{2}\left(\nabla_{\bm{x}}\log p_{\theta}^{t}(\bm{x})+\lambda\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}\mid\bm{x})\right)\right]\mathrm{d}t+g(t)\mathrm{d}\bm{w}$ (22)
-6:Calibrate Classifier Output
-7:Adjust classifier probabilities using power prior:
-$p^{t}_{\theta}\left(\bm{x}\mid y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}\mid\bm{x}\right)p^{t}_{\theta}\left(\bm{x}\right)$ (23)
-8:Train Time-Dependent Classifier via TDKD
-9:Input: Pre-trained Classifier $p_{\theta}(y_{I}\mid\bm{x})$, Synthetic Dataset $\mathcal{D}_{synthetic}$
-10:Output: Time-Dependent Classifier $p_{\theta}^{t}(y_{I}\mid\bm{x}_{t})$
-11:Initialize Time-Dependent Classifier architecture with time-dependent modules
-12:Generate synthetic dataset using DPM:
-$\mathcal{D}_{synthetic}=\{\bm{x}^{(i)}\}_{i=1}^{N_{synthetic}}\sim p_{\theta}(\bm{x})$ (24)
-13:Generate pseudo labels using pre-trained classifier:
-$y_{I}^{(i)}=p_{\theta}(y_{I}\mid\bm{x}^{(i)})$ (25)
-14:Train Time-Dependent Classifier by minimizing KL divergence:
-$\mathcal{L}_{distil}=D_{KL}\left(p_{\theta}(y_{I}\mid\bm{x}^{(i)})\,\|\,p_{\theta}^{t}(y_{I}\mid\bm{x}_{t}^{(i)})\right)$ (26)
-15:Overall SIDE Procedure
-16:Input: Trained Time-Dependent Classifier $p_{\theta}^{t}(y_{I}\mid\bm{x}_{t})$, Target Label $y=c$, Hyperparameter set $\mathcal{S}_{\lambda}$, Number of samples $N_{G}$
-17:Initialize empty dataset $\mathcal{D}_{extracted}$
-18:for each $\lambda\in\mathcal{S}_{\lambda}$ do
-19:     for $i=1$ to $N_{G}$ do
-20:         Sample $\bm{x}_{0}$ conditioned on $y=c$ using Eq. (22)
-21:         Compute gradient of cross-entropy loss:
-$\nabla_{\bm{x}_{t}}\mathcal{L}_{CE}(c,p_{\theta}^{t}(y\mid\bm{x}_{t}))$ (27)
-22:         Reverse diffusion process using computed gradient
-23:         Generate similarity score for $\bm{x}_{t}$
-24:         Append $\bm{x}_{t}$ to $\mathcal{D}_{extracted}$ with similarity score
-25:     end for
-26:end for
-27:Evaluate Attack Performance
-28:Compute evaluation metrics on $\mathcal{D}_{extracted}$
-29:return $\mathcal{D}_{extracted}$
-Appendix F SSCD-related metrics
-In our evaluation of the proposed method, it is essential to include a discussion of the 95th percentile SSCD metric alongside our newly introduced metrics, AMS and UMS.
-While 95th percentile SSCD metric has its limitations, it still serves as a useful reference point for assessing the relative similarity between extracted images and the original training dataset.
+Appendix C SSCD-related metrics
+To comprehensively evaluate the proposed method, it’s crucial to present a discussion of the 95th percentile SSCD metric alongside our newly introduced metrics, AMS and UMS.
+Although the 95th percentile SSCD metric has its limitations, it still serves as a useful reference point for assessing the relative similarity between extracted images and the original training dataset.
 As shown in Table 3, we provide a comprehensive comparison of performance across different datasets.
-In addition, we report detailed metrics in the context of our experiments on the CelebA-HQ-FI dataset, further illustrating our approach’s effectiveness.
-Table 3: Performance Comparison on CelebA-HQ-FI and CelebA-25000 Datasets
-Dataset Method Top 0.01% Top 0.05% Top 0.1% Top 0.5% Top 1.0% Top 5.0%
-CelebA-HQ-FI Uncond 0.636 0.604 0.584 0.524 0.498 0.433
-SIDE(ours) 0.680 0.639 0.618 0.567 0.544 0.485
-CelebA-25000 Uncond 0.545 0.518 0.505 0.471 0.455 0.404
-SIDE(ours) 0.585 0.554 0.539 0.506 0.491 0.450
-Appendix G Extended SIDE for Text-to-Image Diffusion Models
-We extend the SIDE method for a text-to-image diffusion model in a fine-tuning scenario.
+Table 3: Evaluation of Data Extraction Performance using SSCD Percentiles.
+The table shows SSCD scores (Top 0.05% to Top 5.0%) for the Random baseline and SIDE methods (SIDE-A, SIDE-O) across CelebA-HQ-FI, CelebA-25000, CelebA, and ImageNet datasets.
+Dataset Method Top 0.05% Top 0.1% Top 0.5% Top 1.0% Top 5.0%
+CelebA-HQ-FI Random 0.604 0.584 0.524 0.498 0.433
+SIDE-A 0.617 0.611 0.576 0.555 0.485
+SIDE-O 0.639 0.618 0.567 0.544 0.485
+CelebA-25000 Random 0.518 0.505 0.471 0.455 0.404
+SIDE-A 0.534 0.525 0.500 0.485 0.439
+SIDE-O 0.554 0.539 0.506 0.491 0.450
+CelebA Random 0.487 0.477 0.452 0.439 0.404
+SIDE-A 0.523 0.517 0.495 0.479 0.436
+SIDE-O 0.513 0.508 0.471 0.456 0.419
+ImageNet Random 0.330 0.316 0.287 0.275 0.250
+SIDE-A 0.512 0.493 0.378 0.342 0.283
+SIDE-O 0.462 0.419 0.362 0.342 0.291
+Appendix D Cluster-Based SIDE (SIDE-C)
+In the main text, we presented the SIDE algorithm using a classifier as a surrogate condition for simplicity and clarity in the core algorithm description.
+However, we also investigated the use of cluster-based surrogate conditions.
+This appendix details the formulation of the cluster-based SIDE approach and presents a performance analysis.
+In our cluster-based approach, we first extract feature dataset from the synthetic dataset generated by the target DPM using SSCD encoder.
+This extracted data serves as the basis for our clustering.
+We then apply k-means clustering to this SSCD-extracted dataset, partitioning the data into 100 distinct clusters.These cluster assignments are then treated as implicit labels.
+Instead of using a classifier’s predicted class probabilities to guide the diffusion process, we use the cluster labels to train the time-dependent classifier.
+In essence, we replace the class-conditional term in the gradient guidance with a cluster-conditional guidance term.
+This allows the diffusion model to be guided towards generating data points that are representative of specific clusters discovered within the SSCD-extracted data, effectively using these clusters as implicit, self-discovered categories for data extraction.
+Table 4: Performance analysis of SIDE with cluster-based surrogate conditions.
+This table compares AMS (%) and UMS (%) in percentage for different methods, including baselines (Random, OL-TI) and SIDE variations (classifier-based: SIDE-O, SIDE-A; cluster-based: SIDE-C), across low, mid, and high similarity levels on CIFAR-10 and CelebA-25000 datasets.
+Dataset Method Low Similarity Mid Similarity High Similarity
+AMS(%) UMS(%) AMS(%) UMS(%) AMS(%) UMS(%)
+CIFAR-10 Random 2.470 1.770 0.910 0.710 0.510 0.420
+OL-TI 2.460 1.780 0.800 0.680 0.420 0.370
+SIDE-A 5.250 2.020 2.300 0.880 1.620 0.640
+SIDE-O 5.320 2.050 2.490 0.860 1.770 0.560
+SIDE-C 7.830 2.730 3.830 1.190 2.610 0.760
+CelebA-25000 Random 5.000 4.240 0.100 0.100 0.000 0.000
+OL-TI 0.164 0.152 0.000 0.000 0.000 0.000
+SIDE-A 8.712 6.802 0.234 0.234 0.010 0.010
+SIDE-O 8.756 6.940 0.224 0.212 0.012 0.012
+SIDE-C 20.500 11.400 1.842 1.164 0.000 0.000
+SIDE-C (cluster-based SIDE) generally outperforms other methods, especially classifier-based SIDE variants (SIDE-O, SIDE-A), in data extraction, achieving significantly higher AMS and UMS on both CIFAR-10 and CelebA-25000, particularly at low and mid similarity levels.
+The cluster-based approach appears to offer a more effective guidance mechanism, potentially due to implicit labels derived from data-driven clusters.
+However, a notable exception is observed at high similarity for CelebA-25000, where SIDE-C’s performance dramatically drops to 0%, suggesting a potential failure mode, possibly related to cluster overfitting or interaction with strong similarity constraints, unlike the more consistent, though less performant, classifier-based SIDE.
+This highlights SIDE-C’s promise as a powerful data extraction technique, but also points to areas needing further investigation, particularly its robustness under high similarity conditions.
+Appendix E Extended SIDE for Text-to-Image Diffusion Models
+We extend the SIDE method for text-to-image diffusion models in a fine-tuning scenario.
 Instead of using a classifier as a surrogate condition to extract training data, we use random strings as surrogate conditions.
@@
 We name this method as Extended SIDE.
-G.1 Attack Setting
+E.1 Attack Setting
 The attack setting for the Extended SIDE method involves two primary components: the victim’s fine-tuning process and the attacker’s access to the fine-tuned model.
-Below, we provide a detailed description of each component.
-  • Victim Fine-Tuning:
-    The victim fine-tunes a text-to-image diffusion model, denoted as $p_{\theta}(\bm{x}\mid s)$, using a dataset of image-text pairs $\mathcal{D}_{train}=\{(s_{i},\bm{x}_{i})\}_{i=1}^{N}$.
-In this context:
-    • $s_{i}\in\mathcal{S}_{s}$ are random strings crafted by the attacker to serve as surrogate text prompts.
+In the following, we provide a detailed description of each component.
+  • Victim Fine-Tuning Process: The victim fine-tunes a text-to-image diffusion model, denoted as $p_{\theta}(\bm{x}\mid s)$, using a dataset of image-text pairs $\mathcal{D}_{train}=\{(s_{i},\bm{x}_{i})\}_{i=1}^{N}$.
+In this context $s_{i}\in\mathcal{S}_{s}$ are random strings crafted by the attacker to serve as surrogate text prompts.
 $\bm{x}_{i}\in\mathbb{R}^{d}$ represents the corresponding images associated with each random string.
@@
 The use of random strings as surrogate conditions aims to induce memorization within the model, causing it to overfit specific image-text associations during the fine-tuning process.
-  • Attacker’s Access:
-    After the fine-tuning phase, the attacker gains access to the fine-tuned model $p_{\theta}(\bm{x}\mid s)$. This access allows the attacker to perform the following actions:
-    • Input Manipulation: The attacker can input any text prompt $s\in\mathcal{S}_{s}$ into the model to generate corresponding images:
+  • Attacker’s Access: After the fine-tuning phase, the attacker gains access to the fine-tuned model $p_{\theta}(\bm{x}\mid s)$.
+Such access enables the attacker to perform Input Manipulation.
+The attacker can input any text prompt $s\in\mathcal{S}_{s}$ into the model to generate corresponding images:
     $\bm{x}\sim p_{\theta}(\bm{x}\mid s)$ (20)
   • Data Extraction: By systematically querying the model with the same set of random strings $\mathcal{S}_{s}$, the attacker can generate multiple images $\{\bm{x}_{i,j}\}_{j=1}^{N_{G}}$ for each string $s_{i}$.
-The goal is to identify images that closely resemble the memorized training samples:
+Our goal is to identify images that closely resemble the memorized training samples:
     $\bm{x}_{i,j}\sim p_{\theta}(\bm{x}\mid s_{i}),\quad j=1,\dots,N_{G}$ (21)
@@
 If $\sigma_{s_{i}}^{2}$ falls below a predefined threshold $\tau$, the attacker considers $\overline{\bm{x}}_{s_{i}}$ as a potential memorized training sample.
-  • Assumptions:
-    • Knowledge of Surrogate Strings: The attacker is aware of the set of random strings $\mathcal{S}_{s}$ used during the victim’s fine-tuning process.
+  • Assumptions: Knowledge of Surrogate Strings: The attacker is aware of the set of random strings $\mathcal{S}_{s}$ used during the victim’s fine-tuning process.
 Model Query Capability: The attacker can perform a large number of queries to the fine-tuned model with each random string $s_{i}$ to generate sufficient samples for variance analysis.
 No Access to Training Data: The attacker does not have direct access to the victim’s training dataset $\mathcal{D}_{train}$.
-  • Attack Goals:
-    • Data Reconstruction: To reconstruct or approximate the training images $\bm{x}_{i}$ by exploiting the model’s memorization of the random string-image associations.
+  • Attack Goals: Data Reconstruction: To reconstruct or approximate the training images $\bm{x}_{i}$ by exploiting the model’s memorization of the random string-image associations.
 Privacy Breach: To demonstrate the potential privacy risks associated with fine-tuned diffusion models by extracting sensitive or proprietary training data.
-G.2 Extended SIDE Method
+E.2 Extended SIDE Method
 The Extended SIDE method leverages random strings as surrogate conditions to directly extract training images from a fine-tuned diffusion model.
@@
 12:return $\mathcal{D}_{\text{extracted}}$
-The Extended SIDE method operates on the principle that random strings, when used as surrogate conditions during fine-tuning, become tightly coupled with specific training images.
-This follows the results of our theoretical understanding in Theorem 1.
+The Extended SIDE method is based on the principle that random strings, when used as surrogate conditions during fine-tuning, become strongly associated with specific training images.
+This follows from the results of our theoretical understanding in Theorem 1.
 The attacker exploits this behavior to extract training data by performing the following steps:
@@
 This method effectively leverages the deterministic mapping between random strings and training images induced by memorization, allowing the attacker to reconstruct the training dataset without direct access.
-G.3 Algorithmic Implementation
+E.3 Algorithmic Implementation
 The Extended SIDE method is formalized in Algorithm 2.
@@
 The Extended SIDE is shown in Algorithm 2
-G.4 Experiments
+E.4 Experiments
 To verify the effectiveness of the Extended SIDE, we conduct experiments on the subset of the Laion-2B dataset, finetuning the Stable Diffusion v1.5 using the LORA[24] method.
 We use the following metrics to evaluate the attack results: mean SSCD (M-SSCD), AMS (mid-similarity) and LIPIPS.
-The results are shown in Tab.
-4.
+The results are shown in Table 5.
 We use the generated images of original prompts as the Baseline method.
 The extracted examples are shown in Figure 8.
-As the results show, our method could effectively extract training data from fine-tuned DPMs.
 Table 5: Performance Comparison for Extended SIDE method
@@
 Extended SIDE(ours) 0.467 0.672 0.809
-As presented in Tab. 4, the Extended SIDE method outperforms the baseline across all evaluation metrics:
+As presented in Table 5, the Extended SIDE method outperforms the baseline across all evaluation metrics:
   • M-SSCD: The SIDE method achieves an M-SSCD of 0.467 compared to the baseline’s 0.028.